Fontname

version 2.1, July 1996 Filenames for TEX fonts

K. Berry

1 Introduction

This is Fontname 2.1, a naming scheme for (the base part of) external TEX font filenames. This makes at most eight-character names from (almost) arbitrarily complex font names, thus helping portability of TEX documents. We point out some shortcomings of the scheme, and ways of dealing with them.

Please send additions, corrections, or other comments to me at kb@mail.tug.org, or to the

mailing list ‘tex-fonts@mail.tug.org’. Email ‘tex-fonts-request@mail.tug.org’ to join the

list.

Besides this document, the Fontname distribution includes various encoding files, mapping files, and utilities. See the file ‘README’ for details. You can retrieve the most up-to-date version of these lists by anonymous ftp as ‘ftp://ftp.tug.org/tex/fontname’. This directory is mirrored on all CTAN hosts (‘ftp.dante.de’, ‘ftp.tex.ac.uk’, and their mirrors, such as ‘ftp.cdrom.com’) in the directory ‘tex-archive/doc/fontname’. For a list of CTAN mirrors, finger ‘ctan@ftp.tug.org’.

This document is uncopyrighted and may be used freely.

1.1 History

The original inspiration for Fontname was Frank Mittelbach and Rainer Schoepf’s article in TUGboat 11(2) (June 1990), which led to my article in TUGboat 11(4) (November 1990), pages 512–519. Mittelbach wrote another article criticizing the scheme below in TUGboat 13(1) (April 1992), pages 51–53; most of his points are well-taken, but I saw no alternative then, and see no alternative now. Others of his points are addressed in the appropriate sections below.

In August 1995, Fontname version 2 was released, concomitantly with a new release of the PostScript fonts for use with LaTEX and Dvips. After several years of experience, several modifications of the original distribution seemed desirable: more precise specification of encodings (see Section 2.4 [Variants], page 14); machine-readable mapping files (the ‘*.map’ files in the distribution) and standardizing directory names as well as font filenames (included in ‘supplier.map’ and

‘typeface.map’).

In Fontname version 1, some encoding variant characters (‘0’, ‘2’, etc.) have been reassigned to a ‘7’ or ‘8’ sequence (see Section 2.4 [Variants], page 14). Although this is an incompatible change, the advantages of (1) consistency in naming encoding variants, and (2) having some free variants makes it worthwhile. Also, not so coincidentally, new metrics for PostScript fonts in the Cork encoding were issued—same encoding, different metrics. So a new name served a practical purpose as well.

Many people have contributed to this work. In particular, I gratefully acknowledge Nelson Beebe, Barbara Beeton, Rocky Bernstein, Thierry Bouche, Damian Cugley, Berthold K.P. Horn, Alan Jeffrey, Russell Lang, Pierre MacKay, Sebastian Rahtz, Jean Rivlin, and Tom Rokicki.

1.2 References

Related files and distributions:

• The TEX Directory Structure (TDS) standard, available from (‘ftp://ftp.tug.org/tex/tds’ and ‘CTAN:/tds’) specifies a directory tree for all TEX library files. The directory names given in Section 2.1 [Suppliers], page 3 and Section 2.2 [Typefaces], page 4 are intended to mesh with that standard.

• The ‘modes.mf’ file (‘ftp://ftp.tug.org/tex/modes.mf’) contains recommended mode

names to use as directory names.

• The PSfonts distribution (‘CTAN:/fonts/psfonts’) contains TEX support for many fonts, both commercially available and free.

• The Fontinst distribution (‘CTAN:/fonts/utilities/psfonts’) supports quite general creation of virtual fonts, named according to the scheme given here.

• The Dvips translator (original from ‘ftp://labrea.stanford.edu/pub/dvips*.tar.gz’ and
variation from ‘ftp://ftp.cs.umb.edu/pub/tex/dvipsk.tar.gz’) supports resident and

downloadable PostScript fonts. Dvips was the first real application of Fontname 1. Dvips also includes Afm2tfm, another program that can create virtual fonts.

• The Dviljk processor (‘ftp://ftp.cs.umb.edu/pub/tex/dviljk.tar.gz’) contains TEX sup-

port for the fonts built into the LaserJet 4.

• ‘Designing New Typefaces with Metafont’, by Richard Southall, Stanford Computer Science Department Technical Report STAN-CS-85-1074, September 1985, for a comprehensive terminology of font design.

2 Filenames for fonts

We must limit ourselves to eight character names, for compatibility with DOS filesystems and the ISO 9660 standard used for CD-ROM distribution. Names may consist of only the letters (monocase a-z), numerals (0-9), and underscore.

Here is the basic division of the eight characters (the spaces here are merely for readability):

S TT W [V…] [N] [E] [DD] where

S represents the supplier of the font.

TT represents the typeface name.

W represents the weight.

V… represents the variant(s), and is omitted if both it and the width are normal. Many fonts have more than one variant.

N represents the encoding, and is omitted if the encoding is nonstandard. Encodings are subsumed in the section on variants (see Section 2.4 [Variants], page 14).

E represents the width (“expansion”), and is omitted if it is normal.

DD represents the design size (in decimal), and is omitted if the font is linearly scaled. Mittelbach in TUGboat 13(1) proposes using hexadecimal or base-36 notation. I don’t think the increased range makes up for the unreadability of the common sizes (e.g., 10 pt fonts would have a suffix ‘a’ (in hex), or ‘j’ (in base 36)).

The weight, variants, and width are probably all best taken from the original name of the font, instead of trying to relate them to some external standard.

2.1 Suppliers

The supplier is the source of a font, typically a (digital) type foundry.

You should use the supplier letter which matches the supplier you obtained the font from, not the original source; for example, Avant Garde was designed by Herb Lubalin for ITC, but Adobe also sells it. The name of the font that you get from Adobe should start with ‘p’. This is because font resellers typically make modifications to the original design.

Notes on specific suppliers:

‘f’ For fonts that are distributed without any specific attribution to the creator, by individuals, or by small foundries. (Unfortunately, we don’t have enough characters to assign one to every font supplier in the world.)

‘r’ obsolete; specifies raw fonts, in the old distribution of Dvips. New fonts should never use ‘r’. (The right thing to do is specify the correct encoding, variant, or whatever the font’s characteristics actually are.)

‘z’ for fonts that just don’t fit well into the naming scheme. The ‘z’ should be followed by the real supplier letter.

Here is the table from the file ‘supplier.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation, directory name, and comment.

9 unknown
a autologi  Autologic
b bitstrea  Bitstream

c cg Compugraphic d dtc Digital Typeface Corporation

e apple  Apple

f public small foundries g gnu Free Software Foundation h bh Bigelow & Holmes i itc International Typeface Corporation

k softkey  SoftKey
l linotype  Linotype
m monotype  Monotype
n ibm  IBM
o corel  Corel

p adobe Adobe (‘p’ for PostScript) r - raw (for [obsolete] use with virtual fonts)

s sun  Sun
t paragrap  ParaGraph
u urw  URW

z - bizarre (fontname is nonstandard)

- jknappen Joerg Knappen

2.2 Typefaces

For our purposes, a typeface is a collection of related fonts. We sometimes use the same typeface abbreviation for fonts with different supplier names, where

we know (usually by inspection) that the fonts truly are the same. This helps conserve abbreviations.

Notes on specific typefaces:

‘ad’

‘gm’ To explain a common confusion: ‘ad’ is Adobe Garamond (with a ‘FontName’ of ‘AGaramond’). This is quite different from the ‘Garamond’ typeface distributed by Adobe, which is actually ITC Garamond. (Garamond No. 3 (‘f3’), Simoncini Garamond (‘ig’), and Stempel Garamond (‘eg’) are still other variations on the Garamond theme.)

‘lh’
‘ls’
‘lx’

‘r1’ These abbreviations for Lucida typeface subfamilies, et al., solve a problem with names getting too long, or help conserve variant abbreviations. See Section 2.4 [Variants], page 14. Ideally, they would not exist.

‘ns’

‘nt’

‘tm’ Suppliers like to offer as many variations on the Times New Roman design as they do Garamond. Adobe has ‘Times’, ‘TimesNewRoman’, and ‘TimesNewRomanPS’. Although this last has been used by other foundries (e.g., Monotype) for a Times variation exactly compatible with Adobe’s ‘Times’, Adobe itself also offers it, so we must have abbreviations for all three.

‘pi’ For pi fonts. Since names of pi fonts are rather arbitrary, it is not useful to try to assign typeface abbreviations to all of them. Instead, we use the supplier letter, ‘pi’, an arbitarily-assigned three-digit number, a weight and (if necessary) a variant. For example, ‘mpi001r’. Pi font ‘001’ for one supplier has nothing to do with pi font ‘001’ for another.

Here is the table, from the file ‘typeface.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and directory name, possibly followed by PostScript ‘FontName’s and/or a comment in roman. The ‘FontName’ may be preceded by a letter s and a ‘:’, as in ‘b:Dutch’. In that case, the s is a supplier abbreviation, and we’ve used that abbreviation for the given ‘FontName’ from that supplier.

a1 arial        Arial
a2 amasis       Amasis
a3 adlib        AdLib
a4 altgoth2     AlternateGothicNo2
a5 allegro      Allegro
a6 amelia       Amelia
a7 amertext     AmericanText
a8 aurora       Aurora
a9 adsans       Adsans b:Humanist970BT
aa aachen       Aachen
ab aboeckli     ArnoldBoecklin
ac acaslon      ACaslon  Adobe Caslon
ad agaramon     AGaramond  Adobe Garamond
ag avantgar     AvantGarde m:MAvantGardeGothicPS u:URWGothicL
ah ashleysc     AshleyScript
ai abadi        Abadi
al albertus     Albertus b:Flareserif821BT
am amercana     Americana b:Americana
an anna         Anna
ao atqolive     AntiqueOlive b:Incised901BT
ap apollo       Apollo
aq antiqua      Antiqua u:URWAntiqua m:BookAntiqua
ar arcadia      Arcadia
as aldus        Aldus
at atypewri     AmericanTypewriter
au auriol       Auriol b:Freeform721
av avenir       Avenir
az azdzgrot     AkzidenzGrotesk b:Gothic725BT
b0 blackoak     Blackoak

b1 blado        Blado
b2 binnyos      BinnyOldStyle
b3 braggado     Braggadocio
b4 bison        Bison b:Brush738BT
b5 binghams     BinghamScript b:Freehand591BT
b6 block        Block b:Gothic821
b7 bernhard     Bernhard BernhardModern
b8 blippo       Blippo
b9 bremen       Bremen
bb bembo        Bembo b:Aldine401BT
bc banco        Banco
bd bodoni       Bodoni
be bell         Bell
bf biffo        Biffo
bg benguiat     Benguiat
bh bauhaus      Bauhaus
bi birch        Birch
bk bookman      Bookman m:MBookmanPS u:URWBookmanL
bl belwe        Belwe
bm bulmer       Bulmer
bn bernard      Bernard
bo balloon      Balloon
bp bodocamp     BodoniCampanile b:Modern735
br berling      Berling b:Revival565BT
bs bakersig     BakerSignet
bt boton        Boton
bu brush        Brush
bv baskervi     Baskerville
bw broadway     Broadway
by berkeley     Berkeley
c0 carolina     Carolina
c1 calisto      Calisto
c2 casln224     CaslonTwoTwentyFour
c3 caslon3      CaslonThree
c4 chicago      Chicago
c5 casln540     CaslonFiveForty
c6 clarion      Clarion
c7 compacta     Compacta
c8 calvert      Calvert
c9 codex        Codex b:Calligraphic421BT
ca caslon       Caslon
cb cooper       Cooper
cc cascade      Cascade b:Freehand471BT
cd clarendo     Clarendon
ce centenni     Centennial
cf clearfac     Clearface
cg charlema     Charlemagne
ch charter      Charter

ci candida      Candida
cl cloister     Cloister

cm cm Computer Modern

cn century      Century
co cochin       Cochin
cp copperpl     Copperplate
cq cheq         Cheq
cr courier      Courier m:MCourierPS m:CourierTwelveMT u:NumbusMonL
cs cntursbk     CenturySchoolbook
ct cheltenh     Cheltenham
cu centuros     CenturyOldStyle
cv clairvx      Clairvaux
cw cottonwd     Cottonwood
cx caxton       Caxton
cy city         City b:SquareSlabserif711BT
cz choc         Choc b:Staccato555BT
d0 davida       DavidaBold
dc domcasl      Dom DomCasual DomDiagonal
dd ducberry     DucDeBerry
de dellarob     DellaRobbia
di diotima      Diotima b:Calligraphic810BT
dl dolores      Dolores
dm delima       Delima
dr doric        Doric
ds dorchesc     DorchesterScript
dt dante        Dante
e0 embassy      Embassy
e1 empire       Empire
ea lectra       Electra b:Transitional521BT
eb baskerbe     BaskervilleBE
ec applecha     e:AppleChancery
ee egyptian     Egyptian EgyptianExtended
ef egyptien     EgyptienneF
eg stgaramo     StempelGaramond b:OriginalGaramondBT
eh engsft       Engschrift
ei esprit       Esprit
ek eckmann      Eckmann b:Freeform710
el ellingtn     Ellington
en engraver     Engravers
eo amerigo      Amerigo b:AmerigoBT
er eras         Eras
es eurostil     Eurostile b:Square721BT k:Minima
et ehrhardt     Ehrhardt
eu euler        Euler
ex exclsior     Excelsior
f0 flemishs     FlemishScript
f1 fournier     Fournier
f2 facade       Facade

f3 festival     Festival
f4 footligh     Footlight
f5 french       French
fa formata      Formata
fe forte        Forte
ff fairfld      Fairfield b:Transitional551BT
fg frankgth     FranklinGothic
fi figaro       Figaro
fl flora        Flora
fn fenice       Fenice
fo folio        Folio
fq frizquad     FrizQuadrata
fr frutiger     Frutiger b:Humanist777BT
fs falstaff     Falstaff
ft fettefrk     FetteFraktur Fraktur
fu futura       Futura
fx felixttl     FelixTitling
fy flyer        Flyer
g0 gando        Gando
g1 gothic13     GothicThirteen
g2 giddyup      Giddyup
g3 garamon3     GaramondThree b:AmercianGaramondBT  Garamond No. 3
g4 gymodern     GoudyModern
g5 gycatalg     GoudyCatalogue
g6 gorilla      GorillaITC
g9 ugaramon     GaramondRetrospective
gb gothicbb     GothicBBB
ge geneva       e:Geneva
gg garthgra     GarthGraphic
gj granjon      Granjon b:ElegantGaramondBT
gl galliard     Galliard
gm garamond     Garamond
go gyoldsty     GoudyOldStyle
gp glypha       Glypha
gq grotesq      Grotesque u:URWGrotesk
gr gloucest     Gloucester
gs gillsans     GillSans m:GillAltOneMT b:Humanist521BT
gt goudytxt     GoudyText
gu guardi       Guardi
gv giovanni     Giovanni
gy goudy        Goudy
h0 handelgo     HandelGothic
hb hobo         Hobo Tramp
hc herculnm     Herculanum
hd headline     Headline
hg hiroshig     Hiroshige
hi hvinsera     HelveticaInserat b:Swiss921BT
hm charme       Charme

hn hvneue       HelveticaNeue
ho horleyos     HorleyOldStyle
hr hvround      HelveticaRounded
hs hnseati      Hanseatic B:Swiss924BT
ht hoeflert     e:HoeflerText
hv helvetic     Helvetica u:NimbusSans u:NumbusSanL b:Swiss721BT b:Monospace821BT b:Swiss911BT d:Sans Newton n:SonoranSansSerif
hw schadow      Schadow
hx huxleyvr     HuxleyVertical
i0 commscr      CommercialScript
i1 impress      Impress
i2 impuls       Impuls
ia insignia     Insignia
ib ibycus       Ibycus
ic impact       Impact
if inflex       Inflex
ig scgaramo     SimonciniGaramond b:ItalianGaramond
ii imprint      Imprint
il caecilia     Caecilia
im imago        Imago
in industri     Industria
io ionic        Ionic b:News701BT
ip impressm     Impressum
is italnos      ItalianOldStyle
it italia       Italia
iw ironwood     Ironwood
iy iowanold     IowanOldStyle
jf jeffersn     Jefferson b:Freehand575BT
jn janson       Janson
jo joanna       Joanna
jp juniper      Juniper
kb kabel        Kabel
kg bankgoth     BankGothic
kl klang        Klang
kn kino         Kino
ko bookmano     BookmanOldStyle
kr korinna      Korinna
ks kuenstsc     KuenstlerScript
ku kaufmann     Kaufmann
l1 liberty      Liberty
l2 libra        Libra
l3 lucia        Lucia
l4 lucian       Lucian
lb lubaling     LubalinGraph
lc lucida       Lucida
lf life         Life
lg lettrgth     LetterGothic ISILGothic
lh lubright     LucidaBright
li latin        Latin

ln lino         Lino
lo lithos       Lithos
ls lucsans      LucidaSans
lt lutetia      Lutetia
lw leawood      Leawood
lx lucidfax     LucidaFax
ly lydian       Lydian
m0 monospac     Monospace
m1 monolisc     MonolineScript
m2 metro        Metro b:Geometric415BT
m3 mirarae      Mirareae
m4 mrearl       MisterEarl
m5 murrayhl     MurrayHill
m6 mittlsft     Mittelschrift
m7 matura       Matura
m8 monotyos     MonotypeOldStyle
ma mandate      Mandate b:Freehand521BT
mc monaco       e:Monaco
md meridien     Meridien
me mercursc     MercuriusScript
mg marigold     Marigold
mh machine      Machine
mi medici       Medici
ml melior       Melior b:ZapfElliptical711BT
mn minion       Minion
mo modern       Modern b:Modern20BT
mp memphis      Memphis b:GeometricSlab703BT
mq mesquite     Mesquite
mr madrone      Madrone
ms mistral      Mistral b:Staccato222BT
mt minister     Minister
mv malvern      Malvern
mx maximus      Maximus
mz amazone      Amazone
n0 normande     Normande
n1 binner       Binner
na newaster     NewAster
nb nbaskerv     NewBaskerville
nc ncntrsbk     NewCenturySchlbk u:CenturySchL
nd ncaledon     NewCaledonia
nf nofret       Nofret
ng newsgth      NewsGothic
ni nimrod       Nimrod
nk neograph     Neographik
nl neuland      Neuland b:Informal011BT
nm nimbus       Nimbus
nn nclarend     NewClarendon
np nsplanti     NewsPlantin

nr nberolin     NewBerolina
ns timesnps     TimesNewRomanPS
nt timesnew     TimesNewRoman m:TimesNR n:SonoranSerif
nu nuptial      Nuptial
nv novarese     Novarese
nz neuzeits     NeuzeitS b:Geometric706BT
o0 oldreadf     OldDreadfulNo7
o1 oranda       Oranda
o2 orbitb       OrbitB
o3 colonna      Colonna
o4 oceansan     OceanSans
o7 oldsty7      OldStyleSeven
oa ocra         OCR-A
ob ocrb         OCR-B
oc concorde     Concorde
oe oldengli     OldEnglish b:English157
oi ondine       Ondine Mermaid b:FormalScript421BT
ol colmcill     Colmcille
om omnia        Omnia
on corona       Corona
op optima       Optima d:Optimum b:ZapfHumanist601BT Omega
or orator       Orator
os oldstyle     OldStyle
ot coronet      Coronet b:Ribbon131BT
ov octavian     Octavian
ox onyx         Onyx
oy olympian     Olympian l:OlympicLH
oz ozhandcr     OzHandicraft
p0 poplar       Poplar
p1 palacesc     PalaceScript
p2 pepita       Pepita
p3 palette      Palette b:Brush445BT
p4 pioneer      PioneerITC
p5 piranesi     Piranesi
p6 blueprin     Blueprint
pa parkave      ParkAvenue
pb ptbarnum     PtBarnum
pc poetica      PoeticaChancery s:Scriptoria
pd placard      Placard
pe prestige     PrestigeElite
pf profile      Profile Profil b:Decorated035BT
pg peignot      Peignot b:Exotic350BT
ph photina      Photina

pi pi pi fonts—see text and ‘*.map’ files

pl palatino     Palatino d:Palton b:ZapfCalligraphic801BT u:URWPalladioL Palmspring k:Palindrome
pn plantin      Plantin b:Aldine721BT
po ponderos     Ponderosa
pp perpetua     Perpetua b:Lapidary333BT

pq postatqa     PostAntiqua
ps parisian     Parisian
pt present      Present
pu popplaud     PopplLaudatio
pw perrywoo     Perrywood
px popplpfx     PopplPontifex
py playbill     Playbill
pz poliphil     Poliphilus
qr quorum       Quorum
qt boulevar     Boulevard
r0 romana       Romana
r1 rotis        RotisSemi
r2 cantoria     Cantoria
ra raleigh      Raleigh
ro rotis        Rotis
rm carmina      Carmina
rp reporter     Reporter
rq russelsq     RussellSquare
rs arrus        Arrus
rt carta        Carta
ru runic        Runic
rv revue        Revue
rw rockwell     Rockwell Slate b:GeometricSlab712BT
ry ryumin       Ryumin
s0 sanmarco     SanMarco
s1 scotchro     ScotchRoman
s2 script       Script
s3 seagull      Seagull
s4 shotgun      Shotgun ShotgunBlanks
s5 snell        Snell
s6 stuyvesa     Stuyvesant
s7 sassoon      Sassoon
sb sabon        Sabon b:ClassicalGaramondBT
sc slimbach     Slimbach
sf serifa       Serifa
sg serifgth     SerifGothic
sh shannon      Shannon
sk skia         e:Skia
sl stencil      Stencil
sm spectrum     Spectrum
sn spartan      Spartan
so sonata       Sonata
sp serpenti     Serpentine
sr snellrnd     SnellRoundhand
ss stschnei     StempelSchneidler
st stone        Stone
sv souvenir     Souvenir
sw swing        Swing

sx syntax       Syntax
sy symbol       Symbol m:MSymbolPS u:StandardSymL
t0 tango        Tango
t1 thunderb     Thunderbird
t2 typo         TypoUpright
tb bebodoni     BBodoni  Berthold Bodoni
te timeseur     TimesEuropa
tf tiffany      Tiffany
tg tradegth     TradeGothic
ti techphon     TechPhonetic
tj trajan       Trajan
tk tekton       Tekton
tl castella     Castellar
tm times        Times u:NimbusRoman u:NimbusRomNo9L b:Dutch801BT
to torino       Torino b:Industrial736BT
tp tempo        Tempo
tv trumpmed     TrumpMediaeval b:Kuenstler480BT
tw twentyc      TwentiethCentury
ty typewrit     Typewriter
u0 columbus     Columbus
ub babodoni     BauerBodoni
ug bengugth     BenguiatGothic
uh cushing      Cushing
ul universa     Universal
um umbra        Umbra
un univers      Univers b:ZurichBT
ur centaur      Centaur b:Venetian301BT
ut utopia       Utopia
uw usherwoo     Usherwood
uy universi     University
v0 vineta       Vineta
va activa       Activa
vc vectora      Vectora
vd vandijck     VanDijck
vi victoria     Victoria
vj veljovic     Veljovic
vl versaill     Versailles
vr vagrnd       VAGRounded
w0 weddingt     WeddingT
wb walbaum      Walbaum
wd weideman     Weidemann
wg wlklgth      WilhelmKlingsporGotisch
wi wingding     Wingdings
wk wilke        Wilke
wn windsor      Windsor
wr wittenbe     WitternbergerFrakt
ws weiss        Weiss
ww willow       Willow

yg centuryg     CenturyGothic
yh strayhor     Strayhorn
ym stymie       Stymie
za zantiqua     ZAntiqua
zc zapfchan     ZapfChancery m:ZapfChanceryPS u:URWChanceryL
zd zapfding     ZapfDingbats m:ZapfDingbatsPS u:Dingbats
zg neuzeitg     NeuzeitGrotesk
zt zeitgeis     Zeitgeist

2.3 Weights

The weight of a font specifies its boldness. Arranged from lightest to heaviest (more or less): hairline, extra light, light book, regular, medium demibold, semibold, bold, extra bold, heavy black, ultra, poster

Here is the table, from the file ‘weight.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and any parts of a PostScript ‘FontName’ which use that abbreviation.

a Thin Hairline
b Bold
c Black
d Demi
h Heavy Heavyface
j ExtraLight
k Book
l Light
m Medium
p Poster
r Regular Roman
s Semibold
u Ultra UltraBlack
x ExtraBold ExtraBlack

2.4 Variants

Variants include typeface variations (e.g., italic, sans serif), and font encodings (e.g., Adobe standard, TEX text).

A fontname may require multiple variants. To resolve the worst ambiguities, we specify that any encoding variant (‘7’, ‘8’, or ‘9’, see below) come last and any other numeral variant come first (to avoid confusion with a design size). We recommend but do not require that the other variants be given in alphabetical order. (It’s not required because it’s too painful to implement the sorting in TEX, and many existing names already have non-alphabetized variants.)

The letterspacing possibilities introduced by fontinst and virtual fonts have not yet become sufficiently widespread to make standardization beneficial. Likewise for the many possible ways to generate small caps fonts.

Notes on specific variants, both old and new:

‘1’ ‘4’ ‘6’ ‘b’ These variants (‘semisans’, ‘bright’, ‘semiserif’, and ‘fax’) were used only for single

typefaces, so to conserve variant abbreviations, they now have typeface abbreviations instead. Also, for ‘b’, to avoid too-long names; and for the others, to avoid variant vs. design-size ambiguities.

‘g’ ‘h’ Since these variants were not actually used in any font names, they are now obsolete

and may be reassigned in the future.

‘5’ ‘6’ ‘7’ ‘8’ ‘9’ These characters begin a two-character encoding variant. Generally, ‘7’ is for 7-bit

encodings, ‘8’ is for 8-bit encodings, and ‘9’ is for expertised encodings with or without oldstyle digits (see the ‘x’ item below), but this is not an absolute rule. Also, ‘5’ is presently used for phonetic encodings and ‘6’ for Cyrillic encodings. The ‘9s’ (‘SuperFont’) variant contain all of Adobe, Latin 1, and Expert glyphs, perhaps with slightly different metrics than the original fonts. For a font to be named with a certain encoding variant, it’s not necessary that all the characters appear in precisely the same positions as in the encoding definition. It’s enough that the usual TEX macros work. In practice, this means that it’s ok for a font to be labelled ‘7t’ if the only difference from Computer Modern is that the ligatures and the ‘lslash’ and ‘Lslash’ characters are different, since it’s impossible to access or change the ligature table from TEX. Standard PostScript fonts don’t have an ‘lslash’ character the way Computer Modern does, but they do have the ‘Lslash’ and ‘lslash’ characters themselves, so the usual TEX \lslash and \Lslash macros can be made to work via ligatures. See the file ‘T1.etx’ file in the fontinst distribution for details. If a name does not contain a specific encoding variant, its encoding is unspecified. For example, some of the fonts distributed with Dvips(k) have names like ‘ptmr’ for ‘Times-Roman’; they use the Dvips encoding (see Section B.4 [dvips], page 92), which is close to (but not the same as) the TEX text encoding (as in Computer Modern Roman). Similarly, the TFM files distributed with Dvilj(k) for the builtin LaserJet 4 fonts have names like ‘cunm’, since these fonts, while compatible with TEX text, contain many additional characters.

‘9’ These items are needed only because ‘x’ (and possibly ‘j’) followed by a two-character regular variant makes some names too long. That is, ‘9t’ is equivalent in meaning to ‘x7t’, and ‘9d’ is equivalent to ‘jx7t’. (This may not be true for all ‘9x’ encodings, though.)

‘k’ ‘z’ These variants (indicating Greek and Cyrillic fonts, respectively) are obsolete; future

fonts in different scripts should be assigned an appropriate encoding abbreviation. These are not currently standardized, due to ignorance of the appropriate encodings.

Please send suggestions for abbreviations to the address in Chapter 1 [Introduction], page 1.

‘m’ ‘v’ ‘y’ These math-related variants remain for the sake of typeface-specific math encodings,

e.g., Lucida. Fonts that use the Computer Modern math encodings should use ‘7m’ (see Section B.7 [texmital], page 101), ‘7v’ (see Section B.5 [texmext], page 93), and ‘7y’ (see Section B.6 [texmsym], page 97).

‘n’ This is used for several different (but very rare) variants: only the Stone typeface has an “informal” variant, and only a few Monotype fonts have a “schoolbook” variant, with different ‘a’, ‘g’, and ‘y’ shapes.

‘r’ Included only if no other variants, including encodings, apply, and either the width (see Section 2.5 [Widths], page 18) is not ‘r’ or a design size is present. I.e., ‘r’ is only used as a placeholder. When the normal version of the typeface is sans serif (e.g., Helvetica), ‘r’ should be used when necessary, not ‘s’. Use ‘s’ only when the typeface family has both serif and sans serif variants.

‘s’ ‘t’ Mittelbach in TUGboat 13(1) suggests that these variants (for ‘sans’ and ‘typewriter’)

should be identified as part of the typeface name, because there are few typeface families with these variants. I feel that since they are logically variants, it’s best to name them that way. But ‘LucidaSans’ (see Section 2.2 [Typefaces], page 4) and a few others are exceptions, to avoid too-long names.

‘x’

‘8x’

‘9?’ ‘8x’ indicates a font in the ‘Expert’ encoding itself. ‘x’ indicates an expertised font, i.e., a composite (virtual) font that includes characters from an ‘8x’ font. And in fact an ‘xee’ sequence is replaced by ‘9e’, to save characters.

Fontname 1 assignments are shown in brackets in the following table, from the file ‘variant.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and either any parts of a PostScript ‘FontName’ which use that abbreviation or the PostScript ‘Encoding’ name.

obsolete [0=>8a] Adobe standard encoding obsolete [1 SemiSans] obsolete [2=>7c] Fraktur encoding obsolete [3=>7f] Fraction obsolete [4 fax; now typeface ‘lx’, Lucida Fax] 5 escape for (presently) phonetic encodings 6 escape for Cyrillic encodings [was ‘SemiSerif’] 7 escape for (mostly) 7-bit encodings 8 escape for (mostly) 8-bit encodings 9 escape for (presently) expert encodings [was oldstyle digits] a Alt Arrows Alternative [was alternate encoding] obsolete [b bright; now typeface lh, Lucida Bright]

c SmallCaps
d Display Titling Caption Headline TallCaps SwashCaps LombardicCaps
e Engraved Copperplate Elite

f Fraktur Gothic OldEnglish Handtooled (‘gothic’ can also be sans)

g SmallText designed for small sizes [was grooved, as in the IBM logo]

h Shadow
i Italic Kursiv Ital  text italic

j old-style digits [was invisible] k Greek obsolete

l Outline OpenFace Blanks

m math italic

n Informal Fashion Schlbk  for Stone
o Oblique Obl  slanted
p Ornaments

obsolete [q=>8t Cork (TEX extended) encoding] r roman or sans; often omitted, see text s Gothic sans serif t Monospace fixed-width typewriter u unslanted italic

v MathExtension
w Script Handwritten Swash Calligraphy Cursive Tango

x built with Adobe expert encoding [was expert-encoded]

y MathSymbol

z Cyrillic font-dependent Cyrillic

5a PhoneticAlternate
5i PhoneticIPA

5s sil-IPA 6b Cyrillic part of ISO 8859-5, seven bits 6d Cyrillic CP866 encoding 6i ISO 8859-5 6k Cyrillic KOI-8 encoding 6m Cyrillic Macintosh encoding 6w Cyrillic CP1251 encoding 7a A alternate characters only 7c Dfr Fraktur 7d OsF OSF oldstyle digit encoding

7f Fraction
7k OT2Cyrillic

7m TeXMathItalicEncoding see Section B.7 [texmital], page 101 7t TEX text encoding (as in Computer Modern Roman) 7v TeXMathExtensionEncoding see Section B.5 [texmext], page 93 7y TeXMathSymbolEncoding see Section B.6 [texmsym], page 97

82 GreekKeys
83 Ibycus1
84 Ibycus2

8a StandardEncoding Adobe standard encoding, see Section B.1 [8a], page 82

8c TeXTextCompanion  LaTEX’s TS1

8m Macintosh standard encoding 8r TeXBase1Encoding see Section B.2 [8r], page 87 8t CorkEncoding see Section B.3 [cork], page 90

8u XT2Encoding  cmtt + Latin 2, see ‘xl2.enc’

8w Windows 3.1 ANSI encoding 8x Expert expert encoding

8z XL2Encoding  cmr + Latin 2, see ‘xl2.enc’

9c expert + Text companion 9d expert + oldstyle digits + Cork 9e expert + Cork 9o expert + oldstyle digits + TEX text 9s SF SuperFont 9t expert + TEX text

2.5 Widths

The width of a font specifies the compression or expansion of the font. Arranged from narrowest to widest (more or less):

ultra compressed, extra condensed, compressed, condensed, narrow regular extended, expanded, wide

Expansion or compression of fonts is sometimes done automatically (as by the PostScript scale operator), and sometimes done by humans. In the latter case, the human will also presumably choose a font name which includes ‘Extended’ or ‘Expanded’ or ‘Condensed’ or ‘Narrow’ or whatever according to their own preferences; the abbreviation can follow along. When creating a new synthetically expanded or compressed font for use with TEX, e.g., with Afm2tfm or fontinst, use ‘n’ and ‘e’.

Here is the table, from the file ‘width.map’. It is organized alphabetically by abbreviation. Each line consists of an abbreviation and any parts of a PostScript ‘FontName’ which use that abbreviation.

c Condensed Cond
e Expanded
n Narrow
o UltraCondensed
p Compressed Compact
q ExtraCompressed ExtraCondensed
r Normal Medium Regular (usually omitted)
t Thin
u UltraCompressed
v ExtraExpanded more than Expanded, less than Wide
w Wide
x Extended Elongated

3 Long names

Eight characters is not enough to unambiguously represent all fonts. To do that, we have to allow ourselves arbitrarily long names. Although we certainly cannot use such names for the files themselves, we could still use them in TEX documents given a simple mapping file, as discussed below.

3.1 A fontname mapping file

At the moment, most implementations of TEX look up a TFM file (as part of the \font command), by searching for a file with the name given by the user (possibly in any of series of directories). But if we also looked TFM names up in another file (or set of files), which specifies the actual filename, the fontname given in the TEX source file could be almost anything at all, of any length.

In version 5.851d of Web2c, I implemented this mapping file. Each file ‘texfonts.map’ in a search path is read for abbreviations. The file has a straightforward format: each line specifies the filename and the TEX name for one font, separated by whitespace. Extra information on the line is ignored; then more information could be specified for the benefit of DVI-reading programs in the same file. Comments start with % and continue to the end of the line.

Besides allowing long names, this sort of mapping file has other benefits. TEX source or DVI files can be more easily transported, because the font names in a particular file can be made work on every system. Also, when combined with a consistent naming scheme, macros could be written to access any of a number of fonts. Right now, each font family has to have specialized macros written to deal with it.

Incidentally, Professor Knuth has approved this as a legitimate “system-dependent” extension; a TEX with such a feature can still be called “TEX”.

3.2 A naming scheme for long names

Once we allow ourselves long names, we can construct a naming scheme to handle arbitrary fonts without much difficulty. Here is one proposal, based on the X Window System font naming conventions.

supplier-family-weight-variants-width-encoding--size supplier is the usual Adobe, Autologic, etc., as well as unknown or bizarre—this last meaning

the rest of the name is nonstandard. If the supplier is missing, i.e., the name starts with a -, “public domain” is assumed. For fonts made by individuals, the initials of the designer would probably make a good supplier.

family is ComputerModern or Times or whatever.

Everything else is optional. The -- before the size lets one specify a name with, say, a weight and variants, but then skip the width and encoding, but still be able to give a size.

weight and width are as described earlier.

If there is more than one variant, they are separated with some character other than ‘-’, say ‘=’:

     B+H-Lucida-Bold-Sans=Typewriter--10

encoding is what Metafont calls the font_coding_scheme—the layout of the characters in the

font. For example, ‘Cork’ or ‘ISOLatin1’ or ‘AdobeAlternate’.

Names are case-sensitive, for consistency with the rest of TEX and with PostScript, etc. Spaces cannot be used in the name, to make it easier for TEX to parse. Likewise, characters with default category codes other than 11 (letter) or 12 (other) should not be used.

Another possibility is to forget all the above, and simply use the vendor’s name (perhaps prefixed by the vendor): ‘Times-Roman’.

Appendix A Font name lists

These sections contain a list of fonts from various suppliers together with the corresponding names.

A.1 Standard PostScript fonts

Abbreviated names for the 35 standard PostScript fonts. An encoding variant, such as ‘8a’ for the Adobe ‘StandardEncoding’, is omitted.

‘pagk’  ‘AvantGarde-Book’
‘pagko’  ‘AvantGarde-BookOblique’
‘pagd’  ‘AvantGarde-Demi’
‘pagdo’  ‘AvantGarde-DemiOblique’
‘pbkd’  ‘Bookman-Demi’
‘pbkdi’  ‘Bookman-DemiItalic’
‘pbkl’  ‘Bookman-Light’
‘pbkli’  ‘Bookman-LightItalic’
‘pcrb’  ‘Courier-Bold’
‘pcrbo’  ‘Courier-BoldOblique’
‘pcrr’  ‘Courier’
‘pcrro’  ‘Courier-Oblique’
‘phvb’  ‘Helvetica-Bold’
‘phvbo’  ‘Helvetica-BoldOblique’
‘phvbrn’ ‘Helvetica-NarrowBold’
‘phvbon’ ‘Helvetica-NarrowBoldOblique’
‘phvr’  ‘Helvetica’
‘phvro’  ‘Helvetica-Oblique’
‘phvrrn’ ‘Helvetica-Narrow’
‘phvron’ ‘Helvetica-NarrowOblique’
‘pncb’  ‘NewCenturySchlbk-Bold’
‘pncbi’  ‘NewCenturySchlbk-BoldItalic’
‘pncri’  ‘NewCenturySchlbk-Italic’
‘pncr’  ‘NewCenturySchlbk-Roman’
‘pplb’  ‘Palatino-Bold’
‘pplbi’  ‘Palatino-BoldItalic’

‘pplri’  ‘Palatino-Italic’
‘pplr’  ‘Palatino-Roman’
‘psyr’  ‘Symbol’
‘ptmb’  ‘Times-Bold’
‘ptmbi’  ‘Times-BoldItalic’
‘ptmri’  ‘Times-Italic’
‘ptmr’  ‘Times-Roman’
‘pzcmi’  ‘ZapfChancery-MediumItalic’
‘pzdr’  ‘ZapfDingbats’

A.2 Adobe fonts

The ‘adobe.map’ file has abbreviations for Adobe fonts.

ierb8r   ItcEras-Bold
ierd8r   ItcEras-Demi
ierk8r   ItcEras-Book
ierl8r   ItcEras-Light
ierm8r   ItcEras-Medium
ieru8r   ItcEras-Ultra
ikbb8r   ItcKabel-Bold
ikbd8r   ItcKabel-Demi
ikbk8r   ItcKabel-Book
ikbm8r   ItcKabel-Medium
ikbu8r   ItcKabel-Ultra
isyb8r   ItcSymbol-Bold
isybi8r  ItcSymbol-BoldItalic
isyc8r   ItcSymbol-Black
isyci8r  ItcSymbol-BlackItalic
isyk8r   ItcSymbol-Book
isyki8r  ItcSymbol-BookItalic
isym8r   ItcSymbol-Medium
isymi8r  ItcSymbol-MediumItalic
paab8r   Aachen-Bold
pabr8r   ArnoldBoecklin
pacb7a   ACaslon-AltBold
pacb8r   ACaslon-Bold
pacb8x   ACaslonExp-Bold
pacbi7a  ACaslon-AltBoldItalic
pacbi8r  ACaslon-BoldItalic
pacbi8x  ACaslonExp-BoldItalic
pacbwi8r ACaslon-SwashBoldItalic
pacr7a   ACaslon-AltRegular
pacr8r   ACaslon-Regular
pacr8x   ACaslonExp-Regular

pacri7a  ACaslon-AltItalic
pacri8r  ACaslon-Italic
pacri8x  ACaslonExp-Italic
pacrp8r  ACaslon-Ornaments
pacrwi8r ACaslon-SwashItalic
pacs7a   ACaslon-AltSemibold
pacs8r   ACaslon-Semibold
pacs8x   ACaslonExp-Semibold
pacsi7a  ACaslon-AltSemiboldItalic
pacsi8r  ACaslon-SemiboldItalic
pacsi8x  ACaslonExp-SemiboldItalic
pacswi8r ACaslon-SwashSemiboldItalic
padb8r   AGaramond-Bold
padb8x   AGaramondExp-Bold
padbi8r  AGaramond-BoldItalic
padbi8x  AGaramondExp-BoldItalic
padr7a   AGaramondAlt-Regular
padr8r   AGaramond-Regular
padr8x   AGaramondExp-Regular
padrd8r  AGaramond-Titling
padri7a  AGaramondAlt-Italic
padri8r  AGaramond-Italic
padri8x  AGaramondExp-Italic
pads8r   AGaramond-Semibold
pads8x   AGaramondExp-Semibold
padsi8r  AGaramond-SemiboldItalic
padsi8x  AGaramondExp-SemiboldItalic
pagb8r   AvantGarde-Bold
pagbo8r  AvantGarde-BoldObl
pagbr8rc AvantGarde-CondBold
pagd8r   AvantGarde-Demi
pagdo8r  AvantGarde-DemiOblique
pagdr8rc AvantGarde-CondDemi
pagi8r   AvantGarde-ExtraLight
pagio8r  AvantGarde-ExtraLightObl
pagk8r   AvantGarde-Book
pagko8r  AvantGarde-BookOblique
pagkr8rc AvantGarde-CondBook
pagm8r   AvantGarde-Medium
pagmo8r  AvantGarde-MediumObl
pagmr8rc AvantGarde-CondMedium
pamb8r   Americana-Bold
pamr8r   Americana
pamri8r  Americana-Italic
pamx8r   Americana-ExtraBold
panr8r   Anna
paob7a   AntiqueOliveEastA-Bold
paob8r   AntiqueOlive-Bold

paobr8rc AntiqueOlive-BoldCond
paoc7a   AntiqueOliveEastA-Black
paoc8r   AntiqueOlive-Black
paohix8r AntiqueOlive-NordItalic
paohrx8r AntiqueOlive-Nord
paol7a   AntiqueOliveEastA-Light
paol8r   AntiqueOlive-Light
paor7a   AntiqueOliveEastA-Roman
paor8r   AntiqueOlive-Roman
paori7a  AntiqueOliveEastA-Italic
paori8r  AntiqueOlive-Italic
paorr8rn AntiqueOlive-Compact
parr7a   Arcadia-A
parr8r   Arcadia
pasr8r   Aldus-Roman
pasrc8r  Aldus-RomanSC
pasri7d  Aldus-ItalicOsF
pasri8r  Aldus-Italic
patb7aa  AmericanTypewriter-BoldA
patb7ac  AmericanTypewriter-BoldCondA
patb8r   AmericanTypewriter-Bold
patbr8rc AmericanTypewriter-BoldCond
patl7a   AmericanTypewriter-LightA
patl7ac  AmericanTypewriter-LightCondA
patl8r   AmericanTypewriter-Light
patlr8rc AmericanTypewriter-LightCond
patm7a   AmericanTypewriter-MediumA
patm8r   AmericanTypewriter-Medium
patr7ac  AmericanTypewriter-CondA
patrr8rc AmericanTypewriter-Cond
paub8r   Auriol-Bold
paubi8r  Auriol-BoldItalic
pauc8r   Auriol-Black
pauci8r  Auriol-BlackItalic
paur8r   Auriol
pauri8r  Auriol-Italic
pavc8r   Avenir-Black
pavco8r  Avenir-BlackOblique
pavh8r   Avenir-Heavy
pavho8r  Avenir-HeavyOblique
pavk8r   Avenir-Book
pavko8r  Avenir-BookOblique
pavl8r   Avenir-Light
pavlo8r  Avenir-LightOblique
pavm8r   Avenir-Medium
pavmo8r  Avenir-MediumOblique
pavr8r   Avenir-Roman
pavro8r  Avenir-Oblique

pazb8r   AkzidenzGrotesk-Bold
pazc8r   AkzidenzGrotesk-Black
pazl8r   AkzidenzGrotesk-Light
pazr8r   AkzidenzGrotesk-Roman
pb0r8r   Blackoak
pbbb8r   Bembo-Bold
pbbbi8r  Bembo-BoldItalic
pbbr8r   Bembo
pbbri8r  Bembo-Italic
pbbs8r   Bembo-Semibold
pbbsi8r  Bembo-SemiboldItalic
pbbx8r   Bembo-ExtraBold
pbbxi8r  Bembo-ExtraBoldItalic
pbcr8r   Banco
pbdb8r   Bodoni-Bold
pbdbi8r  Bodoni-BoldItalic
pbdbr8rc Bodoni-BoldCondensed
pbdk8r   Bodoni-Book
pbdki8r  Bodoni-BookItalic
pbdp8r   Bodoni-Poster
pbdpi8r  Bodoni-PosterItalic
pbdpr8rp Bodoni-PosterCompressed
pbdr8r   Bodoni
pbdri8r  Bodoni-Italic
pbgb8r   Benguiat-Bold
pbgk8r   Benguiat-Book
pbhb8r   Bauhaus-Bold
pbhd8r   Bauhaus-Demi
pbhh8r   Bauhaus-Heavy
pbhl8r   Bauhaus-Light
pbhm8r   Bauhaus-Medium
pbir8r   Birch
pbkb8r   Bookman-Bold
pbkbi8r  Bookman-BoldItalic
pbkd8r   Bookman-Demi
pbkdi8r  Bookman-DemiItalic
pbkl8r   Bookman-Light
pbkli8r  Bookman-LightItalic
pbkm8r   Bookman-Medium
pbkmi8r  Bookman-MediumItalic
pblb8r   Belwe-Bold
pbll8r   Belwe-Light
pblm8r   Belwe-Medium
pblrr8rc Belwe-Condensed
pbrb8r   Berling-Bold
pbrbi8r  Berling-BoldItalic
pbrr8r   Berling-Roman
pbrri8r  Berling-Italic

pbsr8r   BakerSignet
pburw8r  BrushScript
pbvbz8r  BaskervilleCyrillic-Bold
pbvroz8r BaskervilleCyrillic-Inclined
pbvrz8r  BaskervilleCyrillic-Upright
pbyb8r   Berkeley-Bold
pbybi8r  Berkeley-BoldItalic
pbyc8r   Berkeley-Black
pbyci8r  Berkeley-BlackItalic
pbyk8r   Berkeley-Book
pbyki8r  Berkeley-BookItalic
pbym8r   Berkeley-Medium
pbyri8r  Berkeley-Italic
pc0r7c   Carolina-Dfr
pc0r8r   Carolina
pc2b8r   CaslonTwoTwentyFour-Bold
pc2bi8r  CaslonTwoTwentyFour-BoldIt
pc2c8r   CaslonTwoTwentyFour-Black
pc2ci8r  CaslonTwoTwentyFour-BlackIt
pc2k8r   CaslonTwoTwentyFour-Book
pc2ki8r  CaslonTwoTwentyFour-BookIt
pc2m8r   CaslonTwoTwentyFour-Medium
pc2mi8r  CaslonTwoTwentyFour-MediumIt
pc3r8r   CaslonThree-Roman
pc3rc8r  CaslonThree-RomanSC
pc3ri7d  CaslonThree-ItalicOsF
pc3ri8r  CaslonThree-Italic
pc5r8r   CaslonFiveForty-Roman
pc5rc8r  CaslonFiveForty-RomanSC
pc5ri7d  CaslonFiveForty-ItalicOsF
pc5ri8r  CaslonFiveForty-Italic
pcarl8r  CaslonOpenFace
pcbc8r   CooperBlack
pcbci8r  CooperBlack-Italic
pccrw8r  CascadeScript
pcdb8r   Clarendon-Bold
pcdl8r   Clarendon-Light
pcdr8r   Clarendon
pceb7d   Centennial-BoldOsF
pceb8r   Centennial-Bold
pcebi7d  Centennial-BoldItalicOsF
pcebi8r  Centennial-BoldItalic
pcec7d   Centennial-BlackOsF
pcec8r   Centennial-Black
pceci7d  Centennial-BlackItalicOsF
pceci8r  Centennial-BlackItalic
pcel8r   Centennial-Light
pcelc8r  Centennial-LightSC

pceli7d  Centennial-LightItalicOsF
pceli8r  Centennial-LightItalic
pcer8r   Centennial-Roman
pcerc8r  Centennial-RomanSC
pceri7d  Centennial-ItalicOsF
pceri8r  Centennial-Italic
pcfb8r   Clearface-Bold
pcfbi8r  Clearface-BoldItalic
pcfc8r   Clearface-Black
pcfci8r  Clearface-BlackItalic
pcfh8r   Clearface-Heavy
pcfhi8r  Clearface-HeavyItalic
pcfr8r   Clearface-Regular
pcfri8r  Clearface-RegularItalic
pcgb8r   Charlemagne-Bold
pcgr8r   Charlemagne-Regular
pcib8r   Candida-Bold
pcir8r   Candida-Roman
pciri8r  Candida-Italic
pcnb8r   Century-Bold
pcnbi8r  Century-BoldItalic
pcnbi8rc Century-BoldCondensedItalic
pcnbr8rc Century-BoldCondensed
pcnk8r   Century-Book
pcnki8r  Century-BookItalic
pcnki8rc Century-BookCondensedItalic
pcnkr8rc Century-BookCondensed
pcnl8r   Century-Light
pcnli8r  Century-LightItalic
pcnli8rc Century-LightCondensedItalic
pcnlr8rc Century-LightCondensed
pcnri8rx CenturyExpanded-Italic
pcnrr8rx CenturyExpanded
pcnu8r   Century-Ultra
pcnui8r  Century-UltraItalic
pcnui8rc Century-UltraCondensedItalic
pcnur8rc Century-UltraCondensed
pcob8r   Cochin-Bold
pcobi8r  Cochin-BoldItalic
pcor8r   Cochin
pcori8r  Cochin-Italic
pcqr8r   Cheq
pcrb8r   Courier-Bold
pcrbo8r  Courier-BoldOblique
pcrr8r   Courier
pcrro8r  Courier-Oblique
pctb8r   Cheltenham-Bold
pctbi8r  Cheltenham-BoldItalic

pctbi8rc Cheltenham-BoldCondItalic
pctbr8rc Cheltenham-BoldCond
pctk8r   Cheltenham-Book
pctki8r  Cheltenham-BookItalic
pctki8rc Cheltenham-BookCondItalic
pctkr8rc Cheltenham-BookCond
pctl8r   Cheltenham-Light
pctli8r  Cheltenham-LightItalic
pctli8rc Cheltenham-LightCondItalic
pctlr8rc Cheltenham-LightCond
pctrio8r Cheltenham-UltraCondItalic
pctrro8r Cheltenham-UltraCond
pctu8r   Cheltenham-Ultra
pctui8r  Cheltenham-UltraItalic
pcub8r   CenturyOldStyle-Bold
pcur8r   CenturyOldStyle-Regular
pcuri8r  CenturyOldStyle-Italic
pcvr7c   Clairvaux-Dfr
pcvr8r   Clairvaux
pcwr8r   Cottonwood
pcxb8r   Caxton-Bold
pcxbi8r  Caxton-BoldItalic
pcxk8r   Caxton-Book
pcxki8r  Caxton-BookItalic
pcxl8r   Caxton-Light
pcxli8r  Caxton-LightItalic
pcyb8r   City-Bold
pcybi8r  City-BoldItalic
pcyl8r   City-Light
pcyli8r  City-LightItalic
pcym8r   City-Medium
pcymi8r  City-MediumItalic
pdcb8r   DomCasual-Bold
pdcr8r   DomCasual
pddr7c   DucDeBerry-Dfr
pddr8r   DucDeBerry
pdrb8r   Doric-Bold
pefb8r   EgyptienneF-Bold
pefc8r   EgyptienneF-Black
pefr8r   EgyptienneF-Roman
pefri8r  EgyptienneF-Italic
pegb7d   StempelGaramond-BoldOsF
pegb8r   StempelGaramond-Bold
pegbi7d  StempelGaramond-BoldItalicOsF
pegbi8r  StempelGaramond-BoldItalic
pegr8r   StempelGaramond-Roman
pegrc8r  StempelGaramond-RomanSC
pegri7d  StempelGaramond-ItalicOsF

pegri8r  StempelGaramond-Italic
pehr8r   DINEngschrift
peib8r   Esprit-Bold
peibi8r  Esprit-BoldItalic
peic8r   Esprit-Black
peici8r  Esprit-BlackItalic
peik8r   Esprit-Book
peiki8r  Esprit-BookItalic
peim8r   Esprit-Medium
peimi8r  Esprit-MediumItalic
pesb8r   Eurostile-Bold
pesb8rx2 Eurostile-BoldExtendedTwo
pesbo8r  Eurostile-BoldOblique
pesbr8rc Eurostile-BoldCondensed
pesd8r   Eurostile-Demi
pesdo8r  Eurostile-DemiOblique
pesr28rx Eurostile-ExtendedTwo
pesr8r   Eurostile
pesr8rc2 Eurostile-Condensed
pesro8r  Eurostile-Oblique
pexb7a   ExcelsiorEastA-Bold
pexb8r   Excelsior-Bold
pexbz8r  ExcelsiorCyrillic-Bold
pexr7a   ExcelsiorEastA-Roman
pexr8r   Excelsior
pexri7a  ExcelsiorEastA-Italic
pexri8r  Excelsior-Italic
pexroz8r ExcelsiorCyrillic-Inclined
pexrz8r  ExcelsiorCyrillic-Upright
pfab8r   Formata-Bold
pfabi8r  Formata-BoldItalic
pfal8r   Formata-Light
pfali8r  Formata-LightItalic
pfam8r   Formata-Medium
pfami8r  Formata-MediumItalic
pfar8r   Formata-Regular
pfari8r  Formata-Italic
pfgd8r   FranklinGothic-Demi
pfgdo8r  FranklinGothic-DemiOblique
pfgh8r   FranklinGothic-Heavy
pfgho8r  FranklinGothic-HeavyOblique
pfgk8r   FranklinGothic-Book
pfgko8r  FranklinGothic-BookOblique
pfgr8r   FranklinGothic-Roman
pfgrr8rc FranklinGothic-Condensed
pfgrr8rq FranklinGothic-ExtraCond
pflb8r   Flora-Bold
pflm8r   Flora-Medium

pfnb8r   Fenice-Bold
pfnbo8r  Fenice-BoldOblique
pfnl8r   Fenice-Light
pfnlo8r  Fenice-LightOblique
pfnr8r   Fenice-Regular
pfnro8r  Fenice-RegularOblique
pfnu8r   Fenice-Ultra
pfnuo8r  Fenice-UltraOblique
pfob8r   Folio-Bold
pfobr8rc Folio-BoldCondensed
pfol8r   Folio-Light
pfom8r   Folio-Medium
pfox8r   Folio-ExtraBold
pfqb8r   FrizQuadrata-Bold
pfqr8r   FrizQuadrata
pfrb8r   Frutiger-Bold
pfrbi8r  Frutiger-BoldItalic
pfrbr8rc Frutiger-BoldCn
pfrc8r   Frutiger-Black
pfrci8r  Frutiger-BlackItalic
pfrcr8rc Frutiger-BlackCn
pfrl8r   Frutiger-Light
pfrli8r  Frutiger-LightItalic
pfrlr8rc Frutiger-LightCn
pfrr8r   Frutiger-Roman
pfrri8r  Frutiger-Italic
pfrrr8rc Frutiger-Cn
pfru8r   Frutiger-UltraBlack
pfrxr8rc Frutiger-ExtraBlackCn
pfsrw8r  FreestyleScript
pftrf7c  FetteFraktur-Dfr
pftrf8r  FetteFraktur
pfub8r   Futura-Bold
pfubo8r  Futura-BoldOblique
pfubo8rc Futura-CondensedBoldOblique
pfubr8rc Futura-CondensedBold
pfuh8r   Futura-Heavy
pfuho8r  Futura-HeavyOblique
pfuk8r   Futura-Book
pfuko8r  Futura-BookOblique
pful8r   Futura-Light
pfulo8r  Futura-LightOblique
pfulo8rc Futura-CondensedLightOblique
pfulr8rc Futura-CondensedLight
pfur8r   Futura
pfuro8r  Futura-Oblique
pfuro8rc Futura-CondensedOblique
pfurr8rc Futura-Condensed

pfux8r   Futura-ExtraBold
pfuxo8r  Futura-ExtraBoldOblique
pfuxo8rc Futura-CondExtraBoldObl
pfuxr8rc Futura-CondensedExtraBold
pfycr8rc Flyer-BlackCondensed
pfyxr8rc Flyer-ExtraBlackCondensed
pg1r8r   Gothic-Thirteen
pg3b8r   GaramondThree-Bold
pg3bc8r  GaramondThree-BoldSC
pg3bi7d  GaramondThree-BoldItalicOsF
pg3bi8r  GaramondThree-BoldItalic
pg3r8r   GaramondThree
pg3rc8r  GaramondThree-SC
pg3ri7d  GaramondThree-ItalicOsF
pg3ri8r  GaramondThree-Italic
pggb8r   GarthGraphic-Bold
pggbi8r  GarthGraphic-BoldItalic
pggbr8rc GarthGraphic-BoldCondensed
pggc8r   GarthGraphic-Black
pggr8r   GarthGraphic
pggri8r  GarthGraphic-Italic
pggrr8rc GarthGraphic-Condensed
pggx8r   GarthGraphic-ExtraBold
pgjb7d   Granjon-BoldOsF
pgjb8r   Granjon-Bold
pgjr8r   Granjon
pgjrc8r  Granjon-SC
pgjri7d  Granjon-ItalicOsF
pgjri8r  Granjon-Italic
pglb8r   Galliard-Bold
pglbi8r  Galliard-BoldItalic
pglc8r   Galliard-Black
pglci8r  Galliard-BlackItalic
pglr8r   Galliard-Roman
pglri8r  Galliard-Italic
pglu8r   Galliard-Ultra
pglui8r  Galliard-UltraItalic
pgmb8r   Garamond-Bold
pgmbi8r  Garamond-BoldItalic
pgmbi8rc Garamond-BoldCondensedItalic
pgmbr8rc Garamond-BoldCondensed
pgmk8r   Garamond-Book
pgmki8r  Garamond-BookItalic
pgmki8rc Garamond-BookCondensedItalic
pgmkr8rc Garamond-BookCondensed
pgml8r   Garamond-Light
pgmli8r  Garamond-LightItalic
pgmli8rc Garamond-LightCondensedItalic

pgmlr8rc Garamond-LightCondensed
pgmu8r   Garamond-Ultra
pgmui8r  Garamond-UltraItalic
pgmui8rc Garamond-UltraCondensedItalic
pgmur8rc Garamond-UltraCondensed
pgpa8r   Glypha-Thin
pgpao8r  Glypha-ThinOblique
pgpb8r   Glypha-Bold
pgpbo8r  Glypha-BoldOblique
pgpc8r   Glypha-Black
pgpco8r  Glypha-BlackOblique
pgpl8r   Glypha-Light
pgplo8r  Glypha-LightOblique
pgpr8r   Glypha
pgpro8r  Glypha-Oblique
pgsb8r   GillSans-Bold
pgsbi8r  GillSans-BoldItalic
pgsbr8rc GillSans-BoldCondensed
pgsl8r   GillSans-Light
pgsli8r  GillSans-LightItalic
pgsr8r   GillSans
pgsri8r  GillSans-Italic
pgsrr8rc GillSans-Condensed
pgsu8r   GillSans-UltraBold
pgsur8rc GillSans-UltraBoldCondensed
pgsx8r   GillSans-ExtraBold
pgub8r   Guardi-Bold
pgubi8r  Guardi-BoldItalic
pguc8r   Guardi-Black
pguci8r  Guardi-BlackItalic
pgur8r   Guardi-Roman
pguri8r  Guardi-Italic
pgvb8r   Giovanni-Bold
pgvbi8r  Giovanni-BoldItalic
pgvc8r   Giovanni-Black
pgvci8r  Giovanni-BlackItalic
pgvk8r   Giovanni-Book
pgvki8r  Giovanni-BookItalic
pgyb8r   Goudy-Bold
pgybi8r  Goudy-BoldItalic
pgyh8r   Goudy-Heavyface
pgyhi8r  Goudy-HeavyfaceItalic
pgyr8r   Goudy
pgyrc8r  Goudy-SC
pgyri8r  Goudy-Italic
pgyx8r   Goudy-ExtraBold
phbr8r   Hobo
phcr8r   Herculanum

phgb8r   Hiroshige-Bold
phgbi8r  Hiroshige-BoldItalic
phgc8r   Hiroshige-Black
phgci8r  Hiroshige-BlackItalic
phgk8r   Hiroshige-Book
phgki8r  Hiroshige-BookItalic
phgm8r   Hiroshige-Medium
phgmi8r  Hiroshige-MediumItalic
phir8r   HelveticaInserat-Roman
phmr8r   Charme
phna8r   HelveticaNeue-Thin
phnai8rt HelveticaNeue-ThinItalic
phnao8rc HelveticaNeue-ThinCondObl
phnao8rx HelveticaNeue-ThinExtObl
phnar8rc HelveticaNeue-ThinCond
phnar8rx HelveticaNeue-ThinExt
phnb7a   HelveticaNeueEastA-Bold
phnb8r   HelveticaNeue-Bold
phnbi8r  HelveticaNeue-BoldItalic
phnbl8r  HelveticaNeue-BoldOutline
phnbo7a  HelveticaNeueEastA-BoldObl
phnbo8rc HelveticaNeue-BoldCondObl
phnbo8rx HelveticaNeue-BoldExtObl
phnbr8rc HelveticaNeue-BoldCond
phnbr8rx HelveticaNeue-BoldExt
phnc8r   HelveticaNeue-Black
phnci8r  HelveticaNeue-BlackItalic
phnco8rc HelveticaNeue-BlackCondObl
phnco8rx HelveticaNeue-BlackExtObl
phncr8rc HelveticaNeue-BlackCond
phncr8rx HelveticaNeue-BlackExt
phnh8r   HelveticaNeue-Heavy
phnhi8r  HelveticaNeue-HeavyItalic
phnho8rc HelveticaNeue-HeavyCondObl
phnho8rx HelveticaNeue-HeavyExtObl
phnhr8rc HelveticaNeue-HeavyCond
phnhr8rx HelveticaNeue-HeavyExt
phnj8r   HelveticaNeue-UltraLight
phnji8r  HelveticaNeue-UltraLightItal
phnjo8rc HelveticaNeue-UltraLigCondObl
phnjo8rx HelveticaNeue-UltraLigExtObl
phnjr8rc HelveticaNeue-UltraLigCond
phnjr8rx HelveticaNeue-UltraLigExt
phnl8r   HelveticaNeue-Light
phnli8r  HelveticaNeue-LightItalic
phnlo8rc HelveticaNeue-LightCondObl
phnlo8rx HelveticaNeue-LightExtObl
phnlr8rc HelveticaNeue-LightCond

phnlr8rx HelveticaNeue-LightExt
phnm8r   HelveticaNeue-Medium
phnmi8r  HelveticaNeue-MediumItalic
phnmo8rc HelveticaNeue-MediumCondObl
phnmo8rx HelveticaNeue-MediumExtObl
phnmr8rc HelveticaNeue-MediumCond
phnmr8rx HelveticaNeue-MediumExt
phnr7a   HelveticaNeueEastA-Roman
phnr7at  HelveticaNeueEastA-Thin
phnr8r   HelveticaNeue-Roman
phnri8r  HelveticaNeue-Italic
phnro7a  HelveticaNeueEastA-RomanObl
phnro7at HelveticaNeueEastA-ThinObl
phnro8rc HelveticaNeue-CondensedObl
phnro8rx HelveticaNeue-ExtendedObl
phnrr8rc HelveticaNeue-Condensed
phnrr8rx HelveticaNeue-Extended
phnxo8rc HelveticaNeue-ExtBlackCondObl
phnxr8rc HelveticaNeue-ExtBlackCond
phrb8r   HelveticaRounded-Bold
phrbo8r  HelveticaRounded-BoldObl
phrbo8rc HelveticaRounded-BoldCondObl
phrbr8rc HelveticaRounded-BoldCond
phrc8r   HelveticaRounded-Black
phrco8r  HelveticaRounded-BlackObl
phvb7f   Helvetica-FractionBold
phvb8r   Helvetica-Bold
phvbk8r  HelveticaGreekP-Bold
phvbko8r HelveticaGreekP-BoldInclined
phvbo8r  Helvetica-BoldOblique
phvbo8rc Helvetica-Condensed-BoldObl
phvbo8rn Helvetica-Narrow-BoldOblique
phvboz8r HelveticaCyrillic-BoldIncl
phvbr8rc Helvetica-Condensed-Bold
phvbr8rn Helvetica-Narrow-Bold
phvbr8rp Helvetica-Compressed
phvbz8r  HelveticaCyrillic-Bold
phvc8r   Helvetica-Black
phvco8r  Helvetica-BlackOblique
phvco8rc Helvetica-Condensed-BlackObl
phvcr8rc Helvetica-Condensed-Black
phvl8r   Helvetica-Light
phvlo8r  Helvetica-LightOblique
phvlo8rc Helvetica-Condensed-LightObl
phvlr8rc Helvetica-Condensed-Light
phvr7f   Helvetica-Fraction
phvr8r   Helvetica
phvrk8r  HelveticaGreekP-Upright

phvrko8r HelveticaGreekP-Inclined
phvro8r  Helvetica-Oblique
phvro8rc Helvetica-Condensed-Oblique
phvro8rn Helvetica-Narrow-Oblique
phvroz8r HelveticaCyrillic-Inclined
phvrr8rc Helvetica-Condensed
phvrr8rn Helvetica-Narrow
phvrr8rq Helvetica-ExtraCompressed
phvrr8ru Helvetica-UltraCompressed
phvrz8r  HelveticaCyrillic-Upright
piar7a   Insignia-A
piar8r   Insignia
picr8r   Impact
pigb8r   SimonciniGaramond-Bold
pigr8r   SimonciniGaramond
pigri8r  SimonciniGaramond-Italic
pilb8r   Caecilia-Bold
pilbc8r  Caecilia-BoldSC
pilbci8r Caecilia-BoldItalicSC
pilbi8r  Caecilia-BoldItalic
pilh8r   Caecilia-Heavy
pilhc8r  Caecilia-HeavySC
pilhci8r Caecilia-HeavyItalicSC
pilhi8r  Caecilia-HeavyItalic
pill8r   Caecilia-Light
pillc8r  Caecilia-LightSC
pillci8r Caecilia-LightItalicSC
pilli8r  Caecilia-LightItalic
pilr8r   Caecilia-Roman
pilrc8r  Caecilia-RomanSC
pilrci8r Caecilia-ItalicSC
pilri8r  Caecilia-Italic
pimk8r   Imago-Book
pimki8r  Imago-BookItalic
piml8r   Imago-Light
pimli8r  Imago-LightItalic
pimm8r   Imago-Medium
pimmi8r  Imago-MediumItalic
pimx8r   Imago-ExtraBold
pimxi8r  Imago-ExtraBoldItalic
pinr7a   Industria-SolidA
pinr8r   Industria-Solid
pinrh7a  Industria-InlineA
pinrh8r  Industria-Inline
pipb8r   Impressum-Bold
pipr8r   Impressum-Roman
pipri8r  Impressum-Italic
pitb8r   Italia-Bold

pitk8r   Italia-Book
pitm8r   Italia-Medium
piwr8r   Ironwood
pjnb7d   JansonText-BoldOsF
pjnb8r   JansonText-Bold
pjnbi7d  JansonText-BoldItalicOsF
pjnbi8r  JansonText-BoldItalic
pjnr8r   JansonText-Roman
pjnrc8r  JansonText-RomanSC
pjnri7d  JansonText-ItalicOsF
pjnri8r  JansonText-Italic
pjpr8r   Juniper
pkbc8r   Kabel-Black
pkbh8r   Kabel-Heavy
pkbk8r   Kabel-Book
pkbl8r   Kabel-Light
pkrb8r   Korinna-Bold
pkrbi8r  Korinna-KursivBold
pkrr8r   Korinna-Regular
pkrri8r  Korinna-KursivRegular
pksb28r  KuenstlerScript-TwoBold
pksc8r   KuenstlerScript-Black
pksm8r   KuenstlerScript-Medium
pkub8r   Kaufmann-Bold
pkur8r   Kaufmann
plbd8r   LubalinGraph-Demi
plbdo8r  LubalinGraph-DemiOblique
plbk8r   LubalinGraph-Book
plbko8r  LubalinGraph-BookOblique
plcb8r   Lucida-Bold
plcbi8r  Lucida-BoldItalic
plcbs8r  LucidaSans-Bold
plcbsi8r LucidaSans-BoldItalic
plcr8r   Lucida
plcri8r  Lucida-Italic
plcrm8r  LucidaMath-Italic
plcrs8r  LucidaSans
plcrsi8r LucidaSans-Italic
plcrv8r  LucidaMath-Extension
plcry8r  LucidaMath-Symbol
plfb8r   Life-Bold
plfr8r   Life-Roman
plfri8r  Life-Italic
plgb8r   LetterGothic-Bold
plgbo8r  LetterGothic-BoldSlanted
plgr8r   LetterGothic
plgro8r  LetterGothic-Slanted
plnr8r   Linotext

plnrw8r  Linoscript
plob8r   Lithos-Bold
ploc8r   Lithos-Black
ploi8r   Lithos-ExtraLight
plol8r   Lithos-Light
plor8r   Lithos-Regular
plsr8r   DINMittelschrift
plwb8r   Leawood-Bold
plwbi8r  Leawood-BoldItalic
plwc8r   Leawood-Black
plwci8r  Leawood-BlackItalic
plwk8r   Leawood-Book
plwki8r  Leawood-BookItalic
plwm8r   Leawood-Medium
plwmi8r  Leawood-MediumItalic
pmdb8r   Meridien-Bold
pmdbi8r  Meridien-BoldItalic
pmdm8r   Meridien-Medium
pmdmi8r  Meridien-MediumItalic
pmdr8r   Meridien-Roman
pmdri8r  Meridien-Italic
pmhb8r   Machine-Bold
pmhr8r   Machine
pmirw8r  MediciScript
pmlb8r   Melior-Bold
pmlbi8r  Melior-BoldItalic
pmlr8r   Melior
pmlri8r  Melior-Italic
pmnb8r   Minion-Bold
pmnb8x   MinionExp-Bold
pmnbi8r  Minion-BoldItalic
pmnbi8x  MinionExp-BoldItalic
pmnc8r   Minion-Black
pmnc8x   MinionExp-Black
pmnr8r   Minion-Regular
pmnr8x   MinionExp-Regular
pmnrd8r  Minion-DisplayRegular
pmnrd8x  MinionExp-DisplayRegular
pmnrdi8r Minion-DisplayItalic
pmnrdi8x MinionExp-DisplayItalic
pmnri8r  Minion-Italic
pmnri8x  MinionExp-Italic
pmnrp8r  Minion-Ornaments
pmnrwdi8r Minion-SwashDisplayItalic
pmnrwi8r Minion-SwashItalic
pmns8r   Minion-Semibold
pmns8x   MinionExp-Semibold
pmnsi8r  Minion-SemiboldItalic

pmnsi8x  MinionExp-SemiboldItalic
pmnswi8r Minion-SwashSemiboldItalic
pmpb8r   Memphis-Bold
pmpbi8r  Memphis-BoldItalic
pmpl8r   Memphis-Light
pmpli8r  Memphis-LightItalic
pmpm8r   Memphis-Medium
pmpmi8r  Memphis-MediumItalic
pmpx8r   Memphis-ExtraBold
pmqr8r   Mesquite
pmrr8r   Madrone
pmsr8r   Mistral
pmtb8r   Minister-Bold
pmtbi8r  Minister-BoldItalic
pmtc8r   Minister-Black
pmtci8r  Minister-BlackItalic
pmtk8r   Minister-Book
pmtki8r  Minister-BookItalic
pmtl8r   Minister-Light
pmtli8r  Minister-LightItalic
pmxr8r   Maximus
pnab8r   NewAster-Bold
pnabi8r  NewAster-BoldItalic
pnac8r   NewAster-Black
pnaci8r  NewAster-BlackItalic
pnar8r   NewAster
pnari8r  NewAster-Italic
pnas8r   NewAster-SemiBold
pnasi8r  NewAster-SemiBoldItalic
pnbb8r   NewBaskerville-Bold
pnbbc8r  NewBaskerville-BoldSC
pnbbi7d  NewBaskerville-BoldItalicOsF
pnbbi8r  NewBaskerville-BoldItalic
pnbr8r   NewBaskerville-Roman
pnbrc8r  NewBaskerville-SC
pnbri7d  NewBaskerville-ItalicOsF
pnbri8r  NewBaskerville-Italic
pncb7f   NewCenturySchlbk-FractionBold
pncb8r   NewCenturySchlbk-Bold
pncbi8r  NewCenturySchlbk-BoldItalic
pncbk8r  NewCenturySchlbkGreek-Bd
pncr7f   NewCenturySchlbk-Fraction
pncr8r   NewCenturySchlbk-Roman
pncri8r  NewCenturySchlbk-Italic
pncrik8r NewCenturySchlbkGreek-Ic
pncrk8r  NewCenturySchlbkGreek-Up
pndb8r   NewCaledonia-Bold
pndbc8r  NewCaledonia-BoldSC

pndbi7d  NewCaledonia-BoldItalicOsF
pndbi8r  NewCaledonia-BoldItalic
pndc8r   NewCaledonia-Black
pndci8r  NewCaledonia-BlackItalic
pndr8r   NewCaledonia
pndrc8r  NewCaledonia-SC
pndri7d  NewCaledonia-ItalicOsF
pndri8r  NewCaledonia-Italic
pnds8r   NewCaledonia-SemiBold
pndsi8r  NewCaledonia-SemiBoldItalic
pnfb8r   Nofret-Bold
pnfbi8r  Nofret-BoldItalic
pnfl8r   Nofret-Light
pnfli8r  Nofret-LightItalic
pnfm8r   Nofret-Medium
pnfmi8r  Nofret-MediumItalic
pnfr8r   Nofret-Regular
pnfri8r  Nofret-Italic
pngb8r   NewsGothic-Bold
pngbo8r  NewsGothic-BoldOblique
pngr8r   NewsGothic
pngro8r  NewsGothic-Oblique
pnsb8r   TimesNewRomanPS-Bold
pnsbi8r  TimesNewRomanPS-BoldItalic
pnsr8r   TimesNewRomanPS
pnsri8r  TimesNewRomanPS-Italic
pnurw8r  NuptialScript
pnvb8r   Novarese-Bold
pnvbi8r  Novarese-BoldItalic
pnvk8r   Novarese-Book
pnvki8r  Novarese-BookItalic
pnvm8r   Novarese-Medium
pnvmi8r  Novarese-MediumItalic
pnvu8r   Novarese-Ultra
pnzh8r   NeuzeitS-BookHeavy
pnzk8r   NeuzeitS-Book
po7r8r   OldStyleSeven
po7rc8r  OldStyleSeven-SC
po7ri7d  OldStyleSeven-ItalicOsF
po7ri8r  OldStyleSeven-Italic
poar8r   OCRA
pobr8r   OCRB
pocb8r   Concorde-Bold
pocbi8r  Concorde-BoldItalic
pocr8r   Concorde-Roman
pocri8r  Concorde-Italic
pomr8r   Omnia
ponb8r   Corona-Bold

ponr8r   Corona
ponri8r  Corona-Italic
popb8r   Optima-Bold
popbo8r  Optima-BoldOblique
popr8r   Optima
popro8r  Optima-Oblique
porr8r   Orator
porro8r  Orator-Slanted
poyb8r   Olympian-Bold
poybi8r  Olympian-BoldItalic
poyr8r   Olympian
poyri8r  Olympian-Italic
pp0r8r   Poplar
ppar8r   ParkAvenue
ppeb8r   PrestigeElite-Bold
ppebo8r  PrestigeElite-BoldSlanted
pper8r   PrestigeElite
ppero8r  PrestigeElite-Slanted
ppgb8r   Peignot-Bold
ppgd8r   Peignot-Demi
ppgl8r   Peignot-Light
ppi001   BorderPi-OneFiveOneFiveNine
ppi002   BundesbahnPi-One
ppi003   BundesbahnPi-Two
ppi004   BundesbahnPi-Three
ppi005   EuropeanPi-Four
ppi006   EuropeanPi-One
ppi007   EuropeanPi-Three
ppi008   EuropeanPi-Two
ppi009   LAstrologyPi-One
ppi010   LAstrologyPi-Two
ppi011   LDecorationPi-One
ppi012   LDecorationPi-Two
ppi013   LGamePi-ChessDraughts
ppi014   LGamePi-DiceDominoes
ppi015   LGamePi-EnglishCards
ppi016   LGamePi-FrenchCards
ppi017   LHolidayPi-One
ppi018   LHolidayPi-Two
ppi019   LHolidayPi-Three
ppi020   MathematicalPi-One
ppi021   MathematicalPi-Two
ppi022   MathematicalPi-Three
ppi023   MathematicalPi-Four
ppi024   MathematicalPi-Five
ppi025   MathematicalPi-Six
ppi026   WoodtypeOrnaments-One
ppi027   WoodtypeOrnaments-Two

pplb7d   Palatino-BoldOsF
pplb8r   Palatino-Bold
pplbi7d  Palatino-BoldItalicOsF
pplbi8r  Palatino-BoldItalic
pplr8r   Palatino-Roman
pplrc8r  Palatino-SC
pplri7d  Palatino-ItalicOsF
pplri8r  Palatino-Italic
ppnb8r   Plantin-Bold
ppnb8x   PlantinExpert-Bold
ppnbi8r  Plantin-BoldItalic
ppnbi8x  PlantinExpert-BoldItalic
ppnbr8rc Plantin-BoldCondensed
ppnl8r   Plantin-Light
ppnli8r  Plantin-LightItalic
ppnr8r   Plantin
ppnr8x   PlantinExpert
ppnri8r  Plantin-Italic
ppnri8x  PlantinExpert-Italic
ppns8r   Plantin-Semibold
ppnsi8r  Plantin-SemiboldItalic
ppor8r   Ponderosa
pppb8r   Perpetua-Bold
pppbi8r  Perpetua-BoldItalic
pppr8r   Perpetua
pppri8r  Perpetua-Italic
ppqb8r   PostAntiqua-Bold
ppqr8r   PostAntiqua-Roman
ppsr8r   Parisian
pptr8r   Present
ppxb8r   PopplPontifex-Bold
ppxbr8rc PopplPontifex-BoldCondensed
ppxc8r   PopplPontifex-Black
ppxr8r   PopplPontifex-Roman
ppxr8rc  PopplPontifex-RomanSC
ppxri8r  PopplPontifex-Italic
pqrb8r   Quorum-Bold
pqrc8r   Quorum-Black
pqrk8r   Quorum-Book
pqrl8r   Quorum-Light
pqrm8r   Quorum-Medium
pr1b8r   RotisSemiSerif-Bold
pr1bs8r  RotisSemiSans-Bold
pr1lis8r RotisSemiSans-LightItalic
pr1ls8r  RotisSemiSans-Light
pr1r8r   RotisSemiSerif
pr1ris8r RotisSemiSans-Italic
pr1rs8r  RotisSemiSans

pr1xs8r  RotisSemiSans-ExtraBold
prab8r   Raleigh-Bold
prad8r   Raleigh-DemiBold
pram8r   Raleigh-Medium
prar8r   Raleigh
prob8r   RotisSerif-Bold
probs8r  RotisSansSerif-Bold
prolis8r RotisSansSerif-LightItalic
prols8r  RotisSansSerif-Light
pror8r   RotisSerif
prori8r  RotisSerif-Italic
proris8r RotisSansSerif-Italic
prors8r  RotisSansSerif
proxs8r  RotisSansSerif-ExtraBold
prpr28r  Reporter-Two
prqr8r   RussellSquare
prqro8r  RussellSquare-Oblique
prtr8r   Carta
prvr8r   Revue
prwb8r   Rockwell-Bold
prwbi8r  Rockwell-BoldItalic
prwbr8rc Rockwell-BoldCondensed
prwl8r   Rockwell-Light
prwli8r  Rockwell-LightItalic
prwr8r   Rockwell
prwri8r  Rockwell-Italic
prwrr8rc Rockwell-Condensed
prwx8r   Rockwell-ExtraBold
ps0r7c8r SanMarco-Dfr
ps0r8r   SanMarco
ps7rp    Sassoon-Primary
psbb7d   Sabon-BoldOsF
psbb8r   Sabon-Bold
psbbi7d  Sabon-BoldItalicOsF
psbbi8r  Sabon-BoldItalic
psbr8r   Sabon-Roman
psbrc8r  Sabon-RomanSC
psbri7d  Sabon-ItalicOsF
psbri8r  Sabon-Italic
pscb8r   Slimbach-Bold
pscbi8r  Slimbach-BoldItalic
pscc8r   Slimbach-Black
pscci8r  Slimbach-BlackItalic
psck8r   Slimbach-Book
pscki8r  Slimbach-BookItalic
pscm8r   Slimbach-Medium
pscmi8r  Slimbach-MediumItalic
psfb8r   Serifa-Bold

psfc8r   Serifa-Black
psfl8r   Serifa-Light
psfli8r  Serifa-LightItalic
psfr8r   Serifa-Roman
psfri8r  Serifa-Italic
psgb8r   SerifGothic-Bold
psgc8r   SerifGothic-Black
psgh8r   SerifGothic-Heavy
psgl8r   SerifGothic-Light
psgr8r   SerifGothic
psgx8r   SerifGothic-ExtraBold
pshb8r   Shannon-Bold
pshk8r   Shannon-Book
pshro8r  Shannon-Oblique
pshx8r   Shannon-ExtraBold
pslr8r   Stencil
psnh8r   Spartan-HeavyClassified
psnk8r   Spartan-BookClassified
psor8r   Sonata
pspb8r   Serpentine-Bold
pspbo8r  Serpentine-BoldOblique
pspl8r   Serpentine-Light
psplo8r  Serpentine-LightOblique
pspm8r   Serpentine-Medium
pspmo8r  Serpentine-MediumOblique
psrbw8r  SnellRoundhand-BoldScript
psrcw8r  SnellRoundhand-BlackScript
psrrw8r  SnellRoundhand-Script
pssb8r   StempelSchneidler-Bold
pssbi8r  StempelSchneidler-BoldItalic
pssc8r   StempelSchneidler-Black
pssci8r  StempelSchneidler-BlackItalic
pssl8r   StempelSchneidler-Light
pssli8r  StempelSchneidler-LightItalic
pssm8r   StempelSchneidler-Medium
pssmi8r  StempelSchneidler-MedItalic
pssr8r   StempelSchneidler-Roman
pssri8r  StempelSchneidler-Italic
pstb8r   StoneSerif-Bold
pstbi8r  StoneSerif-BoldItalic
pstbin8r StoneInformal-BoldItalic
pstbis8r StoneSans-BoldItalic
pstbn8r  StoneInformal-Bold
pstbs8r  StoneSans-Bold
pstr5a   StoneSerif-PhoneticAlternate
pstr5i   StoneSerif-PhoneticIPA
pstr8r   StoneSerif
pstri8r  StoneSerif-Italic

pstrin8r StoneInformal-Italic
pstris8r StoneSans-Italic
pstrn8r  StoneInformal
pstrs5a  StoneSans-PhoneticAlternate
pstrs5i  StoneSans-PhoneticIPA
pstrs8r  StoneSans
psts8r   StoneSerif-Semibold
pstsi8r  StoneSerif-SemiboldItalic
pstsin8r StoneInformal-SemiboldItalic
pstsis8r StoneSans-SemiboldItalic
pstsn8r  StoneInformal-Semibold
pstss8r  StoneSans-Semibold
psvb8r   Souvenir-Bold
psvbi8r  Souvenir-BoldItalic
psvd8r   Souvenir-Demi
psvdi8r  Souvenir-DemiItalic
psvl8r   Souvenir-Light
psvli8r  Souvenir-LightItalic
psvm8r   Souvenir-Medium
psvmi8r  Souvenir-MediumItalic
psxb8r   Syntax-Bold
psxc8r   Syntax-Black
psxr8r   Syntax-Roman
psxri8r  Syntax-Italic
psxu8r   Syntax-UltraBlack
psyb     Symbol-Bold
psybi    Symbol-BoldItalic
psyc     Symbol-Black
psyci    Symbol-BlackItalic
psyk     Symbol-Book
psyki    Symbol-BookItalic
psym     Symbol-Medium
psymi    Symbol-MediumItalic
psyr     Symbol
ptbb8r   BBodoni-Bold
ptbbi8r  BBodoni-BoldItalic
ptbl8r   BBodoni-Light
ptbli8r  BBodoni-LightItalic
ptbm8r   BBodoni-Medium
ptbmi8r  BBodoni-MediumItalic
ptbr8r   BBodoni-Regular
ptbri8r  BBodoni-Italic
pteb8r   TimesEuropa-Bold
ptebi8r  TimesEuropa-BoldItalic
pter8r   TimesEuropa-Roman
pteri8r  TimesEuropa-Italic
ptfd8r   Tiffany-Demi
ptfdi8r  Tiffany-DemiItalic

ptfh8r   Tiffany-Heavy
ptfhi8r  Tiffany-HeavyItalic
ptfr8r   Tiffany
ptfri8r  Tiffany-Italic
ptgb28r  TradeGothic-BoldTwo
ptgb2o8r TradeGothic-BoldTwoOblique
ptgb8r   TradeGothic-Bold
ptgb8r20c TradeGothic-BoldCondTwenty
ptgb8r20o TradeGothic-BoldCondTwentyObl
ptgbo8r  TradeGothic-BoldOblique
ptgl8r   TradeGothic-Light
ptglo8r  TradeGothic-LightOblique
ptgr188rc TradeGothic-CondEighteen
ptgr18o8r TradeGothic-CondEighteenObl
ptgr8r   TradeGothic
ptgro8r  TradeGothic-Oblique
ptjb8r   Trajan-Bold
ptjr8r   Trajan-Regular
ptkb8r   Tekton-Bold
ptkbo8r  Tekton-BoldOblique
ptkr8r   Tekton
ptkro8r  Tekton-Oblique
ptmb10   TimesTen-Bold
ptmb7a10 TimesTenEastA-Bold
ptmb7d10 TimesTen-BoldOsF
ptmb7di10 TimesTen-BoldItalicOsF
ptmb8r   Times-Bold
ptmbc8r  Times-BoldSC
ptmbi10  TimesTen-BoldItalic
ptmbi7a10 TimesTenEastA-BoldItalic
ptmbi7d  Times-BoldItalicOsF
ptmbi8r  Times-BoldItalic
ptmbk10  TimesTenGreekP-Bold
ptmbko10 TimesTenGreekP-BoldInclined
ptmboz10 TimesTenCyrillic-BoldInclined
ptmbz10  TimesTenCyrillic-Bold
ptmr10   TimesTen-Roman
ptmr7a10 TimesTenEastA-Roman
ptmr7di10 TimesTen-ItalicOsF
ptmr8r   Times-Roman
ptmrc10  TimesTen-RomanSC
ptmrc8r  Times-RomanSC
ptmri10  TimesTen-Italic
ptmri7a10 TimesTenEastA-Italic
ptmri7d  Times-ItalicOsF
ptmri8r  Times-Italic
ptmrk10  TimesTenGreekP-Upright
ptmrko10 TimesTenGreekP-Inclined

ptmroz10 TimesTenCyrillic-Inclined
ptmrz10  TimesTenCyrillic-Upright
ptms8r   Times-Semibold
ptmsi8r  Times-SemiboldItalic
ptmx8r   Times-ExtraBold
ptphi8rc Tempo-HeavyCondensedItalic
ptphr8rc Tempo-HeavyCondensed
ptvb7d   TrumpMediaeval-BoldOsF
ptvb8r   TrumpMediaeval-Bold
ptvbi7d  TrumpMediaeval-BoldItalicOsF
ptvbi8r  TrumpMediaeval-BoldItalic
ptvr8r   TrumpMediaeval-Roman
ptvrc8r  TrumpMediaeval-SC
ptvri7d  TrumpMediaeval-ItalicOsF
ptvri8r  TrumpMediaeval-Italic
pubb7d   BauerBodoni-BoldOsF
pubb8r   BauerBodoni-Bold
pubbi7d  BauerBodoni-BoldItalicOsF
pubbi8r  BauerBodoni-BoldItalic
pubbr8rc BauerBodoni-BoldCond
pubc8r   BauerBodoni-Black
pubci8r  BauerBodoni-BlackItalic
pubcr8rc BauerBodoni-BlackCond
pubr8r   BauerBodoni-Roman
pubrc8r  BauerBodoni-RomanSC
pubri7d  BauerBodoni-ItalicOsF
pubri8r  BauerBodoni-Italic
pugb8r   BenguiatGothic-Bold
pugbo8r  BenguiatGothic-BoldOblique
pugh8r   BenguiatGothic-Heavy
pugho8r  BenguiatGothic-HeavyOblique
pugk8r   BenguiatGothic-Book
pugko8r  BenguiatGothic-BookOblique
pugm8r   BenguiatGothic-Medium
pugmo8r  BenguiatGothic-MediumOblique
puhb8r   Cushing-Bold
puhbi8r  Cushing-BoldItalic
puhh8r   Cushing-Heavy
puhhi8r  Cushing-HeavyItalic
puhk8r   Cushing-Book
puhki8r  Cushing-BookItalic
puhm8r   Cushing-Medium
puhmi8r  Cushing-MediumItalic
pulrk8r  Universal-GreekwithMathPi
pumr8r   Umbra
punaro8r Univers-ThinUltraCondensed
punb8r   Univers-Bold
punbo8r  Univers-BoldOblique

punbo8rc Univers-CondensedBoldOblique
punbo8rx Univers-BoldExtObl
punbr8rc Univers-CondensedBold
punbr8rx Univers-BoldExt
punc8r   Univers-Black
punco8r  Univers-BlackOblique
punco8rx Univers-BlackExtObl
puncr8rx Univers-BlackExt
punl8r   Univers-Light
punlo8r  Univers-LightOblique
punlo8rc Univers-CondensedLightOblique
punlr8rc Univers-CondensedLight
punlro8r Univers-LightUltraCondensed
punr8r   Univers
punro8r  Univers-Oblique
punro8rc Univers-CondensedOblique
punro8rx Univers-ExtendedObl
punrr8rc Univers-Condensed
punrr8rx Univers-Extended
punrro8r Univers-UltraCondensed
punx8r   Univers-ExtraBlack
punxo8r  Univers-ExtraBlackObl
punxo8rx Univers-ExtraBlackExtObl
punxr8rx Univers-ExtraBlackExt
putb8r   Utopia-Bold
putb8x   UtopiaExp-Bold
putbi8r  Utopia-BoldItalic
putbi8x  UtopiaExp-BoldItalic
putc8r   Utopia-Black
putc8x   UtopiaExp-Black
putr8r   Utopia-Regular
putr8x   UtopiaExp-Regular
putrd8r  Utopia-Titling
putri8r  Utopia-Italic
putri8x  UtopiaExp-Italic
putrp8r  Utopia-Ornaments
puts8r   Utopia-Semibold
puts8x   UtopiaExp-Semibold
putsi8r  Utopia-SemiboldItalic
putsi8x  UtopiaExp-SemiboldItalic
puwb8r   Usherwood-Bold
puwbi8r  Usherwood-BoldItalic
puwc8r   Usherwood-Black
puwci8r  Usherwood-BlackItalic
puwk8r   Usherwood-Book
puwki8r  Usherwood-BookItalic
puwm8r   Usherwood-Medium
puwmi8r  Usherwood-MediumItalic

puyr8r   UniversityRoman
pvjb8r   Veljovic-Bold
pvjbi8r  Veljovic-BoldItalic
pvjc8r   Veljovic-Black
pvjci8r  Veljovic-BlackItalic
pvjk8r   Veljovic-Book
pvjki8r  Veljovic-BookItalic
pvjm8r   Veljovic-Medium
pvjmi8r  Veljovic-MediumItalic
pvlb8r   Versailles-Bold
pvlbi8r  Versailles-BoldItalic
pvlc8r   Versailles-Black
pvlci8r  Versailles-BlackItalic
pvll8r   Versailles-Light
pvlli8r  Versailles-LightItalic
pvlr8r   Versailles-Roman
pvlri8r  Versailles-Italic
pvra8r   VAGRounded-Thin
pvrb8r   VAGRounded-Bold
pvrc8r   VAGRounded-Black
pvrl8r   VAGRounded-Light
pwbb7d   Walbaum-BoldOsF
pwbb8r   Walbaum-Bold
pwbbi7d  Walbaum-BoldItalicOsF
pwbbi8r  Walbaum-BoldItalic
pwbr8r   Walbaum-Roman
pwbrc8r  Walbaum-RomanSC
pwbri7d  Walbaum-ItalicOsF
pwbri8r  Walbaum-Italic
pwdb8r   Weidemann-Bold
pwdbi8r  Weidemann-BoldItalic
pwdc8r   Weidemann-Black
pwdci8r  Weidemann-BlackItalic
pwdk8r   Weidemann-Book
pwdki8r  Weidemann-BookItalic
pwdm8r   Weidemann-Medium
pwdmi8r  Weidemann-MediumItalic
pwgr7c   WilhelmKlingsporGotisch-Dfr
pwgr8r   WilhelmKlingsporGotisch
pwkb8r   Wilke-Bold
pwkbi8r  Wilke-BoldItalic
pwkc8r   Wilke-Black
pwkci8r  Wilke-BlackItalic
pwkr8r   Wilke-Roman
pwkri8r  Wilke-Italic
pwsb8r   Weiss-Bold
pwsr8r   Weiss
pwsri8r  Weiss-Italic

pwsx8r   Weiss-ExtraBold
pwwr8r   Willow
pzcb8r   ZapfChancery-Bold
pzcd8r   ZapfChancery-Demi
pzcl8r   ZapfChancery-Light
pzcli8r  ZapfChancery-LightItalic
pzcmi8r  ZapfChancery-MediumItalic
pzcr8r   ZapfChancery-Roman
pzcri8r  ZapfChancery-Italic
pzdr     ZapfDingbats
pzgbr8rc DINNeuzeitGrotesk-BoldCond
pzgl8r   DINNeuzeitGrotesk-Light
zpcp29ab         Copperplate-TwentyNineAB
zpcp29bc         Copperplate-TwentyNineBC
zpcp30ab         Copperplate-ThirtyAB
zpcp30bc         Copperplate-ThirtyBC
zpcp31ab         Copperplate-ThirtyOneAB
zpcp31bc         Copperplate-ThirtyOneBC
zpcp32ab         Copperplate-ThirtyTwoAB
zpcp32bc         Copperplate-ThirtyTwoBC
zpcp33bc         Copperplate-ThirtyThreeBC
zpgbm7h  GothicBBB-Medium-78-H
zpgbm7v  GothicBBB-Medium-78-V
zpgbm8rh GothicBBB-Medium-83pv-RKSJ-H
zpgbmah  GothicBBB-Medium-Add-H
zpgbmav  GothicBBB-Medium-Add-V
zpgbmeh  GothicBBB-Medium-Ext-H
zpgbmev  GothicBBB-Medium-Ext-V
zpgbmh   GothicBBB-Medium-H
zpgbmv   GothicBBB-Medium-V
zpmicr   MICR
zpryl7h  Ryumin-Light-78-H
zpryl7v  Ryumin-Light-78-V
zpryl8rh Ryumin-Light-83pv-RKSJ-H
zprylah  Ryumin-Light-Add-H
zprylav  Ryumin-Light-Add-V
zpryleh  Ryumin-Light-Ext-H
zprylev  Ryumin-Light-Ext-V
zprylh   Ryumin-Light-H
zprylv   Ryumin-Light-V
zpshals  Shelley-AllegroScript
zpshans  Shelley-AndanteScript
zpshvos  Shelley-VolanteScript
zpsrliolh SerlioLH
zpulrncp Universal-NewswithCommPi

A.3 Apple fonts

The ‘apple.map’ file has abbreviations for Apple fonts.

eecr8r   AppleChancery
eger8r   Geneva
ehtr8r   HoeflerText
emcr8r   Monaco
epi001   Zeal
eskr8r   Skia

A.4 Bitstream fonts

The ‘bitstrea.map’ file has abbreviations for Bitstream fonts.

ba3r8r   AdLibBT-Regular                        1154a___
ba4r8r   AlternateGothicNo2BT-Regular           0542a___
ba5r8r   AllegroBT-Regular                      1138a___
ba6r8r   AmeliaBT-Regular                       0510a___
ba7r8r   AmericanTextBT-Regular                 0211a___
ba8b8rc  AuroraBT-BoldCondensed                 1028a___
ba9b8r   Humanist970BT-BoldC                    0265a___
ba9r8r   Humanist970BT-RomanC                   0264a___
baab8r   AachenBT-Bold                          1103a___
baar8r   AachenBT-Roman                         1102a___
balb8r   Flareserif821BT-Bold                   0918a___
ball8r   Flareserif821BT-Light                  0916a___
balr8r   Flareserif821BT-Roman                  0917a___
bamb8r   AmericanaBT-Bold                       0500a___
bamr8r   AmericanaBT-Roman                      0498a___
bamri8r  AmericanaBT-Italic                     0499a___
bamx8r   AmericanaBT-ExtraBold                  0501a___
bamx8rc  AmericanaBT-ExtraBoldCondensed         0502a___
baob8r   Incised901BT-Bold                      0390a___
baob8rc  Incised901BT-BoldCondensed             0467a___
baoc8r   Incised901BT-Black                     0391a___
baoh8rx  Incised901BT-Nord                      0469a___
baohi8rx Incised901BT-NordItalic                0470a___
baol8r   Incised901BT-Light                     0387a___
baop8r   Incised901BT-Compact                   0468a___
baor8r   Incised901BT-Roman                     0388a___
baori8r  Incised901BT-Italic                    0389a___
baub8r   Freeform721BT-Bold                     0999a___
baubi8r  Freeform721BT-BoldItalic               1000a___
bauc8r   Freeform721BT-Black                    1001a___
bauci8r  Freeform721BT-BlackItalic              1002a___
baur8r   Freeform721BT-Roman                    0997a___
bauri8r  Freeform721BT-Italic                   0998a___
bazb8r   Gothic725BT-BoldA                      0541a___
bazc8r   Gothic725BT-BlackA                     1007a___

bb4r8r   Brush738BT-RegularA                    0806a___
bb5r8r   Freehand591BT-RegularA                 1043a___
bb6r8rc  Gothic821CondensedBT-Regular           0807a___
bb7b8r   BernhardModernBT-Bold                  0838a___
bb7b8rc  BernhardBoldCondensedBT-Regular        1003a___
bb7bi8r  BernhardModernBT-BoldItalic            0839a___
bb7r8r   BernhardModernBT-Roman                 0607a___
bb7ri8r  BernhardModernBT-Italic                0837a___
bb7rn8r  BernhardFashionBT-Regular              1051a___
bb7rw8r  BernhardTangoBT-Regular                        0986a___
bb8c8r   BlippoBT-Black                         0511a___
bb9b8r   BremenBT-Bold                          1231a___
bb9c8r   BremenBT-Black                         1232a___
bbbb8a   Aldine401BT-BoldA                      0304a___
bbbbi8a  Aldine401BT-BoldItalicA                0305a___
bbbr8a   Aldine401BT-RomanA                     0302a___
bbbri8a  Aldine401BT-ItalicA                    0303a___
bbdb8r   BodoniBT-Bold                          0037a___
bbdb8rc  BodoniBT-BoldCondensed                 0973a___
bbdbi8r  BodoniBT-BoldItalic                    0038a___
bbdk8r   BodoniBT-Book                          0035a___
bbdki8r  BodoniBT-BookItalic                    0036a___
bbdp8r   PosterBodoniBT-Roman                   0129a___
bbdpi8r  PosterBodoniBT-Italic                  0130a___
bbdr8r   BodoniBT-Roman                         0055a___
bbdri8r  BodoniBT-Italic                        0056a___
bblb8r   BelweBT-Bold                           1024a___
bbll8r   BelweBT-Light                          1022a___
bblm8r   BelweBT-Medium                         1023a___
bblr8rc  BelweBT-RomanCondensed                 1025a___
bbob8r   BalloonBT-Bold                         0728a___
bbol8r   BalloonBT-Light                        0121a___
bbox8r   BalloonBT-ExtraBold                    0729a___
bbpr8r   Modern735BT-RomanA                     0793a___
bbrb8r   Revival565BT-Bold                      1070a___
bbrbi8r  Revival565BT-BoldItalic                1071a___
bbrr8r   Revival565BT-Roman                     1068a___
bbrri8r  Revival565BT-Italic                    1069a___
bbsr8r   BakerSignetBT-Roman                    1152a___
bburw8r  BrushScriptBT-Regular                  0199a___
bbwr8r   BroadwayBT-Regular                     0131a___
bbwre8r  BroadwayEngravedBT-Regular             0802a___
bc5r8r   Caslon540BT-Roman                      0118a___
bc5ri8r  Caslon540BT-Italic                     0119a___
bc7b8r   CompactaBT-Bold                        0786a___
bc7bi8r  CompactaBT-BoldItalic                  0787a___
bc7c8r   CompactaBT-Black                       0788a___
bc7r8r   CompactaBT-Roman                       0784a___

bc7ri8r  CompactaBT-Italic                      0785a___
bc9r8r   Calligraphic421BT-RomanB               1139a___
bcab8r   CaslonBT-Bold                          0443a___
bcabi8r  CaslonBT-BoldItalic                    0444a___
bcarl8r  CaslonOpenfaceBT-Regular               0581a___
bcbb8r   CooperBT-Bold                          0579a___
bcbbi8r  CooperBT-BoldItalic                    0580a___
bcbbl8r  CooperBT-BlackOutline                  1034a___
bcbc8r   CooperBT-Black                         0630a___
bcbci8r  CooperBT-BlackItalic                   0631a___
bcbcil8r CooperBT-BlackItalicHeadline           0232a___
bcbcl8r  CooperBT-BlackHeadline                 0231a___
bcbl8r   CooperBT-Light                         0575a___
bcbli8r  CooperBT-LightItalic                   0576a___
bcbm8r   CooperBT-Medium                        0577a___
bcbmi8r  CooperBT-MediumItalic                  0578a___
bccr8r   Freehand471BT-Regular                  0418a___
bcdb8r   ClarendonBT-Bold                       0284a___
bcdb8rc  ClarendonBT-BoldCondensed              0228a___
bcdc8r   ClarendonBT-Black                      0351a___
bcdh8r   ClarendonBT-Heavy                      0416a___
bcdl8r   ClarendonBT-Light                      0282a___
bcdr8r   ClarendonBT-Roman                      0283a___
bcdr8rc  ClarendonBT-RomanCondensed             0846a___
bchb8r   CharterBT-Bold
bchbi8r  CharterBT-BoldItalic
bchr8r   CharterBT-Roman
bchri8r  CharterBT-Italic
bclc8r   CloisterBlackBT-Regular                0757a___
bcnb8rx    CenturyExpandedBT-Bold                 0041a___
bcnbi8rx CenturyExpandedBT-BoldItalic           0042a___
bcnr8rx  CenturyExpandedBT-Roman                0039a___
bcnri8rx CenturyExpandedBT-Italic               0040a___
bcpbs8r  CopperplateGothicBT-Bold               0421a___
bcpbs8rc CopperplateGothicBT-BoldCond           0424a___
bcphs8r  CopperplateGothicBT-Heavy              0422a___
bcprs8r  CopperplateGothicBT-Roman              0420a___
bcprs8rc CopperplateGothicBT-RomanCond          0423a___
bcsb8r   CenturySchoolbookBT-Bold               0085a___
bcsb8rc  CenturySchoolbookBT-BoldCond           0574a___
bcsbi8r  CenturySchoolbookBT-BoldItalic         0086a___
bcsr8r   CenturySchoolbookBT-Roman              0083a___
bcsri8r  CenturySchoolbookBT-Italic             0084a___
bcsrt8r  CenturySchoolbookBT-Monospace          0600a___
bctb8r   CheltenhamBT-Bold                      0622a___
bctb8rc  CheltenhamBT-BoldCondensed             1065a___
bctb8rq  CheltenhamBT-BoldExtraCondensed        1067a___
bctbd8r  CheltenhamBT-BoldHeadline              0508a___

bctbdi8r CheltenhamBT-BoldItalicHeadline        0509a___
bctbi8r  CheltenhamBT-BoldItalic                0623a___
bctbi8rc CheltenhamBT-BoldCondItalic            1066a___
bctr8r   CheltenhamBT-Roman                     0620a___
bctri8r  CheltenhamBT-Italic                    0621a___
bcub8r   CenturyOldstyleBT-Bold                 0287a___
bcur8r   CenturyOldstyleBT-Roman                0285a___
bcuri8r  CenturyOldstyleBT-Italic               0286a___
bcyb8r   SquareSlabserif711BT-Bold              0412a___
bcyl8r   SquareSlabserif711BT-Light             0410a___
bcym8r   SquareSlabserif711BT-Medium            0411a___
bczr8r   Staccato555BT-RegularA                 1153a___
bd0b8r   DavidaBoldBT-Regular                   0195a___
bdcb8r   DomBoldBT-Regular                      0769a___
bdcbi8r  DomDiagonalBT-Bold                     0921a___
bdcr8r   DomCasualBT-Regular                    0604a___
bdcri8r  DomDiagonalBT-Regular                  0920a___
bdeb8r   DellaRobbiaBT-Bold                     1113a___
bder8r   DellaRobbiaBT-Roman                    1112a___
bdir8r   Calligraphic810BT-Roman                0811a___
bdiri8r  Calligraphic810BT-Italic               0812a___
be0r8r   EmbassyBT-Regular                      0588a___
be1r8r   EmpireBT-Regular                       1176a___
beab8r   Transitional521BT-BoldA                0915a___
bear8r   Transitional521BT-RomanA               0913a___
bearw8r  Transitional521BT-CursiveA             0914a___
beeb8r   Egyptian505BT-Bold                     0964a___
beel8r   Egyptian505BT-Light                    0961a___
beem8r   Egyptian505BT-Medium                   0963a___
beer8r   Egyptian505BT-Roman                    0962a___
begb8r   OriginalGaramondBT-Bold                0865a___
begbi8r  OriginalGaramondBT-BoldItalic          0866a___
begr8r   OriginalGaramondBT-Roman               0863a___
begri8r  OriginalGaramondBT-Italic              0864a___
bekr8r   Freeform710BT-Regular                  0714a___
benb8r   EngraversRomanBT-Bold                  0974a___
benbf8r  EngraversOldEnglishBT-Bold             1045a___
benr8r   EngraversRomanBT-Regular               0587a___
benrf8r  EngraversOldEnglishBT-Regular          0212a___
benrs8r  EngraversGothicBT-Regular              0586a___
beob8a   AmerigoBT-BoldA                        0645a___
beobi8a  AmerigoBT-BoldItalicA                  0772a___
beom8a   AmerigoBT-MediumA                      0770a___
beomi8a  AmerigoBT-MediumItalicA                0771a___
beor8a   AmerigoBT-RomanA                       0643a___
beori8a  AmerigoBT-ItalicA                      0644a___
besb8r   Square721BT-Bold                       0247a___
besb8rc  Square721BT-BoldCondensed              0249a___

besb8rx  Square721BT-BoldExtended               0251a___
besr8r   Square721BT-Roman                      0246a___
besr8rc  Square721BT-RomanCondensed             0248a___
besr8rx  Square721BT-RomanExtended              0250a___
bf0r8r   FlemishScriptBT-Regular                0591a___
bffm8r   Transitional551BT-MediumB              0884a___
bffmi8r  Transitional551BT-MediumItalicB        0885a___
bfrb8r   Humanist777BT-BoldB                    0859a___
bfrbi8r  Humanist777BT-BoldItalicB              0860a___
bfrc8r   Humanist777BT-BlackB                   0861a___
bfrci8r  Humanist777BT-BlackItalicB             0862a___
bfrl8r   Humanist777BT-LightB                   0855a___
bfrli8r  Humanist777BT-LightItalicB             0856a___
bfrr8r   Humanist777BT-RomanB                   0857a___
bfrri8r  Humanist777BT-ItalicB                  0858a___
bftr8r   FrakturBT-Regular                      0983a___
bfuc8r   FuturaBlackBT-Regular                  0206a___
bg0r8r   GandoBT-Regular                        0625a___
bg3b8r   AmericanGaramondBT-Bold                0045a___
bg3bi8r  AmericanGaramondBT-BoldItalic          0046a___
bg3r8r   AmericanGaramondBT-Roman               0043a___
bg3ri8r  AmericanGaramondBT-Italic              0044a___
bg5r8r   GoudyCatalogueBT-Regular               1054a___
bg6r8r   GorillaITCbyBT-Regular                 0664a___
bgjb8r   ElegantGaramondBT-Bold                 0713a___
bgjr8r   ElegantGaramondBT-Roman                0711a___
bgjri8r  ElegantGaramondBT-Italic               0712a___
bgob8r   GoudyOldStyleBT-Bold                   0109a___
bgobi8r  GoudyOldStyleBT-BoldItalic             0110a___
bgor8r   GoudyOldStyleBT-Roman                  0107a___
bgori8r  GoudyOldStyleBT-Italic                 0108a___
bgox8r   GoudyOldStyleBT-ExtraBold              0111a___
bgsb8r   Humanist521BT-Bold                     0292a___
bgsb8rc  Humanist521BT-BoldCondensed            0496a___
bgsbi8r  Humanist521BT-BoldItalic               0293a___
bgsl8r   Humanist521BT-Light                    0288a___
bgsli8r  Humanist521BT-LightItalic              0289a___
bgsr8r   Humanist521BT-Roman                    0290a___
bgsr8rc  Humanist521BT-RomanCondensed           0495a___
bgsri8r  Humanist521BT-Italic                   0291a___
bgsu8r   Humanist521BT-UltraBold                0295a___
bgsx8r   Humanist521BT-ExtraBold                0294a___
bgsx8rc  Humanist521BT-XtraBoldCondensed        0497a___
bgyh8r   GoudyHeavyfaceBT-Regular               1049a___
bgyrf8r  GoudyHandtooledBT-Regular              1053a___
bgyrrc8r GoudyHeavyfaceBT-RegularCond           1050a___
bh0r8r   HandelGothicBT-Regular                 0768a___
bhbr8r   HoboBT-Regular                         0057a___

bhir8r   Swiss921BT-RegularA                    0562a___
bhrb8r   Swiss721BT-BoldRounded                 0813a___
bhrc8r   Swiss721BT-BlackRounded                0815a___
bhsr8r   Swiss924BT-RegularB                    0969a___
bhva8r   Swiss721BT-Thin                        0100a___
bhvai8r  Swiss721BT-ThinItalic                  0101a___
bhvb8r   Swiss721BT-Bold                        0005a___
bhvb8rc  Swiss721BT-BoldCondensed               0010a___
bhvb8rx  Swiss721BT-BoldExtended                0373a___
bhvbi8r  Swiss721BT-BoldItalic                  0006a___
bhvbi8rc Swiss721BT-BoldCondensedItalic         0172a___
bhvbit8r Monospace821BT-BoldItalic              0599a___
bhvbl8r  Swiss721BT-BoldOutline                 0175a___
bhvbl8rc Swiss721BT-BoldCondensedOutline        0767a___
bhvbt8r  Monospace821BT-Bold                    0598a___
bhvc8r   Swiss721BT-Black                       0007a___
bhvc8rc  Swiss721BT-BlackCondensed              0173a___
bhvc8rx  Swiss721BT-BlackExtended               0374a___
bhvci8r  Swiss721BT-BlackItalic                 0008a___
bhvci8rc Swiss721BT-BlackCondensedItalic        0174a___
bhvcl8r  Swiss721BT-BlackOutline                0563a___
bhvh8r   Swiss721BT-Heavy                       0102a___
bhvhi8r  Swiss721BT-HeavyItalic                 0103a___
bhvl8r   Swiss721BT-Light                       0001a___
bhvl8rc  Swiss721BT-LightCondensed              0169a___
bhvl8rx  Swiss721BT-LightExtended               0371a___
bhvli8r  Swiss721BT-LightItalic                 0002a___
bhvli8rc Swiss721BT-LightCondensedItalic        0170a___
bhvm8r   Swiss721BT-Medium                      0759a___
bhvmi8r  Swiss721BT-MediumItalic                0760a___
bhvr8r   Swiss721BT-Roman                       0003a___
bhvr8rc  Swiss721BT-RomanCondensed              0009a___
bhvr8rq  Swiss911BT-ExtraCompressed             0105a___
bhvr8ru  Swiss911BT-UltraCompressed             0106a___
bhvr8rx  Swiss721BT-RomanExtended               0372a___
bhvri8r  Swiss721BT-Italic                      0004a___
bhvri8rc Swiss721BT-ItalicCondensed             0171a___
bhvrit8r Monospace821BT-Italic                  0597a___
bhvrt8r  Monospace821BT-Roman                   0596a___
bhwc8rc  SchadowBT-BlackCondensed               1114a___
bhxr8r   HuxleyVerticalBT-Regular               1177a___
bi0r8r   CommercialScriptBT-Regular             0976a___
bi1r8r   ImpressBT-Regular                      0209a___
bi2r8r   ImpulsBT-Regular                       0725a___
biob8r   News701BT-BoldA                        0320a___
bior8r   News701BT-RomanA                       0318a___
biori8r  News701BT-ItalicA                      0319a___
biyb8r   IowanOldStyleBT-Bold                   1200a___

biybi8r  IowanOldStyleBT-BoldItalic             1201a___
biyc8r   IowanOldStyleBT-Black                  1202a___
biyci8r  IowanOldStyleBT-BlackItalic            1203a___
biyr8r   IowanOldStyleBT-Roman                  1198a___
biyri8r  IowanOldStyleBT-Italic                 1199a___
bjfr8r   Freehand575BT-RegularB                 1046a___
bkbb8r   Geometric231BT-BoldC                   1128a___
bkbh8r   Geometric231BT-HeavyC                  1129a___
bkbl8r   Geometric231BT-LightC                  1126a___
bkbr8r   Geometric231BT-RomanC                  1127a___
bkgl8r   BankGothicBT-Light                     0589a___
bkgm8r   BankGothicBT-Medium                    0590a___
bkub8r   KaufmannBT-Bold                        0727a___
bkur8r   KaufmannBT-Regular                     0726a___
bl1r8r   LibertyBT-Regular                      0987a___
bl2r8r   LibraBT-Regular                        0690a___
bl3r8r   LuciaBT-Regular                        0847a___
bl4b8r   LucianBT-Bold                          1143a___
bl4r8r   LucianBT-Roman                         1142a___
blgb8r   LetterGothic12PitchBT-Bold             0533a___
blgbi8r  LetterGothic12PitchBT-BoldItal         0543a___
blgr8r   LetterGothic12PitchBT-Roman            0414a___
blgri8r  LetterGothic12PitchBT-Italic           0532a___
blir8rc  LatinExtraCondensedBT-Regular          0507a___
blyb8r   LydianBT-Bold                          0843a___
blybi8r  LydianBT-BoldItalic                    0844a___
blyr8r   LydianBT-Roman                         0841a___
blyri8r  LydianBT-Italic                        0842a___
blyrw8r  LydianCursiveBT-Regular                0845a___
bm2c8r   Geometric415BT-BlackA                  0524a___
bm2ci8r  Geometric415BT-BlackItalicA            0525a___
bm2l8r   Geometric415BT-LiteA                   0520a___
bm2li8r  Geometric415BT-LiteItalicA             0521a___
bm2m8r   Geometric415BT-MediumA                 0522a___
bm2mi8r  Geometric415BT-MediumItalicA           0523a___
bm3b8r   MiraraeBT-Bold                         0831a___
bm3r8r   MiraraeBT-Roman                        0830a___
bm4r8r   MisterEarlBT-Regular                   1221a___
bm5b8r   MurrayHillBT-Bold                      0886a___
bmar8r   Freehand521BT-RegularC                 1018a___
bmlb8r   ZapfElliptical711BT-Bold               0053a___
bmlbi8r  ZapfElliptical711BT-BoldItalic         0054a___
bmlr8r   ZapfElliptical711BT-Roman              0051a___
bmlri8r  ZapfElliptical711BT-Italic             0052a___
bmor8r   Modern20BT-RomanB                      1168a___
bmori8r  Modern20BT-ItalicB                     1169a___
bmpb8r   GeometricSlab703BT-Bold                0310a___
bmpb8rc  GeometricSlab703BT-BoldCond            1020a___

bmpbi8r  GeometricSlab703BT-BoldItalic          0311a___
bmpl8r   GeometricSlab703BT-Light               0306a___
bmpli8r  GeometricSlab703BT-LightItalic         0307a___
bmpm8r   GeometricSlab703BT-Medium              0308a___
bmpm8rc  GeometricSlab703BT-MediumCond          1019a___
bmpmi8r  GeometricSlab703BT-MediumItalic        0309a___
bmpx8r   GeometricSlab703BT-XtraBold            0312a___
bmpx8rc  GeometricSlab703BT-XtraBoldCond        1021a___
bmpxi8r  GeometricSlab703BT-XtraBoldItal        0313a___
bmsr8r   Staccato222BT-Regular                  0610a___
bmzr8r   AmazoneBT-Regular                      1040a___
bn0r8r   NormandeBT-Roman                       1011a___
bn0ri8r  NormandeBT-Italic                      1012a___
bnlc8r   Informal011BT-Black                    1116a___
bnlr8r   Informal011BT-Roman                    1115a___
bnur8r   NuptialBT-Regular                      0981a___
bnzb8rc  Geometric706BT-BoldCondensedB          0662a___
bnzc8r   Geometric706BT-BlackB                  0661a___
bnzc8rc  Geometric706BT-BlackCondensedB         0663a___
bnzm8r   Geometric706BT-MediumB                 0660a___
bo0r8r   OldDreadfulNo7BT-Regular               7009a___
bo1b8r   OrandaBT-Bold                          1206a___
bo1bi8r  OrandaBT-BoldItalic                    1207a___
bo1brc8r OrandaBT-BoldCondensed                 1209a___
bo1r8r   OrandaBT-Roman                         1204a___
bo1ri8r  OrandaBT-Italic                        1205a___
bo1rrc8r OrandaBT-RomanCondensed                1208a___
bo2r8r   OrbitBbyBT-Regular                     0213a___
boar8r   OCRAbyBT-Regular                       0395x___
bobr8r   OCRB10PitchBT-Regular                  0646x___
boer8r   English157BT-Regular                   0840a___
boir8r   FormalScript421BT-Regular              0609a___
bopb8r   ZapfHumanist601BT-Bold                 0017a___
bopbi8r  ZapfHumanist601BT-BoldItalic           0018a___
bopd8r   ZapfHumanist601BT-Demi                 0134a___
bopdi8r  ZapfHumanist601BT-DemiItalic           0135a___
bopr8r   ZapfHumanist601BT-Roman                0015a___
bopri8r  ZapfHumanist601BT-Italic               0016a___
bopu8r   ZapfHumanist601BT-Ultra                0136a___
bopui8r  ZapfHumanist601BT-UltraItalic          0137a___
borr108r OratorBT-TenPitch                      0755a___
borr158r OratorBT-FifteenPitch                  0758a___
botb8r   Ribbon131BT-Bold                       0208a___
botr8r   Ribbon131BT-Regular                    0207a___
boxr8r   OnyxBT-Regular                         0519a___
bozr8r   OzHandicraftBT-Roman                   1178a___
bp3r8r   Brush445BT-Regular                     0990a___
bp4r8r   PioneerITCbyBT-Regular                 0214a___

bp5r8r   PiranesiItalicBT-Regular               0592a___
bpar8r   ParkAvenueBT-Regular                   0362a___
bpbr8r   PTBarnumBT-Regular                     0720a___
bpfr8r   Decorated035BT-Regular                 0605a___
bpgb8r   Exotic350BT-Bold                       0628a___
bpgd8r   Exotic350BT-DemiBold                   0627a___
bpgl8r   Exotic350BT-Light                      0626a___
bpi001   NewspaperPiBT-Regular                  9836x___
bpi002   SymbolProportionalBT-Regular           9830x___
bpi003   UniversalMath1BT-Regular               9014x___
bpi004   CommercialPiBT-Regular                 9025x___
bplb8r   ZapfCalligraphic801BT-Bold             0025a___
bplbi8r  ZapfCalligraphic801BT-BoldItal         0026a___
bplr8r   ZapfCalligraphic801BT-Roman            0023a___
bplri8r  ZapfCalligraphic801BT-Italic           0024a___
bpnb8r   Aldine721BT-Bold                       0114a___
bpnbi8r  Aldine721BT-BoldItalic                 0115a___
bpnl8r   Aldine721BT-Light                      0804a___
bpnli8r  Aldine721BT-LightItalic                0805a___
bpnr8r   Aldine721BT-Roman                      0112a___
bpnri8r  Aldine721BT-Italic                     0113a___
bppb8r   Lapidary333BT-Bold                     0717a___
bppbi8r  Lapidary333BT-BoldItalic               0718a___
bppc8r   Lapidary333BT-Black                    1052a___
bppr8r   Lapidary333BT-Roman                    0715a___
bppri8r  Lapidary333BT-Italic                   0716a___
bpsr8r   ParisianBT-Regular                     1064a___
bpyr8r   PlaybillBT-Regular                     0867a___
br0b8r   RomanaBT-Bold                          0897a___
br0r8r   RomanaBT-Roman                         0896a___
brab8r   RaleighBT-Bold                         1081a___
brad8r   RaleighBT-DemiBold                     1080a___
bral8r   RaleighBT-Light                        1077a___
bram8r   RaleighBT-Medium                       1079a___
brar8r   RaleighBT-Roman                        1078a___
brax8r   RaleighBT-ExtraBold                    1082a___
brmb8r   CarminaBT-Bold                         0723a___
brmbi8r  CarminaBT-BoldItalic                   0724a___
brmc8r   CarminaBT-Black                        0668a___
brmci8r  CarminaBT-BlackItalic                  0719a___
brml8r   CarminaBT-Light                        0667a___
brmli8r  CarminaBT-LightItalic                  0624a___
brmm8r   CarminaBT-Medium                       0721a___
brmmi8r  CarminaBT-MediumItalic                 0722a___
brsb8r   ArrusBT-Bold                           1147a___
brsbi8r  ArrusBT-BoldItalic                     1148a___
brsc8r   ArrusBT-Black                          1149a___
brsci8r  ArrusBT-BlackItalic                    1150a___

brsr8r   ArrusBT-Roman                          1145a___
brsri8r  ArrusBT-Italic                         1146a___
brvr8r   RevueBT-Regular                        0895a___
brwb8r   GeometricSlab712BT-BoldA               0487a___
brwl8r   GeometricSlab712BT-LightA              0483a___
brwli8r  GeometricSlab712BT-LightItalicA        0484a___
brwm8r   GeometricSlab712BT-MediumA             0485a___
brwmi8r  GeometricSlab712BT-MediumItalA         0486a___
brwx8r   GeometricSlab712BT-ExtraBoldA          0488a___
bs3b8r   SeagullBT-Bold                         0819a___
bs3h8r   SeagullBT-Heavy                                0820a___
bs3l8r   SeagullBT-Light                                0817a___
bs3m8r   SeagullBT-Medium                       0818a___
bs4r8r   ShotgunBT-Regular                      0167a___
bs4rl8r  ShotgunBlanksBT-Regular                0234a___
bs5b8r   SnellBT-Bold                           0197a___
bs5c8r   SnellBT-Black                          0198a___
bs5r8r   SnellBT-Regular                        0196a___
bs6r8r   StuyvesantBT-Regular                   0980a___
bsbb8r   ClassicalGaramondBT-Bold               0279a___
bsbbi8r  ClassicalGaramondBT-BoldItalic         0972a___
bsbr8r   ClassicalGaramondBT-Roman              0277a___
bsbri8r  ClassicalGaramondBT-Italic             0278a___
bslr8r   StencilBT-Regular                      0191a___
bt0r8r   TangoBT-Regular                        1151a___
bt1r8r   ThunderbirdBT-Regular                  0606a___
bt2r8r   TypoUprightBT-Regular                  0939a___
btmb8r   Dutch801BT-Bold                        0013a___
btmbi8r  Dutch801BT-BoldItalic                  0014a___
btmr8r   Dutch801BT-Roman                       0011a___
btmrd8r  Dutch801BT-RomanHeadline               1035a___
btmri8r  Dutch801BT-Italic                      0012a___
btms8r   Dutch801BT-SemiBold                    0834a___
btmsi8r  Dutch801BT-SemiBoldItalic              0835a___
btmx8r   Dutch801BT-ExtraBold                   0549a___
btmxi8r  Dutch801BT-ExtraBoldItalic             0971a___
btor8r   Industrial736BT-Roman                  1047a___
btori8r  Industrial736BT-Italic                 1048a___
btvb8r   Kuenstler480BT-Bold                    0763a___
btvbi8r  Kuenstler480BT-BoldItalic              0764a___
btvc8r   Kuenstler480BT-Black                   1026a___
btvr8r   Kuenstler480BT-Roman                   0761a___
btvri8r  Kuenstler480BT-Italic                  0762a___
bumr8r   UmbraBT-Regular                        1074a___
bunb8r   ZurichBT-Bold                          0180a___
bunb8rc  ZurichBT-BoldCondensed                 0300a___
bunb8rq  ZurichBT-BoldExtraCondensed            0960a___
bunb8rx  ZurichBT-BoldExtended                  0261a___

bunbi8r  ZurichBT-BoldItalic                    0181a___
bunbi8rc ZurichBT-BoldCondensedItalic           0301a___
bunc8r   ZurichBT-Black                         0182a___
bunc8rx  ZurichBT-BlackExtended                 0262a___
bunci8r  ZurichBT-BlackItalic                   0183a___
bunl8r   ZurichBT-Light                         0176a___
bunl8rc  ZurichBT-LightCondensed                0296a___
bunl8rq  ZurichBT-LightExtraCondensed           0958a___
bunli8r  ZurichBT-LightItalic                   0177a___
bunli8rc ZurichBT-LightCondensedItalic          0297a___
bunr8r   ZurichBT-Roman                         0178a___
bunr8rc  ZurichBT-RomanCondensed                0298a___
bunr8rq  ZurichBT-ExtraCondensed                0959a___
bunr8rx  ZurichBT-RomanExtended                 0260a___
bunri8r  ZurichBT-Italic                        0179a___
bunri8rc ZurichBT-ItalicCondensed               0299a___
bunu8rx  ZurichBT-UltraBlackExtended            0263a___
bunx8r   ZurichBT-ExtraBlack                    0970a___
burb8r   Venetian301BT-Bold                     1099a___
burbi8r  Venetian301BT-BoldItalic               1100a___
burd8r   Venetian301BT-Demi                     1097a___
burdi8r  Venetian301BT-DemiItalic               1098a___
burr8r   Venetian301BT-Roman                    1095a___
burri8r  Venetian301BT-Italic                   1096a___
buyb8r   UniversityRomanBT-Bold                 0317a___
buyr8r   UniversityRomanBT-Regular              0316a___
bv0r8r   VinetaBT-Regular                       0215a___
bvrr8r   VAGRoundedBT-Regular                   0756a___
bw0r8r   WeddingTextBT-Regular                  0985a___
bwnl8r   WindsorBT-Light                        0909a___
bwnl8rc  WindsorBT-LightCondensed               0326a___
bwnr8r   WindsorBT-Roman                        0122a___
bwnr8rx  WindsorBT-Elongated                    0327a___
bwnrl8r  WindsorBT-Outline                      0988a___
zbshals  ShelleyAllegroBT-Regular               0594a___
zbshans  ShelleyAndanteBT-Regular               0593a___
zbshvos  ShelleyVolanteBT-Regular               0595a___

A.5 DTC fonts

The ‘dtc.map’ file has abbreviations for DTC (Digital Typeface Corporation) fonts.

dhvc8rc  SansBlackCondensed-DTC
dhvco8rc SansBlackCondensed-Oblique-DTC
dhvm8rx  SansExtended-Medium-DTC
dhvr8rx  SansExtended-Regular-DTC
dhvro8rc SansExtended-RegularOblique-DTC
dhvx8rc  SansExtraBlackCondensed-DTC
dhvl8rx  SansLightExtended-DTC

dhvh8rc  SansThinCondensed-DTC
dhvho8rc SansThinCondensed-Oblique-DTC
dnbb8r   NewBaskerville-Bold-DTC
dnbbi8r  NewBaskerville-BoldItalic-DTC
dnbri8r  NewBaskerville-RomanItalic-DTC
dnbr8r   NewBaskerville-Roman-DTC
dopr8r   Optimum-Roman-DTC
dopb8r   Optimum-Bold-DTC
dopbo8r  Optimum-BoldItalic-DTC
doro8r   Optimum-RomanItalic-DTC
dpll8r   Palton-Light-DTC

A.6 Linotype fonts

The ‘linotype.map’ file has abbreviations for Linotype fonts.

lenb8r   EngraversLH-BoldFace
lpi001   OlympicLH-Pi
ltgb8rx  TradeGothicLH-BoldExtended
ltgr8rx  TradeGothicLH-Extended
lvcc8r   VectoraLH-Black
lvcci8r  VectoraLH-BlackItalic
lvcb8r   VectoraLH-Bold
lvcbi8r  VectoraLH-BoldItalic
lvcri8r  VectoraLH-Italic
lvcl8r   VectoraLH-Light
lvcli8r  VectoraLH-LightItalic
lvcr8r   VectoraLH-Roman
lpi002   WarningLH-Pi
lpi003   AudioLH-Pi
lcfcs8r  ClearfaceGothicLH-Black
lcfbs8r  ClearfaceGothicLH-Bold
lcfls8r  ClearfaceGothicLH-Light
lcfms8r  ClearfaceGothicLH-Medium
lcfrs8r  ClearfaceGothicLH-Roman
lpi004   EECLH-Pi
lffb8r   FairfieldLH-Bold
lffbi8r  FairfieldLH-BoldItalic
lffbc8r  FairfieldLH-BoldSC
lffbd8r  FairfieldLH-CaptionBold
lffhd8r  FairfieldLH-CaptionHeavy
lffld8r  FairfieldLH-CaptionLight
lffmd8r  FairfieldLH-CaptionMedium
lffh8r   FairfieldLH-Heavy
lffhi8r  FairfieldLH-HeavyItalic
lffhc8r  FairfieldLH-HeavySC
lffl8r   FairfieldLH-Light
lffli8r  FairfieldLH-LightItalic
lfflc8r  FairfieldLH-LightSC

lffm8r   FairfieldLH-Medium
lffmi8r  FairfieldLH-MediumItalic
lffmc8r  FairfieldLH-MediumSC
lffbiw7d FairfieldLH-SwBoldItalicOsF
lffhiw7d FairfieldLH-SwHeavyItalicOsF
lffliw7d FairfieldLH-SwLightItalicOsF
lffmiw7d FairfieldLH-SwMediumItalicOsF

A.7 Monotype fonts

The ‘monotype.map’ file has abbreviations for Monotype fonts.

ma1b8r   ArialMT-Bold
ma1b8rn  ArialNarrowMT-Bold MArial-NarrowBold
ma1b9s   ArialSFMT-Bold
ma1bi8r  ArialMT-BoldItalic
ma1bi8rn ArialNarrowMT-BoldItalic MArial-NarrowBoldItalic
ma1bi9s  ArialSFMT-BoldItalic
ma1bot8r ArialMonospacedMT-BoldObliq
ma1br8rc ArialMT-CondensedBold
ma1bt8r  ArialMonospacedMT-Bold
ma1c8r   ArialMT-Black
ma1ci8r  ArialMT-BlackItalic
ma1l8r   ArialMT-Light
ma1l8rc  ArialMT-CondensedLight
ma1li8r  ArialMT-LightItalic
ma1m8r   ArialMT-Medium
ma1mi8r  ArialMT-MediumItalic
ma1r8r   ArialMT
ma1r8rc  ArialMT-Condensed
ma1r8rn  ArialNarrowMT MArial-Narrow
ma1r9s   ArialSFMT
ma1ri8r  ArialMT-Italic
ma1ri9s  ArialSFMT-Italic
ma1rin8r ArialNarrowMT-Italic MArial-NarrowItalic
ma1ro8r  ArialMonospacedMT-Oblique
ma1rt8r  ArialMonospacedMT
ma1x8r   ArialMT-ExtraBold
ma1x8rc  ArialMT-CondensedExtraBold
ma1xi8r  ArialMT-ExtraBoldItalic
ma2b8r   AmasisMT-Bold
ma2bi8r  AmasisMT-BoldItalic
ma2c8r   AmasisMT-Black
ma2ci8r  AmasisMT-BlackItalic
ma2l8r   AmasisMT-Light
ma2li8r  AmasisMT-LightItalic
ma2m8r   AmasisMT-Medium
ma2mi8r  AmasisMT-MediumItalic
ma2r8r   AmasisMT

ma2ri8r  AmasisMT-Italic
magd8r   MAvantGardeGothicPS-Demi
magdo8r  MAvantGardeGothicPS-DemiObl
magk8r   MAvantGardeGothicPS-Book
magko8r  MAvantGardeGothicPS-BookObl
mahr8r   AshleyScriptMT
maib8r   AbadiMT-Bold MAbadi-Bold
maib8rc  AbadiMT-CondensedBold
maibi8r  AbadiMT-BoldItalic MAbadi-BoldItalic
maij8r   AbadiMT-ExtraLight MAbadi-ExtraLight
maiji8r  AbadiMT-ExtraLightItalic MAbadi-ExtraLightItalic
mail8r   AbadiMT-Light MAbadi-Light
mail8rc  AbadiMT-CondensedLight
maili8r  AbadiMT-LightItalic MAbadi-LightItalic
mair8r   AbadiMT MAbadi
mair8rc  AbadiMT-Condensed
mairi8r  AbadiMT-Italic MAbadi-Italic
maix8r   AbadiMT-ExtraBold MAbadi-ExtraBold
maix8rc  AbadiMT-CondensedExtraBold
maixi8r  AbadiMT-ExtraBoldItalic MAbadi-ExtraBoldItalic
mall8r   MAlbertus-Light AlbertusMT-Light
malr8r   MAlbertus AlbertusMT
malri8r  MAlbertus-Italic AlbertusMT-Italic
mapr8r   ApolloMT MApollo
maprc8r  ApolloMT-SC
mapri7d  ApolloMT-ItalicOsF
mapri8r  ApolloMT-Italic MApollo-Italic
mapri8x  ApolloMT-ItalicExpert
maps7d   ApolloMT-SemiBoldOsF
maps8r   ApolloMT-SemiBold MApollo-SemiBold
maps8x   ApolloMT-SemiBoldExpert
maqb8r   BookAntiqua-Bold
maqbi8r  BookAntiqua-BoldItalic
maqr8r   BookAntiqua
maqri8r  BookAntiqua-Italic
mb1ri8r  BladoMT-Italic MBlado-Italic
mb1ri8x  BladoExpertMT-Italic
mb2r8r   BinnyOldStyleMT
mb2ri8r  BinnyOldStyleMT-Italic
mb3r8r   MBraggadocio
mbbb7a   Bembo-AltBold
mbbb8r   Bembo-Bold
mbbb8x   BemboExpert-Bold MBemboExpert-Bold
mbbbi7a  Bembo-AltBoldItalic
mbbbi8r  Bembo-BoldItalic
mbbbi8x  BemboExpert-BoldItalic MBemboExpert-BoldItalic
mbbbin8r BemboMT-SchlbkBoldItalic
mbbbn8r  BemboMT-SchlbkBold

mbbr7a   Bembo-Alt
mbbr8r   Bembo
mbbr8x   BemboExpert MBemboExpert
mbbrd8r  BemboTitlingMT
mbbrdi8r BemboTitlingMT-Italic
mbbri7a  Bembo-AltItalic
mbbri8r  Bembo-Italic
mbbri8x  BemboExpert-Italic MBemboExpert-Italic
mbbrin8r BemboMT-SchlbkItalic
mbbrn8r  BemboMT-Schlbk
mbbs7a   Bembo-AltSemiBold
mbbs8r   Bembo-Semibold
mbbs8x   BemboExpert-SemiBold MBemboExpert-SemiBold
mbbsi7a  Bembo-AltSemiBoldItalic
mbbsi8r  Bembo-SemiboldItalic
mbbsi8x  BemboExpert-SemiBoldItalic MBemboExpert-SemiBoldItalic
mbbx7a   Bembo-AltExtraBold
mbbx8r   Bembo-ExtraBold
mbbx8x   BemboExpert-ExtraBold MBemboExpert-ExtraBold
mbbxi8r  Bembo-ExtraBoldItalic
mbbxi8x  BemboExpert-ExtraBoldItalic MBemboExpert-ExtraBoldItalic
mbdb8r   BodoniMT-Bold
mbdb8rc  BodoniMT-CondensedBold MBodoni-BoldCondensed
mbdbi8r  BodoniMT-BoldItalic
mbdbi8rc BodoniMT-CondensedBoldItalic MBodoni-BoldCondensedItalic
mbdc8r   BodoniMT-Black MBodoni-Black
mbdci8r  BodoniMT-BlackItalic MBodoni-BlackItalic
mbdk8r   BodoniMT-Book MBodoni-Book
mbdki8r  BodoniMT-BookItalic MBodoni-BookItalic
mbdr8r   BodoniMT
mbdri8r  BodoniMT-Italic
mbdu8r   BodoniMT-UltraBold MBodoni-UltraBold
mbdui8r  BodoniMT-UltraBoldItalic MBodoni-UltraBoldItalic
mbeb8r   BellMT-Bold MBell-Bold
mbeb8x   BellExpertMT-Bold
mbebi8r  BellMT-BoldItalic
mbebi8x  BellExpertMT-BoldItalic
mber7a   BellMT-Alt
mber8r   BellMT MBell
mber8x   BellExpertMT
mbera8x  BellExpertMT-Alt
mberi7a  BellMT-AltItalic
mberi8r  BellMT-Italic MBell-Italic
mberi8x  BellExpertMT-Italic
mbes7a   BellMT-AltSemiBold
mbes8r   BellMT-SemiBold MBell-SemiBold
mbes8x   BellExpertMT-SemiBold
mbesa8x  BellExpertMT-AltSemiBold

mbesi7a  BellMT-AltSemiBoldItalic
mbesi8r  BellMT-SemiBoldItalic
mbesi8x  BellExpertMT-SemiBoldItalic
mbfr8r   BiffoMT
mbkd8r   MBookmanPS-Demi
mbkdi8r  MBookmanPS-DemiItalic
mbkl8r   MBookmanPS-Light
mbkli8r  MBookmanPS-LightItalic
mbmb7a   BulmerMT-BoldAlt
mbmb8r   BulmerMT-Bold
mbmb8x   BulmerMT-BoldExpert
mbmbd7a  BulmerMT-BoldDisplayAlt
mbmbd8r  BulmerMT-BoldDisplay
mbmbdi7a BulmerMT-BoldItalDisplayAlt
mbmbdi8r BulmerMT-BoldItalicDisplay
mbmbi7a  BulmerMT-BoldItalicAlt
mbmbi8r  BulmerMT-BoldItalic
mbmbi8x  BulmerMT-BoldItalicExpert
mbmr7a   BulmerMT-RegularAlt
mbmr8r   BulmerMT-Regular
mbmr8x   BulmerMT-RegularExpert
mbmrc8r  BulmerMT-RegularSC
mbmrd7a  BulmerMT-RegularDisplayAlt
mbmrd8r  BulmerMT-RegularDisplay
mbmrdi7a BulmerMT-ItalicDisplayAlt
mbmrdi8r BulmerMT-ItalicDisplay
mbmri7a  BulmerMT-ItalicAlt
mbmri8r  BulmerMT-Italic
mbmri8x  BulmerMT-ItalicExpert
mbms7a   BulmerMT-SemiBoldAlt
mbms8r   BulmerMT-SemiBold
mbms8x   BulmerMT-SemiBoldExpert
mbmsi7a  BulmerMT-SemiBoldItalicAlt
mbmsi8r  BulmerMT-SemiBoldItalic
mbmsi8x  BulmerMT-SemiBoldItalExpert
mbmss8r  BulmerMT-SemiBoldSC
mbnr8rc  BernardMT-Condensed
mbtm     Boton-Medium
mbtmi    Boton-MediumItalic
mbtr     Boton-Regular
mbtri    Boton-Italic
mburw8r  BrushScriptMT
mbvb8r   BaskervilleMT-Bold
mbvb8x   BskvillExpMT-Bold MBaskervilleExpert-Bold
mbvbi8r  BaskervilleMT-BoldItalic
mbvbi8x  BskvillExpMT-BoldItalic MBaskervilleExpert-BoldItalic
mbvr8r   BaskervilleMT
mbvr8x   BskvillExpMT MBaskervilleExpert

mbvri8r  BaskervilleMT-Italic
mbvri8x  BskvillExpMT-Italic MBaskervilleExpert-Italic
mbvs8r   BaskervilleMT-SemiBold
mbvs8x   BskvillExpMT-SemiBold MBaskervilleExpert-SemiBold
mbvsi8r  BaskervilleMT-SemiBoldItalic
mbvsi8x  BskvillExpMT-SemiBoldItalic MBaskervilleExpt-SemiBoldIta
mc1b8r   CalistoMT-Bold
mc1bi8r  CalistoMT-BoldItalic
mc1r8r   CalistoMT
mc1ri8r  CalistoMT-Italic
mc6b8r   ClarionMT-Bold MClarion-Bold
mc6r8r   ClarionMT MClarion
mc6ri8r  ClarionMT-Italic MClarion-Italic
mc7b8r   CompactaMT-Bold
mc8b8r   CalvertMT-Bold
mc8l8r   CalvertMT-Light
mc8m8r   CalvertMT-Medium
mc8r8r   CalvertMT
mcard8r  CaslonTitlingMT
mcdr8r   ClarendonMT MtClarendon
mcfb8r   ClearfaceMT-Bold
mcfbs8r  ClearfaceGothicMT-Bold
mcfds8r  ClearfaceGothicMT-DemiBold
mcnb8r   CenturyMT-Bold MCentury-Bold
mcnbi8r  CenturyMT-BoldItalic MCentury-BoldItalic
mcnr8rx  CenturyMT-Expanded MCentury-Expanded
mcnri8rx CenturyMT-ExpandedItalic MCentury-ExpandedItalic
mcr8r12  CourierTwelveMT
mcrb8r   CourierPS-Bold MCourierPS-Bold
mcrbi8r  CourierPS-BoldItalic MCourierPS-BoldItalic
mcrr8r   CourierPS MCourierPS
mcrri8r  CourierPS-Italic MCourierPS-Italic
mcsb8r   CenturySchbookMT-Bold MCenturySchoolbook-Bold
mcsbi8r  CenturySchbookMT-BoldItalic MCenturySchoolbook-BoldItalic
mcsr8r   CenturySchbookMT MCenturySchoolbook
mcsri8r  MCenturySchoolbook-Italic
mcub8r   CenturyOSMT-Bold MCenturyOldStyle-Bold
mcubi8r  CenturyOSMT-BoldItalic MCenturyOldStyle-BoldItalic
mcur8r   CenturyOSMT MCenturyOldStyle
mcuri8r  CenturyOSMT-Italic MCenturyOldStyle-Italic
mdmb8r   DelimaMT-Bold
mdmbi8r  DelimaMT-BoldItalic
mdml8r   DelimaMT-Light
mdmli8r  DelimaMT-LightItalic
mdmr8r   DelimaMT
mdmri8r  DelimaMT-Italic
mdms8r   DelimaMT-SemiBold
mdmsi8r  DelimaMT-SemiBoldItalic

mdmx8r   DelimaMT-ExtraBold
mdmxi8r  DelimaMT-ExtraBoldItalic
mdsr8r   DorchesterScriptMT
mdtb7a   DanteMT-AltBold
mdtb8r   DanteMT-Bold
mdtb8x   DanteExpertMT-Bold
mdtbi7a  DanteMT-AltBoldItalic
mdtbi8r  DanteMT-BoldItalic
mdtbi8x  DanteExpertMT-BoldItalic
mdtm7a   DanteMT-AltMedium
mdtm8r   DanteMT-Medium
mdtm8x   DanteExpertMT-Medium
mdtmi7a  DanteMT-AltMediumItalic
mdtmi8r  DanteMT-MediumItalic
mdtmi8x  DanteExpertMT-MediumItalic
mdtr7a   DanteMT-Alt
mdtr8r   DanteMT
mdtr8x   DanteExpertMT
mdtrd8r  DanteTitlingMT
mdtri7a  DanteMT-AltItalic
mdtri8r  DanteMT-Italic
mdtri8x  DanteExpertMT-Italic
mebm8r   BaskervilleBE-Medium
mebmi8r  BaskervilleBE-MediumItalic
mebr8r   BaskervilleBE-Regular
mebri8r  BaskervilleBE-Italic
meer8r   EgyptianMT-Extended MEgyptianExtended
melb8r   EllingtonMT-Bold MEllington-Bold
melbi8r  EllingtonMT-BoldItalic MEllington-BoldItalic
mell8r   EllingtonMT-Light MEllington-Light
melli8r  EllingtonMT-LightItalic MEllington-LightItalic
melr8r   EllingtonMT MEllington
melri8r  EllingtonMT-Italic MEllington-Italic
melx8r   EllingtonMT-ExtraBold MEllington-ExtraBold
melxi8r  EllingtonMT-ExtraBoldItalic MEllington-ExtraBoldItalic
menb8r   EngraversMT-Bold
menr8r   EngraversMT
menrf8r  EngraversOldEnglishMT
metr8r   EhrhardtMT
metr8x   EhrhardtExpMT MEhrhardtExpt
metri8r  EhrhardtMT-Italic
metri8x  EhrhardtExpMT-Italic MEhrhardtExpt-Italic
mets8r   EhrhardtMT-SemiBold
mets8x   EhrhardtExpMT-SemiBold MEhrhardtExpt-SemiBold
metsi8r  EhrhardtMT-SemiBoldItalic
metsi8x  EhrhardtExpMT-SemiBoldItal MEhrhardtExpt-SemiBoldItalic
mf1r7a   FournierMT-Alt
mf1r8r   FournierMT

mf1r8x   FournierExpertMT
mf1rd8r  FournierMT-TallCaps
mf1rdi8r FournierMT-TallCapsItalic
mf1ri7a  FournierMT-AltItalic
mf1ri8r  FournierMT-Italic
mf1ri8x  FournierExpertMT-Italic
mf1rp8r  FournierOrnamentsMT
mf2r8rc  FacadeMT-Condensed
mf3rd8r  FestivalTitlingMT
mf4b8r   FootlightMT-Bold
mf4bi8r  FootlightMT-BoldItalic
mf4l8r   FootlightMT-Light
mf4li8r  FootlightMT-LightItalic
mf4r8r   FootlightMT
mf4ri8r  FootlightMT-Italic
mf4x8r   FootlightMT-ExtraBold
mf4xi8r  FootlightMT-ExtraBoldItalic
mfer8r   ForteMT MForte
mfgr8rq  FrankGothMT-ExtraCondensed
mfir8r   FigaroMT MFigaro
mfsr8r   FalstaffMT MFalstaff
mfsrd8r  FalstaffFestivalMT
mfxr8r   FelixTitlingMT
mg2rw    Giddyup
mg4r8r   MGoudyModern
mg4ri8r  MGoudyModern-Italic
mg5r8r   GoudyCatalogueMT MGoudyCatalogue
mg5ri8r  GoudyCatalogueMT-Italic MGoudyCatalogue-Italic
mgmb8r   GaramondMT-Bold
mgmb8x   GaramondExpertMT-Bold MGaramondExpert-Bold
mgmbi8r  GaramondMT-BoldItalic
mgmbi8x  GaramondExpertMT-BoldItalic MGaramondExpert-BoldItalic
mgmr8r   GaramondMT
mgmr8x   GaramondExpertMT MGaramondExpert
mgmrd8r  GaramondMT-SwashCaps
mgmri7a  GaramondMT-AlternativeItalic
mgmri8r  GaramondMT-Italic
mgmri8x  GaramondExpertMT-Italic MGaramondExpert-Italic
mgmrw8r  GaramondMT-Swash
mgor8r   GoudyOldStyleMT MGoudyOldStyle
mgori8r  GoudyOldStyleMT-Italic MGoudyOldStyle-Italic
mgqb8r   GrotesqueMT-Bold MGrotesque-Bold
mgqb8rx  GrotesqueMT-BoldExtended
mgqbd8rc GrotesqueDisplayMT-BoldCond
mgqbd8rx GrotesqueDisplayMT-BoldExt
mgqc8r   GrotesqueMT-Black MGrotesque-Black
mgql8r   GrotesqueMT-Light MGrotesque-Light
mgql8rc  GrotesqueMT-LightCondensed MGrotesque-LightCondensed

mgqli8r  GrotesqueMT-LightItalic MGrotesque-LightItalic
mgqr8r   GrotesqueMT MGrotesque
mgqr8rc  GrotesqueMT-Condensed MGrotesque-Condensed
mgqr8rq  GrotesqueMT-ExtraCondensed MGrotesque-ExtraCondensed
mgqri8r  GrotesqueMT-Italic MGrotesque-Italic
mgrb8r   GloucesMT-Bold MGloucester-Bold
mgrb8rc  GloucesMT-CondensedBold MGloucester-BoldCondensed
mgrb8ro  GloucesMT-ExtraCondensedBold MGloucester-BoldExtraCondensed
mgrb8rx  GloucesMT-ExtendedBold MGloucester-BoldExtended
mgrr8r   MGloucester-OldStyle
mgs8rc   GillSans-Condensed
mgsb7a   GillAltOneMT-Bold
mgsb8r   GillSans-Bold
mgsb8rc  GillSans-BoldCondensed
mgsb8rq  GillSansMT-ExCondensedBold
mgsbd8r  GillSansDisplayMT-Bold
mgsbd8rc GillSansDisplayMT-CondBold
mgsbi7a  GillAltOneMT-BoldItalic
mgsbi8r  GillSans-BoldItalic
mgsbin8r GillSansMT-SchlbkBoldItalic
mgsbn8r  GillSansMT-SchlbkBold
mgsh7a   GillAltOneMT-Heavy
mgsh8r   GillSans-Heavy
mgshi7a  GillAltOneMT-HeavyItalic
mgshi8r  GillSans-HeavyItalic
mgsk7a   GillAltOneMT-Book
mgsk8r   GillSans-Book
mgski7a  GillAltOneMT-BookItalic
mgski8r  GillSans-BookItalic
mgsl7a   GillAltOneMT-Light
mgsl8r   GillSans-Light
mgsli7a  GillAltOneMT-LightItalic
mgsli8r  GillSans-LightItalic
mgsr7a   GillAltOneMT
mgsr8r   GillSans
mgsr8rc  GillSans-Condensed
mgsrh8r  GillSansMT-Shadow
mgsri7a  GillAltOneMT-Italic
mgsri8r  GillSans-Italic
mgsrin8r GillSansMT-SchlbkItalic
mgsrl8r  GillSansShadowedMT-Light
mgsrn8r  GillSansMT-Schlbk
mgsu8r   GillSans-UltraBold
mgsu8rc  GillSans-UltraBoldCondensed
mgsx8r   GillSans-ExtraBold
mgsxd8r  GillSansDisplayMT-ExtraBold
mgtr7a   GoudyTextMT-Alternate
mgtr7c   GoudyTextMT-Dfr

mgtr8r   MGoudyText GoudyTextMT
mgtrd8r  GoudyTextMT-LombardicCapitals
mgyb8r   GoudyMT-Bold MGoudyBold
mgybi8r  GoudyMT-BoldItalic
mgyrp8r  GoudySorts
mgyx8r   GoudyMT-ExtraBold MGoudyExtraBold
mhdr8r   HeadlineMT-Bold MHeadlineBold
mhob8r   HorleyOSMT-Bold MHorleyOldStyle-Bold
mhobi8r  HorleyOSMT-BoldItalic MHorleyOldStyle-BoldItalic
mhol8r   HorleyOSMT-Light MHorleyOldStyle-Light
mholi8r  HorleyOSMT-LightItalic MHorleyOldStyle-LightItalic
mhor8r   HorleyOSMT MHorleyOldStyle
mhori8r  HorleyOSMT-Italic MHorleyOldStyle-Italic
mhos8r   HorleyOSMT-SemiBold MHorleyOldStyle-SemiBold
mhosi8r  HorleyOSMT-SemiBoldItalic MHorleyOldStyle-SemiBolItalic
mhrb8r   ArialRoundedMT-Bold
mhrl8r   ArialRoundedMT-Light
mhrr8r   ArialRoundedMT
mhrx8r   ArialRoundedMT-ExtraBold
mi0r8r   CommercialScriptMT
micr8r   ImpactMT
mifb8r   InflexMT-Bold MInflex-Bold
miib8r   ImprintMT-Bold MImprint-Bold
miib8x   ImprintExpertMT-Bold
miibi8r  ImprintMT-BoldItalic MImprint-BoldItalic
miibi8x  ImprintExpertMT-BoldItalic
miir8r   ImprintMT MImprint
miir8x   ImprintExpertMT
miirh8r  ImprintMT-Shadow
miirhi8r ImprintMT-ShadowItalic
miiri8r  ImprintMT-Italic MImprint-Italic
miiri8x  ImprintExpertMT-Italic
miirp    ImprintExtras
miob8r   IonicMT-Bold MIonic-Bold
mior8r   IonicMT MIonic
miori8r  IonicMT-Italic MIonic-Italic
misb8r   ItalianOldStyleMT-Bold MItalianOldStyle-Bold
misbi8r  ItalianOldStyleMT-BoldItalic MItalianOldStyle-BoldItalic
misr8r   ItalianOldStyleMT MItalianOldStyle
misri8r  ItalianOldStyleMT-Italic MItalianOldStyle-Italic
mjnb8r   JansonMT-Bold MJanson-Bold
mjnb8x   JansonExpertMT-Bold
mjnbi8r  JansonMT-BoldItalic MJanson-BoldItalic
mjnbi8x  JansonExpertMT-BoldItalic
mjnr8r   JansonMT MJanson
mjnr8x   JansonExpertMT
mjnri8r  JansonMT-Italic MJanson-Italic
mjnri8x  JansonExpertMT-Italic

mjob7d   JoannaMT-BoldOSF
mjob8r   JoannaMT-Bold
mjob8x   JoannaMT-BoldExpert
mjobi7d  JoannaMT-BoldItalicOSF
mjobi8r  JoannaMT-BoldItalic
mjobi8x  JoannaMT-BoldItalicExpert
mjor8r   JoannaMT
mjor8x   JoannaMT-Expert
mjorc7d  JoannaMT-SCOSF
mjori7d  JoannaMT-ItalicOSF
mjori8r  JoannaMT-Italic
mjori8x  JoannaMT-ItalicExpert
mjos8r   JoannaMT-SemiBold
mjos8x   JoannaMT-SemiBoldExpert
mjosi7d  JoannaMT-SemiBoldItalicOSF
mjosi8r  JoannaMT-SemiBoldItalic
mjosi8x  JoannaMT-SemiBoldItaExpert
mjoss7d  JoannaMT-SemiBoldSCOSF
mjox7d   JoannaMT-ExtraBoldOSF
mjox8r   JoannaMT-ExtraBold
mjox8x   JoannaMT-ExtraBoldExpert
mklr8r   KlangMT MKlang
mkn8r    KinoMT
mkob8r   BookmanOldStyle-Bold
mkobi8r  BookmanOldStyle-BoldItalic
mkor8r   BookmanOldStyle
mkori8r  BookmanOldStyle-Italic
mlgb8r   LetterGothicMT-Bold
mlgbo8r  LetterGothicMT-BoldOblique
mlgr8r   LetterGothicMT
mlgro8r  LetterGothicMT-Oblique
mlir8rc  LatinMT-Condensed
mlyr8r   LydianMT
mlyrw8r  LydianCursiveMT
mm1r8r   MonolineScriptMT
mm7r8r   MaturaMT
mm7rw8r  MaturaMT-ScriptCapitals
mm8bl8r  MonotypeOldStyleMT-BoldOut
mmeb8r   MercuriusMT-BoldScript MercuriusScriptMT-Bold
mmob8r   ModernMT-Bold MModern-Bold
mmobi8r  ModernMT-BoldItalic MModern-BoldItalic
mmor8r   ModernMT-Extended MModernExtended
mmor8rc  ModernMT-Condensed
mmor8rw  ModernMT-Wide
mmori8r  ModernMT-CondensedItalic MModernExtended-Italic
mmori8rw ModernMT-WideItalic
mmori8rx ModernMT-ExtendedItalic
mn1p8r   BinnerPosterMT

mn1r8r   BinnerGothicMT
mncb8r   MCenturySchoolbookPS-Bold
mncbi8r  CentSchlbookPS-BoldItalic MCenturySchoolbookPS-BoldIta
mncr8r   CentSchlbookPS MCenturySchoolbookPS
mncri8r  CentSchlbookPS-Italic MCenturySchoolbookPS-Ita
mngb8r   NewsGothicMT-Bold
mngb8rc  NewsGothicMT-CondensedBold
mngr8r   NewsGothicMT
mngr8rc  NewsGothicMT-Condensed
mngri8r  NewsGothicMT-Italic
mnib8r   NimrodMT-Bold
mnibi8r  NimrodMT-BoldItalic
mnir8r   NimrodMT
mniri8r  NimrodMT-Italic
mnkr8r   NeographikMT
mnnb8r   NewClarendonMT-Bold MNewClarendon-Bold
mnnr8r   NewClarendonMT MNewClarendon
mnpb8r   NewsPlantinMT-Bold
mnpbi8r  NewsPlantinMT-BoldItalic
mnpr8r   NewsPlantinMT
mnpri8r  NewsPlantinMT-Italic
mnrr8r   NewBerolinaMT
mnsb8r   TimesNewRomanPS-Bold
mnsb8x   TimesNRPSExpert-Bold
mnsbi8r  TimesNewRomanPS-BoldItalic
mnsbi8x  TimesNRPSExpert-BoldItalic
mnsr8r   TimesNewRomanPS
mnsr8x   TimesNRPSExpert
mnsri8r  TimesNewRomanPS-Italic
mnsri8x  TimesNRPSExpert-Italic
mntb78r  TimesNRSevenMT-Bold
mntb8r   TimesNRMT-Bold
mntb8rc  TimesNewRomanMT-BoldCond TimesNRMT-CondensedBold
mntb8x   TimesNRExpertMT-Bold MTimesNewRomanExpert-Bold
mntb9s   TimesNewRomanSF-Bold
mntbi78r TimesNRSevenMT-BoldItalic
mntbi8r  TimesNRMT-BoldItalic
mntbi8x  TimesNRExpertMT-BoldItalic MTimesNewRomanExpt-BoldItalic
mntbi9s  TimesNewRomanSF-BoldItalic
mntbk8r  MTimesNewRomanGreek-Bold
mntbko8r MTimesNewRomanGreek-BoldInc
mntboz8r MTimesNewRomanCyrillic-BolInc
mntbz8r  MTimesNewRomanCyrillic-Bold
mntm8r   TimesNRMT-Medium MTimesNewRoman-Medium
mntmi8r  TimesNRMT-MediumItalic MTimesNewRoman-MediumItalic
mntr78r  TimesNRSevenMT
mntr8r   TimesNRMT
mntr8rc  TimesNewRomanMT-Cond TimesNRMT-Condensed

mntr8x   TimesNRExpertMT MTimesNewRomanExpert
mntr9s   TimesNewRomanSF
mntri78r TimesNRSevenMT-Italic
mntri8r  TimesNRMT-Italic
mntri8rc TimesNewRomanMT-CondItalic TimesNRMT-CondensedItalic
mntri8x  TimesNRExpertMT-Italic MTimesNewRomanExpert-Italic
mntri9s  TimesNewRomanSF-Italic
mntrk8r  MTimesNewRomanGreek
mntrko   MTimesNewRomanGreek-Inclined
mntroz8r MTimesNewRomanCyrillic-Inc
mntrz8r  MTimesNewRomanCyrillic
mnts8r   TimesNRMT-SemiBold
mntsi8r  TimesNRMT-SemiBoldItalic
mntx8r   TimesNRMT-ExtraBold
mo3r8r   ColonnaMT
mo4b8r   OceanSansMT-Bold
mo4bi8r  OceanSansMT-BoldItalic
mo4ji8r  OceanSansMT-BookItalic
mo4k8r   OceanSansMT-Book
mo4l8r   OceanSansMT-Light
mo4li8r  OceanSansMT-LightItalic
mo4s8r   OceanSansMT-SemiBold
mo4se8r  OceanSansMT-SemiBoldItalic
mo4x8r   OceanSansMT-ExtraBold
mo4xi8r  OceanSansMT-ExtraBoldItalic
moer8r   OldEnglishTextMT MOldEnglishText
molb7a   ColmcilleAltMT-Bold
molb8r   ColmcilleMT-Bold
molbi7a  ColmcilleAltMT-BoldItalic
molbi8r  ColmcilleMT-BoldItalic
molr7a   ColmcilleAltMT
molr8r   ColmcilleMT
molri7a  ColmcilleAltMT-Italic
molri8r  ColmcilleMT-Italic
molrp8r  ColmcilleOrnamentsMT
mosb8r   OldStyleMT-Bold MOldStyle-Bold
mosbi8r  OldStyleMT-BoldItalic MOldStyle-BoldItalic
mosbl8r  MOldStyle-BoldOutline
mosr8r   OldStyleMT MOldStyle
mosri8r  OldStyleMT-Italic MOldStyle-Italic
motb8r   CoronetMT-Bold
movr8r   OctavianMT
movr8x   OctavianMT-Expert
movrc8r  OctavianMT-SC
movri7d  OctavianMT-ItalicOsF
movri8r  OctavianMT-Italic
movri8x  OctavianMT-ItalicExpert
moxr8r   MOnyx OnyxMT

mp1r8r   PalaceScriptMT
mp1rs8r  PalaceScriptMT-SemiBold
mp2r8r   PepitaMT
mp6b8r   BlueprintMT-Bold
mp6bi8r  BlueprintMT-BoldItalic
mp6r8r   BlueprintMT
mp6ri8r  BlueprintMT-Italic
mpdb8rc  PlacardMT-CondensedBold
mpdr8rc  PlacardMT-Condensed
mphb8r   PhotinaMT-Bold
mphbi8r  PhotinaMT-BoldItalic
mphr8r   PhotinaMT
mphri8r  PhotinaMT-Italic
mphs8r   PhotinaMT-SemiBold MPhotina-SemiBold
mphsi8r  PhotinaMT-SemiBoldItalic MPhotina-SemiBoldItalic
mphu8r   PhotinaMT-UltraBold MPhotina-UltraBold
mphui8r  PhotinaMT-UltraBoldItalic MPhotina-UltraBoldItalic
mpi001   RococoOrnamentsOne
mpi002   RococoOrnamentsTwo
mpi003   ArabesqueOrnamentsOne
mpi101   AlmanacMT Almanac
mpi102   BotanicalMT Botanical
mpi103   DirectionsMT Directions
mpi104   SportsOne SportsOneMT
mpi105   SportsTwo SportsTwoMT
mpi106   TransportMT Transport
mpi107   VacationMT Vacation
mpi108   SportsThree SportsThreeMT
mpi109   SportsFour SportsFourMT
mpi201   CircleFrameMT CircleFrame
mpi202   CircleFrameNegMT CircleFrameNeg
mpi203   SquareFrameMT SquareFrame
mpi204   SquareFrameNegMT SquareFrameNeg
mpi205   DiamondBodoni DiamondBodoni
mpi206   DiamondPositive DiamondPositive
mpi207   DiamondNegative DiamondNegative
mpi208   OvalFrameMT circled numerals
mpi209   OvalFrameNegativeMT ditto
mpi300   Giddyup-Thangs
mpi301   Petrucchio
mpi302   AquaLifeMT fish
mpi303   BonApetit food
mpi304   DinosaursMT dinosaurs
mpi305   FrontierMT cowboys & indians
mpi306   Holidays clip art
mpi307   KeystrokesExtrasMT keyboards
mpi308   KeystrokesMT ditto

mpi309 Melodies musical instruments

mpi310   MilestonesOneMT greeting card illos
mpi311   MilestonesTwoMT ditto
mpi312   MonkeyBusiness animals
mpi313   MonotypeSorts dingbats
mpi314   NatureMT animals
mpi315   Parties clip art
mpi319   PlaytimeMT toys
mpi320   SignsMT signs
mpi321   SouthwestOrnamentsMT geometrics
mpi322   KeystrokesShadowExtrasMT
mpi323   KeystrokesShadowMT
mpnb8r   Plantin-Bold
mpnb8rc  Plantin-BoldCondensed
mpnb8x   PlantinExp-Bold PlantinExpert-Bold
mpnbd8rc PlantinHeadlineMT-BoldCond
mpnbi8r  Plantin-BoldItalic
mpnbi8x  PlantinExp-BoldItalic PlantinExpert-BoldItalic
mpnbin8r PlantinMT-SchlbkBoldItalic
mpnbn8r  PlantinMT-SchlbkBold
mpnl8r   Plantin-Light
mpnl8x   PlantinExp-Light MPlantinExpt-Light
mpnld8rc PlantinHeadlineMT-LightCond
mpnli8r  Plantin-LightItalic
mpnli8x  PlantinExp-LightItalic MPlantinExpt-LightItalic
mpnmd8rc PlantinHeadlineMT-MedCond
mpnr8r   Plantin
mpnr8x   PlantinExp PlantinExpert
mpnra8x  PlantinExp-AltFigures
mpnri8r  Plantin-Italic
mpnri8x  PlantinExp-Italic PlantinExpert-Italic
mpnrin8r PlantinMT-SchlbkItalic
mpnrn8r  PlantinMT-Schlbk
mpns8r   Plantin-Semibold
mpns8x   PlantinExp-SemiBold MPlantinExpt-SemiBold
mpnsi8r  Plantin-SemiboldItalic Plantin-SemiboldItalic
mpnsi8x  PlantinExp-SemiBoldItalic MPlantinExpt-SemiBoldItalic
mppb8r   Perpetua-Bold
mppb8x   PerpetuaExpert-Bold MPerpetuaExpert-Bold
mppbd8r  PerpetuaTitlingMT-Bold
mppbi8r  Perpetua-BoldItalic
mppbi8x  PerpetuaExpert-BoldItalic MPerpetuaExpert-BoldItalic
mppld8r  PerpetuaTitlingMT-Light
mppr8r   Perpetua
mppr8x   PerpetuaExpert MPerpetuaExpert
mpprd8r  PerpetuaTitlingMT
mppri8r  Perpetua-Italic
mppri8x  PerpetuaExpert-Italic MPerpetuaExpert-Italic
mpum     PopplLaudatio-Medium

mpumi    PopplLaudatio-MediumItalic
mpur     PopplLaudatio-Regular
mpuri    PopplLaudatio-Italic
mpwb8r   PerrywoodMT-Bold
mpwbi8r  PerrywoodMT-BoldItalic
mpwl8r   PerrywoodMT-Light
mpwli8r  PerrywoodMT-LightItalic
mpwr8r   PerrywoodMT
mpwri8r  PerrywoodMT-Italic
mpws8r   PerrywoodMT-SemiBold
mpwsi8r  PerrywoodMT-SemiBoldItalic
mpwx8r   PerrywoodMT-ExtraBold
mpwxi8r  PerrywoodMT-ExtraBoldItalic
mpzr8r   PoliphilusMT MPoliphilus
mpzr8x   PoliphilusExpertMT
mqtr     Boulevard
mr2b8r   CantoriaMT-Bold
mr2bi8r  CantoriaMT-BoldItalic
mr2l8r   CantoriaMT-Light
mr2li8r  CantoriaMT-LightItalic
mr2r8r   CantoriaMT
mr2ri8r  CantoriaMT-Italic
mr2s8r   CantoriaMT-SemiBold
mr2si8r  CantoriaMT-SemiBoldItalic
mr2x8r   CantoriaMT-ExtraBold
mr2xi8r  CantoriaMT-ExtraBoldItalic
mr5rw8r  FrenchScriptMT
mrur8rc  RunicMT-Condensed
mrwb8r   Rockwell-Bold
mrwb8rc  Rockwell-BoldCondensed
mrwbi8r  Rockwell-BoldItalic
mrwl8r   Rockwell-Light
mrwli8r  Rockwell-LightItalic
mrwr8r   Rockwell
mrwr8rc  Rockwell-Condensed
mrwri8r  Rockwell-Italic
mrwx8r   Rockwell-ExtraBold
ms1r8r   ScotchRomanMT MScotchRoman
ms1ri8r  ScotchRomanMT-Italic MScotchRoman-Italic
ms2b8r   ScriptMT-Bold
ms7bai   SassoonInfant-BoldAlt
ms7bais  SassoonSanSlope-BoldAlt
ms7bap   SassoonPrimary-BoldAlt
ms7bas   SassoonSans-BoldAlt
ms7bi    SassoonInfant-Bold
ms7bis   SassoonSanSlope-Bold
ms7bp    SassoonPrimary-Bold
ms7bs    SassoonSans-Bold

ms7mais  SassoonSanSlope-MediumAlt
ms7mas   SassoonSans-MediumAlt
ms7mis   SassoonSanSlope-Medium
ms7ms    SassoonSans-Medium
ms7rai   SassoonInfant-Alt
ms7rais  SassoonSanSlope-Alt
ms7rap   SassoonPrimary-Alt
ms7ras   SassoonSans-Alt
ms7ri    SassoonInfant
ms7ris   SassoonSanSlope
ms7rp    SassoonPrimary
ms7rs    SassoonSans
msbr8r   SabonMT MSabon
msbrc7d  SabonMT-SCOSF
msbri7d  SabonMT-ItalicOSF
msbri8r  SabonMT-Italic MSabon-Italic
msbs7c   SabonMT-SemiBoldOSF
msbs8r   SabonMT-SemiBold MSabon-SemiBold
msmr8r   SpectrumMT MSpectrum
msmr8x   SpectrumMT-Expert
msmrc8r  SpectrumMT-SC
msmri7d  SpectrumMT-ItalicOsF
msmri8r  SpectrumMT-Italic MSpectrum-Italic
msmri8x  SpectrumMT-ItalicExpert
msms7d   SpectrumMT-SemiBoldOsF
msms8r   SpectrumMT-SemiBold MSpectrum-SemiBold
msms8x   SpectrumMT-SemiBoldExpert
msn8r1   MSpartan-SizeOne
msn8r2   MSpartan-SizeTwo
msn8r3   MSpartan-SizeThree
msn8r4   MSpartan-SizeFour   SpartanFourMT
msnb8r1  MSpartan-BoldSizeOne
msnb8r12 SpartanOneTwoMT-Bold
msnb8r13 SpartanOneThreeMT-Bold
msnb8r2  MSpartan-BoldSizeTwo
msnb8r24 SpartanTwoFourMT-Bold
msnb8r3  MSpartan-BoldSizeThree
msnb8r34 SpartanThreeFourMT-Bold
msnb8r4  SpartanFourMT-Bold MSpartan-BoldSizeFour
msni8r12 SpartanOneTwoMT-Italic
msni8r13 SpartanOneThreeMT-Italic
msni8r24 SpartanTwoFourMT-Italic
msni8r34 SpartanThreeFourMT-Italic
msnr8r12 SpartanOneTwoMT
msnr8r13 SpartanOneThreeMT
msnr8r24 SpartanTwoFourMT
msnr8r34 SpartanThreeFourMT
msnri8r1 MSpartan-ItalicSizeOne

msnri8r2 MSpartan-ItalicSizeTwo
msnri8r3 MSpartan-ItalicSizeThree
msnri8r4 SpartanFourMT-Italic MSpartan-ItalicSizeFour
mswb8r   SwingMT-Bold
msyr     MSymbolPS SymbolPS
mtlr8r   MCastellar CastellarMT
mtmbg8r  TimesSmallTextMT-Bold MTimesSmallText-Bold
mtmrg8r  TimesSmallTextMT MTimesSmallText
mtmrgi8r TimesSmallTextMT-Italic MTimesSmallText-Italic
mtwb8r   TwCenMT-Bold MTwentiethCentury-Bold
mtwb8rc  TwCenMT-CondensedBold MTwentiethCentury-BoldCondensed
mtwbi8r  TwCenMT-BoldItalic MTwentiethCentury-BoldItalic
mtwl8r   TwCenMT-Light MTwentiethCentury-Light
mtwli8r  TwCenMT-LightItalic MTwentiethCentury-LightItalic
mtwm8r   TwCenMT-Medium MTwentiethCentury-Medium
mtwm8rc  TwCenMT-CondensedMedium MTwentiethCentury-MediumCondensed
mtwmi8r  TwCenMT-MediumItalic MTwentiethCentury-MediumItalic
mtws8r   TwCenMT-SemiBold MTwentiethCentury-SemiBold
mtwsm8r  TwCenMT-SemiMedium MTwentiethCentury-SemiMedium
mtwu8r   TwCenMT-UltraBold MTwentiethCentury-UltraBold
mtwu8rc  TwCenMT-CondensedUltraBold MTwentiethCentury-UltraBoldCondensed
mtwui8r  TwCenMT-UltraBoldItalic MTwentiethCentury-UltraBoldItalic
mtwx8r   TwCenMT-ExtraBold MTwentiethCentury-ExtraBold
mtwx8rc  TwCenMT-CondensedExtraBold MTwentiethCentury-ExtraBoldCondensed
mtwxi8r  TwCenMT-ExtraBoldItalic MTwentiethCentury-ExtraBoldItalic
mtyr8r   TypewriterMT
mtyre8r  TypewriterEliteMT
mtyrs8r  TypewriterGothicMT
mu0b8r   ColumbusMT-Bold
mu0b8x   ColumbusExpMT-Bold
mu0bi8r  ColumbusMT-BoldItalic
mu0bi8x  ColumbusExpMT-BoldItalic
mu0r8r   ColumbusMT
mu0r8x   ColumbusExpMT
mu0ri8r  ColumbusMT-Italic
mu0ri8x  ColumbusExpMT-Italic
mu0rp8r1 ColumbusOrnamentsOneMT
mu0rp8r2 ColumbusOrnamentsTwoMT
mu0s8r   ColumbusMT-SemiBold
mu0s8x   ColumbusExpMT-SemiBold
mu0si8r  ColumbusMT-SemiBoldItalic
mu0si8x  ColumbusExpMT-SemiBoldItal
murb7a   CentaurMT-AltBold
murb8r   CentaurMT-Bold
murb8x   CentaurExpertMT-Bold
murbi7a  CentaurMT-AltBoldItalic
murbi8r  CentaurMT-BoldItalic
murbi8x  CentaurExpertMT-BoldItalic

murr7a   CentaurMT-Alt
murr8r   CentaurMT
murr8x   CentaurExpertMT
murrai   CentaurMT-ItalicA
murri7a  CentaurMT-AltItalic
murri8r  CentaurMT-Italic
murri8x  CentaurExpertMT-Italic
murrw8r  CentaurSwashMT
mvdr8r   VanDijckMT MVanDijck
mvdr8x   VanDijckExpertMT
mvdra8x  VanDijckExpertMT-AltFigs
mvdrai8x VanDijckExpertMT-AltFigsIta
mvdri8r  VanDijckMT-Italic
mvdri8x  VanDijckExpertMT-Italic
mvird8rc VictoriaTitlingMT-Condensed
mwbm8r   WalbaumMT-Medium MWalbaum-Medium
mwbm8x   WalbaumExpMT-Medium
mwbmi8r  WalbaumMT-MediumItalic MWalbaum-MediumItalic
mwbmi8x  WalbaumExpMT-MediumItalic
mwbr8r   WalbaumMT MWalbaum
mwbr8x   WalbaumExpMT
mwbri8r  WalbaumMT-Italic MWalbaum-Italic
mwbri8x  WalbaumExpMT-Italic
mwnr8r   WindsorMT
mwrb8r   WitternbergerFraktMT-Bold
mwrr8r   WitternbergerFraktMT
mygb8r   CenturyGothic-Bold
mygbi8r  CenturyGothic-BoldItalic
mygr8r   CenturyGothic
mygri8r  CenturyGothic-Italic
myhb8r   StrayhornMT-Bold
myhbi8r  StrayhornMT-BoldItalic
myhl8r   StrayhornMT-Light
myhli8r  StrayhornMT-LightItalic
myhr8r   StrayhornMT-Regular
myhri8r  StrayhornMT-Italic
myhx8r   StrayhornMT-ExtraBold
myhxi8r  StrayhornMT-ExtraBoldItalic
mzab8r   MZAntiqua-Bold
mzabi8r  MZAntiqua-BoldItalic
mzar8r   MZAntiqua
mzari8r  MZAntiqua-Italic
mzcb8r   ZapfChancery-Bold
mzcd8r   ZapfChancery-Demi
mzcmi8r  MZapfChanceryPS-MediumItalic
mzdr     MZapfDingbatsPS
mztb8r   ZeitgeistMT-Bold
mztb8x   ZeitgeistExpertMT-Bold

mztr7a   ZeitgeistMT-Alternative
mztr8r   ZeitgeistMT
mztr8rc  ZeitgeistMT-Condensed
mztr8x   ZeitgeistExpertMT
mztr8xc  ZeitgeistExpertMT-Condensed
mztri8r  ZeitgeistMT-Italic
mztri8x  ZeitgeistExpertMT-Italic
zmztcm8r ZeitgeistMT-Cameo
zmztcz8r ZeitgeistMT-CrazyPaving
zmztr8xc ZeitgeistExpertMT-Cameo

A.8 URW fonts

The ‘urw.map’ file has abbreviations for URW fonts.

uagd8r   URWGothicL-Demi
uagdo8r  URWGothicL-DemiObli
uagk8r   URWGothicL-Book
uagko8r  URWGothicL-BookObli
ubkd8r   URWBookmanL-DemiBold
ubkdi8r  URWBookmanL-DemiBoldItal
ubkl8r   URWBookmanL-Ligh
ubkli8r  URWBookmanL-LighItal
ucrb8r   NimbusMonL-Bold
ucrbo8r  NimbusMonL-BoldObli
ucrr8r   NimbusMonL-Regu
ucrro8r  NimbusMonL-ReguObli
uhvb8r   NimbusSanL-Bold
uhvbo8r  NimbusSanL-BoldItal
uhvboc8r NimbusSanL-ReguCondItal
uhvbrc8r NimbusSanL-BoldCond
uhvr8r   NimbusSanL-Regu
uhvro8r  NimbusSanL-ReguItal
uhvro8rc NimbusSanL-BoldCondItal
uhvr8rc  NimbusSanL-ReguCond
uncb8r   CenturySchL-Bold
uncbi8r  CenturySchL-BoldItal
uncr8r   CenturySchL-Roma
uncri8r  CenturySchL-Ital
uplb8r   URWPalladioL-Bold
uplbi8r  URWPalladioL-BoldItal
uplr8r   URWPalladioL-Roma
uplri8r  URWPalladioL-Ital
usyr     StandardSymL
utmb8r   NimbusRomNo9L-Medi
utmbi8r  NimbusRomNo9L-MediItal
utmr8r   NimbusRomNo9L-Regu
utmri8r  NimbusRomNo9L-ReguItal
uzcmi8r  URWChanceryL-MediItal

Appendix A: Font name lists 81

     uzdr     Dingbats

Appendix B Encodings

The following sections define various encodings as PostScript encoding vectors. These files can be installed in the Dvips header directory and then used in ‘psfonts.map’ to change font encodings of downloadable or builtin fonts (see section “psfonts.map” in Dvips).

You can also specify these ‘.enc’ files to Afm2tfm when building a font (see section “Changing font encodings” in Dvips).

B.1 ‘8a.enc’: Adobe standard encoding

% @psencodingfile{
%   author = "Adobe",
%   version = "1.0",
%   date = "22 June 1996",
%   filename = "8a.enc",
%   email = "kb@mail.tug.org",
%   address = "135 Center Hill Rd. // Plymouth, MA 02360",
%   codetable = "ISO/ASCII",
%   checksum = "269     317    2399",
%   docstring = "Adobe StandardEncoding as an afm2tfm/dvips .enc file."
% }
/StandardEncoding [
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/space
/exclam
/quotedbl
/numbersign
/dollar
/percent
/ampersand
/quoteright
/parenleft
/parenright
/asterisk
/plus
/comma
/hyphen
/period
/slash
/zero
/one
/two
/three
/four
/five
/six
/seven
/eight
/nine
/colon
/semicolon
/less
/equal
/greater
/question
/at
/A
/B
/C
/D
/E
/F
/G
/H

/I
/J
/K
/L
/M
/N
/O
/P
/Q
/R
/S
/T
/U
/V
/W
/X
/Y
/Z
/bracketleft
/backslash
/bracketright
/asciicircum
/underscore
/quoteleft
/a
/b
/c
/d
/e
/f
/g
/h
/i
/j
/k
/l
/m
/n
/o
/p
/q
/r
/s
/t
/u
/v
/w

/y
/z
/braceleft
/bar
/braceright
/asciitilde
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/exclamdown
/cent
/sterling
/fraction
/yen
/florin
/section
/currency

/quotesingle
/quotedblleft
/guillemotleft
/guilsinglleft
/guilsinglright
/fi
/fl
/.notdef
/endash
/dagger
/daggerdbl
/periodcentered
/.notdef
/paragraph
/bullet
/quotesinglbase
/quotedblbase
/quotedblright
/guillemotright
/ellipsis
/perthousand
/.notdef
/questiondown
/.notdef
/grave
/acute
/circumflex
/tilde
/macron
/breve
/dotaccent
/dieresis
/.notdef
/ring
/cedilla
/.notdef
/hungarumlaut
/ogonek
/caron
/emdash
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef

/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/AE
/.notdef
/ordfeminine
/.notdef
/.notdef
/.notdef
/.notdef
/Lslash
/Oslash
/OE
/ordmasculine
/.notdef
/.notdef
/.notdef
/.notdef
/.notdef
/ae
/.notdef
/.notdef
/.notdef
/dotlessi
/.notdef
/.notdef
/lslash
/oslash
/oe
/germandbls
/.notdef
/.notdef
/.notdef
/.notdef
] def

B.2 ‘8r.enc’: TEX base encoding

% @psencodingfile{
%   author = "S. Rahtz, P. MacKay, Alan Jeffrey, B. Horn, K. Berry",
%   version = "0.6",
%   date = "22 June 1996",
%   filename = "8r.enc",

%   email = "kb@mail.tug.org",
%   address = "135 Center Hill Rd. // Plymouth, MA 02360",
%   codetable = "ISO/ASCII",
%   checksum = "119     662    4424",
%   docstring = "Encoding for TrueType or Type 1 fonts to be used with TeX."
% }
%
% Idea is to have all the characters normally included in Type 1 fonts
% available for typesetting. This is effectively the characters in Adobe
% Standard Encoding + ISO Latin 1 + extra characters from Lucida.
%
% Character code assignments were made as follows:
%
% (1) the Windows ANSI characters are almost all in their Windows ANSI
% positions, because some Windows users cannot easily reencode the
% fonts, and it makes no difference on other systems. The only Windows
% ANSI characters not available are those that make no sense for
% typesetting -- rubout (127 decimal), nobreakspace (160), softhyphen
% (173). quotesingle and grave are moved just because it's such an
% irritation not having them in TeX positions.
%
% (2) Remaining characters are assigned arbitrarily to the lower part
% of the range, avoiding 0, 10 and 13 in case we meet dumb software.
%
% (3) Y&Y Lucida Bright includes some extra text characters; in the
% hopes that other PostScript fonts, perhaps created for public
% consumption, will include them, they are included starting at 0x12.
%
% (4) Remaining positions left undefined are for use in (hopefully)
% upward-compatible revisions, if someday more characters are generally
% available.
%
% (5) hyphen appears twice for compatibility with both ASCII and Windows.
%
/TeXBase1Encoding [
% 0x00 (encoded characters from Adobe Standard not in Windows 3.1)
  /.notdef /dotaccent /fi /fl
  /fraction /hungarumlaut /Lslash /lslash
  /ogonek /ring /.notdef
  /breve /minus /.notdef
% These are the only two remaining unencoded characters, so may as
% well include them.
  /Zcaron /zcaron
% 0x10
 /caron /dotlessi
% (unusual TeX characters available in, e.g., Lucida Bright)
 /dotlessj /ff /ffi /ffl
 /.notdef /.notdef /.notdef /.notdef

/.notdef /.notdef /.notdef /.notdef
% very contentious; it's so painful not having quoteleft and quoteright
% at 96 and 145 that we move the things normally found there down to here.
/grave /quotesingle
% 0x20 (ASCII begins)
/space /exclam /quotedbl /numbersign
/dollar /percent /ampersand /quoteright
/parenleft /parenright /asterisk /plus /comma /hyphen /period /slash
% 0x30
/zero /one /two /three /four /five /six /seven
/eight /nine /colon /semicolon /less /equal /greater /question
% 0x40
/at /A /B /C /D /E /F /G /H /I /J /K /L /M /N /O
% 0x50
/P /Q /R /S /T /U /V /W
/X /Y /Z /bracketleft /backslash /bracketright /asciicircum /underscore
% 0x60
/quoteleft /a /b /c /d /e /f /g /h /i /j /k /l /m /n /o
% 0x70
/p /q /r /s /t /u /v /w
/x /y /z /braceleft /bar /braceright /asciitilde
/.notdef % rubout; ASCII ends
% 0x80
/.notdef /.notdef /quotesinglbase /florin
/quotedblbase /ellipsis /dagger /daggerdbl
/circumflex /perthousand /Scaron /guilsinglleft
/OE /.notdef /.notdef /.notdef
% 0x90
/.notdef /.notdef /.notdef /quotedblleft
/quotedblright /bullet /endash /emdash
/tilde /trademark /scaron /guilsinglright
/oe /.notdef /.notdef /Ydieresis
% 0xA0
/.notdef % nobreakspace
/exclamdown /cent /sterling
/currency /yen /brokenbar /section
/dieresis /copyright /ordfeminine /guillemotleft
/logicalnot
/hyphen % Y&Y (also at 45); Windows' softhyphen
/registered
/macron
% 0xD0
/degree /plusminus /twosuperior /threesuperior
/acute /mu /paragraph /periodcentered
/cedilla /onesuperior /ordmasculine /guillemotright
/onequarter /onehalf /threequarters /questiondown
% 0xC0
/Agrave /Aacute /Acircumflex /Atilde /Adieresis /Aring /AE /Ccedilla

 /Egrave /Eacute /Ecircumflex /Edieresis
 /Igrave /Iacute /Icircumflex /Idieresis
% 0xD0
 /Eth /Ntilde /Ograve /Oacute
 /Ocircumflex /Otilde /Odieresis /multiply
 /Oslash /Ugrave /Uacute /Ucircumflex
 /Udieresis /Yacute /Thorn /germandbls
% 0xE0
 /agrave /aacute /acircumflex /atilde
 /adieresis /aring /ae /ccedilla
 /egrave /eacute /ecircumflex /edieresis
 /igrave /iacute /icircumflex /idieresis
% 0xF0
 /eth /ntilde /ograve /oacute
 /ocircumflex /otilde /odieresis /divide
 /oslash /ugrave /uacute /ucircumflex
 /udieresis /yacute /thorn /ydieresis
] def

B.3 ‘cork.enc’: Cork encoding

%   This is the Cork encoding, aka EC, Cork, aka TeX Latin 1, as
%   described in TTN 1(4), December 1992, and TUGboat 11(4), November
%   1990, page 514. EC is the name of the Computer Modern-based fonts in
%   this encoding. NFSS and Fontinst know it by `T1'.
%
%   These are the ligs and kerns in addition to those found in the
%   AFM file.  (The only ligs in the Times-Roman.afm file are the
%   fi and fl ligatures.)
%
% LIGKERN space l =: lslash ; space L =: Lslash ;
% LIGKERN question quoteleft =: questiondown ; exclam quoteleft =: exclamdown ;
% LIGKERN hyphen hyphen =: endash ; endash hyphen =: emdash ;
% LIGKERN quoteleft quoteleft =: quotedblleft ;
% LIGKERN quoteright quoteright =: quotedblright ;
%
%   We blow away kerns to and from spaces (TeX doesn't have a
%   space) and also remove any kerns from the numbers (although
%   the only kern pair that mentions a number in Times-Roman.afm
%   is one one.)
%
% LIGKERN space  * ; *  space ; zero  * ; *  zero ;
% LIGKERN one  * ; *  one ; two  * ; *  two ;
% LIGKERN three  * ; *  three ; four  * ; *  four ;
% LIGKERN five  * ; *  five ; six  * ; *  six ;
% LIGKERN seven  * ; *  seven ; eight  * ; *  eight ;

% LIGKERN nine  * ; *  nine ;
% LIGKERN comma comma =: quotedblbase ; less less =: guillemotleft ;
% LIGKERN greater greater =: guillemotright ;
%
/CorkEncoding [          % now 256 chars follow
% 0x00
/grave /acute /circumflex /tilde /dieresis /hungarumlaut /ring /caron
/breve /macron /dotaccent /cedilla
/ogonek /quotesinglbase /guilsinglleft /guilsinglright
% 0x10
/quotedblleft /quotedblright /quotedblbase /guillemotleft
/guillemotright /endash /emdash /compworkmark
/perthousandzero /dotlessi /dotlessj /ff /fi /fl /ffi /ffl
% 0x20
/visualspace /exclam /quotedbl /numbersign
/dollar /percent /ampersand /quoteright
/parenleft /parenright /asterisk /plus /comma /hyphen /period /slash
% 0x30
/zero /one /two /three /four /five /six /seven
/eight /nine /colon /semicolon /less /equal /greater /question
% 0x40
/at /A /B /C /D /E /F /G /H /I /J /K /L /M /N /O
% 0x50
/P /Q /R /S /T /U /V /W
/X /Y /Z /bracketleft /backslash /bracketright /asciicircum /underscore
% 0x60
/quoteleft /a /b /c /d /e /f /g /h /i /j /k /l /m /n /o
% 0x70
/p /q /r /s /t /u /v /w
/x /y /z /braceleft /bar /braceright /asciitilde /hyphen
% 0x80
/Abreve /Aogonek /Cacute /Ccaron /Dcaron /Ecaron /Eogonek /Gbreve
/Lacute /Lcaron /Lslash /Nacute /Ncaron /Ng /Ohungarumlaut /Racute
% 0x90
/Rcaron /Sacute /Scaron /Scedilla
/Tcaron /Tcedilla /Uhungarumlaut /Uring
/Ydieresis /Zacute /Zcaron /Zdotaccent /IJ /Idotaccent /dbar /section
% 0xA0
/abreve /aogonek /cacute /ccaron /dcaron /ecaron /eogonek /gbreve
/lacute /lcaron /lslash /nacute /ncaron /ng /ohungarumlaut /racute
% 0xB0
/rcaron /sacute /scaron /scedilla
/tquoteright /tcedilla /uhungarumlaut /uring
/ydieresis /zacute /zcaron /zdotaccent
/ij /exclamdown /questiondown /sterling
% 0xC0
/Agrave /Aacute /Acircumflex /Atilde /Adieresis /Aring /AE /Ccedilla
/Egrave /Eacute /Ecircumflex /Edieresis

  /Igrave /Iacute /Icircumflex /Idieresis
% 0xD0
  /Eth /Ntilde /Ograve /Oacute /Ocircumflex /Otilde /Odieresis /OE
  /Oslash /Ugrave /Uacute /Ucircumflex /Udieresis /Yacute /Thorn /Germandbls
% 0xE0
  /agrave /aacute /acircumflex /atilde /adieresis /aring /ae /ccedilla
  /egrave /eacute /ecircumflex /edieresis
  /igrave /iacute /icircumflex /idieresis
% 0xF0
  /eth /ntilde /ograve /oacute /ocircumflex /otilde /odieresis /oe
  /oslash /ugrave /uacute /ucircumflex /udieresis /yacute /thorn /germandbls
] def

B.4 ‘dvips.enc’: Dvips encoding

% @psencodingfile{
%   author = "T. Rokicki, K. Berry",
%   version = "2.0",
%   date = "22 June 1996",
%   filename = "dvips.enc",
%   email = "kb@mail.tug.org",
%   address = "135 Center Hill Rd. // Plymouth, MA 02360",
%   codetable = "ISO/ASCII",
%   checksum = "75     515    3715",
%   docstring = "An encoding for PostScript fonts to be used with TeX."
% }
%
% This is almost the default encoding produced by the afm2tfm program
% that comes with the dvips distribution, when it is applied to a
% StandardEncoding PostScript font.  The positions of the accented
% characters (which afm2tfm assigned more or less randomly in the
% previous set of fonts) have been held fixed.
%
% It is almost upward-compatible with TeX text.  (The primary exceptions
% are dotaccent and hungarumlaut.)  Some of the ligatures are also in
% different positions, but this is not important, since ligature
% positions are defined in the TFM file, not by macros.
%
% The character positions were assigned in this order:
%   1) non-composite characters in the afm2tfm encoding (except dotlessj
%      was included in its plain TeX position);
%   2) available characters in the Cork (EC) encoding;
%   3) remaining composite characters in the base encoding (starting at
%      position 130 decimal).  We don't include all the other base
%      characters because that changes the height/depth/width tables in
%      the TFM files (because more characters mean more distinct values).

%
% We leave the initial positions open for the Greek letters in Times-Roman.
%
/DvipsEncoding [
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef
  /.notdef /quotesingle /exclamdown /questiondown
/dotlessi /dotlessj /grave /acute /caron /breve /macron /ring
/cedilla /germandbls /ae /oe /oslash /AE /OE /Oslash
/space /exclam /quotedbl /numbersign /dollar /percent /ampersand /quoteright
/parenleft /parenright /asterisk /plus /comma /hyphen /period /slash
/zero /one /two /three /four /five /six /seven
/eight /nine /colon /semicolon /less /equal /greater /question
/at /A /B /C /D /E /F /G
/H /I /J /K /L /M /N /O
/P /Q /R /S /T /U /V /W
/X /Y /Z /bracketleft /backslash /bracketright /circumflex /underscore
/quoteleft /a /b /c /d /e /f /g
/h /i /j /k /l /m /n /o
/p /q /r /s /t /u /v /w
/x /y /z /braceleft /bar /braceright /tilde /dieresis
/asciicircum /asciitilde /Ccedilla /Iacute
    /Icircumflex /atilde /edieresis /egrave
/scaron /zcaron /Eth /ff /ffi /ffl /.notdef /.notdef
/.notdef /.notdef /Scaron /.notdef /.notdef /.notdef /.notdef /.notdef
/Ydieresis /.notdef /Zcaron  /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /cent /sterling /fraction /yen /florin /section
/currency /copyright /quotedblleft /guillemotleft
    /guilsinglleft /guilsinglright /fi /fl
/degree /endash /dagger /daggerdbl /periodcentered /.notdef /paragraph /bullet
/quotesinglbase /quotedblbase /quotedblright /guillemotright
    /ellipsis /perthousand /.notdef /.notdef
/Agrave /Aacute /Acircumflex /Atilde /Adieresis /Aring /.notdef /dotaccent
/Egrave /Eacute /Ecircumflex /Edieresis
    /Igrave /hungarumlaut /ogonek /Idieresis
/emdash /Ntilde /Ograve /Oacute /Ocircumflex /Otilde /Odieresis /.notdef
/.notdef /Ugrave /Uacute /Ucircumflex /Udieresis /Yacute /Thorn /.notdef
/agrave /aacute /acircumflex /ordfeminine
    /adieresis /aring /.notdef /ccedilla
/Lslash /eacute /ecircumflex /ordmasculine
    /igrave /iacute /icircumflex /idieresis
/.notdef /ntilde /ograve /oacute /ocircumflex /otilde /odieresis /.notdef
/lslash /ugrave /uacute /ucircumflex /udieresis /yacute /thorn /ydieresis
] def

B.5 ‘texmext.enc’: TEX math symbol encoding

% TeX math extension encoding.

% No ligatures or kerns.
%
/TeXMathExtensionEncoding [          % now 256 chars follow
% 0x0
/parenleftbig
/parenrightbig
/bracketleftbig
/bracketrightbig
/floorleftbig
/floorrightbig
/ceilingleftbig
/ceilingrightbig
% 0x8
/braceleftbig
/bracerightbig
/angbracketleftbig
/angbracketrightbig
/vextendsingle
/vextenddouble
/slashbig
/backslashbig
% 0x10
/parenleftBig
/parenrightBig
/parenleftbigg
/parenrightbigg
/bracketleftbigg
/bracketrightbigg
/floorleftbigg
/floorrightbigg
% 0x18
/ceilingleftbigg
/ceilingrightbigg
/braceleftbigg
/bracerightbigg
/angbracketleftbigg
/angbracketrightbigg
/slashbigg
/backslashbigg
% 0x20
/parenleftBigg
/parenrightBigg
/bracketleftBigg

/bracketrightBigg
/floorleftBigg
/floorrightBigg
/ceilingleftBigg
/ceilingrightBigg
% 0x28
/braceleftBigg
/bracerightBigg
/angbracketleftBigg
/angbracketrightBigg
/slashBigg
/backslashBigg
/slashBig
/backslashBig
% 0x30
/parenlefttp
/parenrighttp
/bracketlefttp
/bracketrighttp
/bracketleftbt
/bracketrightbt
/bracketleftex
/bracketrightex
% 0x38
/bracelefttp
/bracerighttp
/braceleftbt
/bracerightbt
/braceleftmid
/bracerightmid
/braceex
/arrowvertex
% 0x40
/parenleftbt
/parenrightbt
/parenleftex
/parenrightex
/angbracketleftBig
/angbracketrightBig
/unionsqtext
/unionsqdisplay
% 0x48
/contintegraltext

/contintegraldisplay
/circledottext
/circledotdisplay
/circleplustext
/circleplusdisplay
/circlemultiplytext
/circlemultiplydisplay
% 0x50
/summationtext
/producttext
/integraltext
/uniontext
/intersectiontext
/unionmultitext
/logicalandtext
/logicalortext
% 0x58
/summationdisplay
/productdisplay
/integraldisplay
/uniondisplay
/intersectiondisplay
/unionmultidisplay
/logicalanddisplay
/logicalordisplay
% 0x60
/coproducttext
/coproductdisplay
/hatwide
/hatwider
/hatwidest
/tildewide
/tildewider
/tildewidest
% 0x68
/bracketleftBig
/bracketrightBig
/floorleftBig
/floorrightBig
/ceilingleftBig
/ceilingrightBig
/braceleftBig
/bracerightBig

% 0x70
/radicalbig
/radicalBig
/radicalbigg
/radicalBigg
/radicalbt
/radicalvertex
/radicaltp
/arrowvertexdbl
% 0x78
/arrowtp
/arrowbt
/bracehtipdownleft
/bracehtipdownright
/bracehtipupleft
/bracehtipupright
/arrowdbltp
/arrowdblbt
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def

B.6 ‘texmsym.enc’: TEX math symbol encoding

% TeX math symbols encoding.
% No ligatures or kerns.
%
/TeXMathSymbolEncoding [          % now 256 chars follow
% 0x0
/minus

/periodcentered
/multiply
/asteriskmath
/divide
/diamondmath
/plusminus
/minusplus
% 0x8
/circleplus
/circleminus
/circlemultiply
/circledivide
/circledot
/circlecopyrt
/openbullet
/bullet
% 0x10
/equivasymptotic
/equivalence
/reflexsubset
/reflexsuperset
/lessequal
/greaterequal
/precedesequal
/followsequal
% 0x18
/similar
/approxequal
/propersubset
/propersuperset
/lessmuch
/greatermuch
/precedes
/follows
% 0x20
/arrowleft
/arrowright
/arrowup
/arrowdown
/arrowboth
/arrownortheast
/arrowsoutheast
/similarequal

% 0x28
/arrowdblleft
/arrowdblright
/arrowdblup
/arrowdbldown
/arrowdblboth
/arrownorthwest
/arrowsouthwest
/proportional
% 0x30
/prime
/infinity
/element
/owner
/triangle
/triangleinv
/negationslash
/mapsto
% 0x38
/universal
/existential
/logicalnot
/emptyset
/Rfractur
/Ifractur
/latticetop
/perpendicular
% 0x40
/aleph
/A
/B
/C
/D
/E
/F
/G
% 0x48
/H
/I
/J
/K
/L
/M
/N

/O
% 0x50
/P
/Q
/R
/S
/T
/U
/V
/W
% 0x58
/X
/Y
/Z
/union
/intersection
/unionmulti
/logicaland
/logicalor
% 0x60
/turnstileleft
/turnstileright
/floorleft
/floorright
/ceilingleft
/ceilingright
/braceleft
/braceright
% 0x68
/angbracketleft
/angbracketright
/bar
/bardbl
/arrowbothv
/arrowdblbothv
/backslash
/wreathproduct
% 0x70
/radical
/coproduct
/nabla
/integral
/unionsq

/intersectionsq
/subsetsqequal
/supersetsqequal
% 0x78
/section
/dagger
/daggerdbl
/paragraph
/club
/diamond
/heart
/spade
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def

B.7 ‘texmital.enc’: TEX math italic encoding

% TeX math italic encoding.
% No ligatures or kerns.
%
/TeXMathItalicEncoding [          % now 256 chars follow
% 0x0
/Gamma
/Delta
/Theta
/Lambda
/Xi
/Pi
/Sigma

/Upsilon
% 0x8
/Phi
/Psi
/Omega
/alpha
/beta
/gamma
/delta
/epsilon1
% 0x10
/zeta
/eta
/theta
/iota
/kappa
/lambda
/mu
/nu
% 0x18
/xi
/pi
/rho
/sigma
/tau
/upsilon
/phi
/chi
% 0x20
/psi
/omega
/epsilon
/theta1
/pi1
/rho1
/sigma1
/phi1
% 0x28
/arrowlefttophalf
/arrowleftbothalf
/arrowrighttophalf
/arrowrightbothalf
/arrowhookleft

/arrowhookright
/triangleright
/triangleleft
% 0x30
/zerooldstyle
/oneoldstyle
/twooldstyle
/threeoldstyle
/fouroldstyle
/fiveoldstyle
/sixoldstyle
/sevenoldstyle
% 0x38
/eightoldstyle
/nineoldstyle
/period
/comma
/less
/slash
/greater
/star
% 0x40
/partialdiff
/A
/B
/C
/D
/E
/F
/G
% 0x48
/H
/I
/J
/K
/L
/M
/N
/O
% 0x50
/P
/Q
/R

/S
/T
/U
/V
/W
% 0x58
/X
/Y
/Z
/flat
/natural
/sharp
/slurbelow
/slurabove
% 0x60
/lscript
/a
/b
/c
/d
/e
/f
/g
% 0x68
/h
/i
/j
/k
/l
/m
/n
/o
% 0x70
/p
/q
/r
/s
/t
/u
/v
/w
% 0x78

/y
/z
/dotlessi
/dotlessj
/weierstrass
/vector
/tie
% Have to have 128 .notdef's now to fill things out.
%
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
/.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef /.notdef
] def

Appendix C Font legalities

This section owes a great deal to Charles Bigelow (co-designer with Kris Holmes of the Lucida typeface family, among others), who has generously answered our many queries about fonts and the law around the world with remarkable patience and understanding. (But he is naturally not responsible for any errors here, much less our opinions.)

Fonts have always been treated rather strangely under the law, as befits their rather strange nature: letterforms are indivisibly both useful and artistic. In most countries—in all countries until recently—utility has taken precedence; i.e., it has been legal to copy fonts without permission or fee.

In any case, to the best of my knowledge, the situation in those countries which have adopted any sort of typeface protection is as follows:

United States Typeface designs can be patented, but not copyrighted. Only a few designs have been patented. (Lucida and Stone are the only ones I know of. I don’t know what the grounds were for patenting Stone, but Lucida had some novel features in its design which make it reasonable to patent, given the acceptance of the patent system in the first place.) Particular programs which instantiate a font can be copyrighted just as any other computer program can. This is arguably wrong, since font programs are nothing but a description of the shapes, possibly with some simple hints, and there’s only one basic way to describe the shapes in any given language. Thus, the creativity lies in making the shape right, not in making the computer program right, so it would seem that to be consistent, the copyright laws should protect the design, not the program—the opposite of the current situation.

Germany Typeface designs have been copyrightable as original works of art since 1981. As a result, many (perhaps most) designs have been copyrighted there. The law is not retroactive, but German courts have upheld the intellectual property rights of designers of earlier fonts as well: In one case the heirs of Paul Bauer (designer of Futura) won a suit against the Bauer foundry for arbitrarily discontinuing a portion of their royalties.

England A copyright law passed in 1989 covers typeface designs first published in England (or published in Britain within 30 days of its publication elsewhere), and it is retroactive. It’s unclear how far back the law extends, but Times Roman, designed in the late 1920’s and 1930’s by Stanley Morison and cut by Victor Lardent for Monotype is probably covered. This does not mean GNU (for example) cannot have a Times Roman; it just means we cannot start with an English version, as the law does not forbid importing foreign versions of English typefaces.

France The Romain du Roi typeface designed by Philippe Grandjean in 1702 for the French royal family is protected, and perhaps other such “royal” designs. There is no protection for designs in general.

In 1973 the international Vienna treaty on typeface design protection was proposed. France ratified it in 1974 or 1975, and Germany in 1981. The English law might constitute ratification, but this has not been settled. In any case, since at least four countries have to ratify it before it takes effect (and even then it takes effect only in those countries which ratify it), it is still of no consequence for now.

Index 107

Index

3

35 standard PostScript fonts . . . . . . . . . . . . . . . . . . . . . . . 21

7

7-bit encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

8

8-bit encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8a.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 8r.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

9

9s variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

A

Adobe standard encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 82 adobe.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Afm2tfm encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 apple.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 attribution of fonts, missing . . . . . . . . . . . . . . . . . . . . . . . . . 3 automatic expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Avant Garde. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

B

base encoding, for TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Bauer, Paul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Beebe, Nelson. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Beeton, Barbara . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Bernstein, Rocky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Bigelow, Charles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 bitstrea.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 bizarre font names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Bouche, Thierry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 bright variant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

C

category codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 CD-ROM distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 character scaling, automatic. . . . . . . . . . . . . . . . . . . . . . . . 18 compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 contributors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 copyright on fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Cork encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 cork.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Cugley, Damian. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Cyrillic encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Cyrillic fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

D

design size specification . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 DOS filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 dtc.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Dviljk translator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Dvips encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Dvips translator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 dvips.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

E

EC encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 eight-character limitation . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 encoding files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 encoding vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 encoding, Dvips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 encodings, Adobe standard. . . . . . . . . . . . . . . . . . . . . . . . . 82 encodings, Cork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 encodings, EC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 encodings, TEX base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 encodings, TEX Latin 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 encodings, TEX math italic . . . . . . . . . . . . . . . . . . . . . . . 101 encodings, TEX math symbol . . . . . . . . . . . . . . . . . . . 93, 97 encodings, unspecified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 England, font design protection law . . . . . . . . . . . . . . . 106 escape for nonstandard names . . . . . . . . . . . . . . . . . . . . . . . 3 expansion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 expert encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 expertised encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 expertised font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 extension, TEX math encoding . . . . . . . . . . . . . . . . . . . . . 93

F

fax variant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 filenames, basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 font copyrighting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 font scaling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 font coding scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Fontinst distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 FontName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 fontname mapping file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Fontname version 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 foundries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 France, font design protection law. . . . . . . . . . . . . . . . . 106 ftp.cs.umb.edu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ftp.tug.org . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Futura . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

G

Garamond, versions of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Germany, font design protection law . . . . . . . . . . . . . . 106 Grandjean, Philippe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Greek fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

H

hexadecimal design size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 history. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Holmes, Kris . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Horn, Berthol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

I

individual font designers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 informal variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ISO 9660 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 italic, TEX math encoding . . . . . . . . . . . . . . . . . . . . . . . . 101

J

Jeffrey, Alan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

K

Knuth, Donald E.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

L

Lang, Russell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Lardent, Victor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 LaserJet 4 fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 legalisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 letterspacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ligatures and encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 linotype.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 long naming scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 lslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Lslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Lubalin, Herb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Lucida Bright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Lucida, protection of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

M

MacKay, Pierre . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 mapping file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 mapping files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 math extension, TEX encoding . . . . . . . . . . . . . . . . . . . . . 93 math italic, TEX encoding . . . . . . . . . . . . . . . . . . . . . . . . 101 math symbol, TEX encoding . . . . . . . . . . . . . . . . . . . . . . . 97 math variants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Mittelbach, Frank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 3, 16 modes.mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Monotype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 monotype.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Morison, Stanley . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 multiple variants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

N

nonstandard names, escape for . . . . . . . . . . . . . . . . . . . . . . 3

P

patenting of fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 phonetic encodings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 pi fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 PostScript encoding vectors . . . . . . . . . . . . . . . . . . . . . . . . 82 PostScript fonts, standard . . . . . . . . . . . . . . . . . . . . . . . . . 21 PSfonts distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

Rahtz, Sebastian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 raw fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 related files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Rivlin, Jean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Rokicki, Tom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 romain du roi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

S

sans variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Schoepf, Rainer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 schoolbook variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 semisans variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 semiserif variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 small caps fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 small foundries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 standard PostScript fonts . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Stone Informal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Stone, protection of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 SuperFont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 supplier.map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 suppliers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 symbol, TEX math encoding . . . . . . . . . . . . . . . . . . . . . . . 97

T

T1.etx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 TDS standard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 terminology. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 tex-fonts@mail.tug.org . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 TEX base encoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 TEX Directory Structure standard . . . . . . . . . . . . . . . . . . . 1 TEX Latin 1 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 TEX math extension encoding . . . . . . . . . . . . . . . . . . . . . . 93 TEX math italic encoding . . . . . . . . . . . . . . . . . . . . . . . . . 101 TEX math symbol encoding . . . . . . . . . . . . . . . . . . . . . . . . 97 texfonts.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 texmext.enc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 texmital.enc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 texmsym.enc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Times Roman, protection of . . . . . . . . . . . . . . . . . . . . . . 106 type foundries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 typeface design protection . . . . . . . . . . . . . . . . . . . . . . . . 106 typeface.map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 typewriter variant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

U

unattributed fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 United States, font design protection law . . . . . . . . . . 106 urw.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

V

variant.map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Vienna treaty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

W

weight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 weight.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 width.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Windows ANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

X

X Window System font names . . . . . . . . . . . . . . . . . . . . . 19

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Filenames for fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Suppliers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Typefaces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Weights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.4 Variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.5 Widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3 Long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

3.1 A fontname mapping file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2 A naming scheme for long names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Appendix A Font name lists . . . . . . . . . . . . . . . . . . . . . 21

A.1 Standard PostScript fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 A.2 Adobe fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 A.3 Apple fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 A.4 Bitstream fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 A.5 DTC fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 A.6 Linotype fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 A.7 Monotype fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 A.8 URW fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

Appendix B Encodings . . . . . . . . . . . . . . . . . . . . . . . . . . 82

B.1 ‘8a.enc’: Adobe standard encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 B.2 ‘8r.enc’: TEX base encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 B.3 ‘cork.enc’: Cork encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 B.4 ‘dvips.enc’: Dvips encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 B.5 ‘texmext.enc’: TEX math symbol encoding . . . . . . . . . . . . . . . . . . . . . 93 B.6 ‘texmsym.enc’: TEX math symbol encoding . . . . . . . . . . . . . . . . . . . . . 97 B.7 ‘texmital.enc’: TEX math italic encoding. . . . . . . . . . . . . . . . . . . . . 101

Appendix C Font legalities. . . . . . . . . . . . . . . . . . . . . . 106

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107